### v1

# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

Abstract
Aiming to accelerate the training of large deep neural models in an energy-efficient way, an analog in-memory computing (AIMC) accelerator emerges as a solution with immense potential. In AIMC accelerators, trainable weights are kept in memory without the need to move from memory to processors during the training, significantly reducing overhead. However, although the in-memory feature enables efficient computation, it also constrains the use of data parallelism since copying weights from one AIMC to another is expensive. To enable parallel training using AIMC, we propose synchronous and asynchronous pipeline parallelism for AIMC accelerators inspired by the pipeline in digital domains. This paper provides a theoretical convergence guarantee for both synchronous and asynchronous pipelines in terms of both sampling and clock cycle complexity, which is non-trivial since the physical characteristic of AIMC accelerators leads to analog updates that suffer from asymmetric bias. The simulations of training deep models on real datasets verify the efficiency of pipeline analog training.
1 Introduction
Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications. As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive. For example, it requires 184 thousand GPU hours to train an LLAMA2 7 billion model, and this time increases to 1.7 million GPU hours for its 70 billion version [1]. Stemming from the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators emerges. In principle, ideal specialized hardware for DNN should be efficient, support large-scale parallelism, and remain lower energy-economic simultaneously. As the demand for efficient processing intensifies, the fundamental backbone of both training and inference largely relies on matrix-vector multiplication (MVM) operation, which is the core component of the ubiquitous fully connected and convolutional layers. Typically, each forward and backward pass requires on the order of millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
The promise of analog computing. To this end, a promising solution to accelerate the MVM operations is the analog in-memory computating (AIMC) accelerator, wherein both the input and output of MVM operation are analog signals, and the trainable weights of DNN layers are stored in analog crossbar tiles [2]. Different from traditional Von Neumann accelerators like GPU and TPU, which require weight movement from memory to processor before computation, the AIMC accelerator keeps the weight stationary and performs MVM operations in memory. Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Figure 1.
The curse of data parallelism. While efficient computation is a benefit, in-memory computing imposes limitations on scalability. Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training. In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different computation devices to support gradient descent with large batches [4, 5]. On the contrary, in AIMC architecture, the trainable weights are stored in the analog tiles in the form of device conductances, and thus, copying weights from one analog tile to another is expensive and prone to error. Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
Due to the physical limitations of the AIMC accelerator, we instead resort to pipeline parallelism [6], another prevalent parallel mechanism. In pipeline parallelism, a large model is partitioned into a series of smaller stages and gets distributed across multiple devices, with each stage corresponding to the computation of a few layers. After a device finishes the process of one datum (or one batch), it passes the intermediate outputs to the next devices and begins the process of the next data. In this way, the model that is too large to fit into one device can be separated into different AIMC devices, enlarging the efficient memory. Meanwhile, the computation on different devices can happen simultaneously, and the systems can process more data in a given period of time.
While pipelining is a simple idea, implementing it in analog training faces a series of crucial challenges. First, AIMC accelerators usually suffer from noisy gradient signals and asymmetric updates [7, 8], which introduce errors on all forward, backward, and update passes. Second, to fully realize the potential of pipeline parallelism, the weight update pass usually involves stale gradients. Both of them incur errors in the training, leading to the following question:
How to design analog pipeline parallel training approaches that do not impair the training?
This paper answers this question by analyzing the sample complexity of different pipeline algorithms and wall-clock time. Before that, we will briefly review prior work that attempted to solve the issues.
Figure 1: (Left) Illustration of MVM computation in AIMC accelerators. The weight $W^{(m)}$ on layer $m$ is stored in a crossbar tile consisting of an array of resistors. The $(i,j)$-th element of $W^{(m)}$ is represented by the conductance of the $(i,j)$-th resistor. To perform an MVM operation $z^{(m)}=W^{(m)}x^{(m)}$, voltage $[x^{(m)}]_{j}$ is applied between $j$-th and $(j+1)$-th row. By Ohm’s law, the current is $I_{ij}=[W^{(m)}]_{ij}[x^{(m)}]_{j}$; and by Kirchhoff’s law, the total current on the $i$-th column is $\sum_{j}I_{ij}=\sum_{j}[W^{(m)}]_{ij}[x^{(m)}]_{j}$. Unlike the digital counterpart, no movement of $W^{(m)}$ is required. (Right) Speedup of the proposed pipeline training methods as the number of stages increases compared with single machine training case (stage=1); see Section 5 for more details.
Gradient-based training on AIMC accelerators. Analog SGD implemented by pulse update is extremely efficient [8]. However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [7]. To overcome these issues, a family of analog training algorithms, Tiki-Taka (TT) has been proposed. TT-v1 [9] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error. To demonstrate its efficiency, [10] establishes a model to characterize the dynamic of the analog training and provides a theoretical convergence guarantee. To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [11] introduces a digital array. Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [12] are proposed to mitigate the impact of inaccurate symmetric point correction [13]. Despite the above efforts, analog training of large models on huge datasets is still a major challenge.
Synchronous and asynchronous parallelism. Typically, existing pipeline designs can be classified into two categories. If the weights are updated using the latest gradients, it is synchronous pipeline; otherwise, it is asynchronous pipeline. As the computing requirement of training large models emerges, a plethora of pipeline parallel libraries has been developed, including Gpipe [14, 15], Deepspeed [16], FSDP [17], Megetron-LM [18], Colossal-AI [19], just to name a few. Nevertheless, they are synchronous pipeline that enjoys the same dynamic as SGD, so they have the same convergence guarantee as SGD. Apart from that, researchers also made some effort to explore the asynchronous pipeline. For instance, [20] studied the asynchronous pipeline setting and proposed weight-prediction approaches to compensate for the delay. PipeDream [21] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup. To mitigate the negative effect of stale gradients, [22] and [23] have proposed only overlapping gradient computation and communication so that delay is always one. However, these algorithms are not for analog pipeline training. Researchers also attempted to design analog pipeline circuits to accelerate the training [24, 25], but they focus on the synchronous pipeline. On the theory side, there are a few studies on pipeline training on digital devices such as [26], but these results rely on weight stashing, which is expensive for analog accelerators. As [21] points out, delay affects both forward and backward passes in the asynchronous pipeline training. It renders the increment no longer a valid gradient without the weight stashing and makes the analysis challenging.
In the context of prior work, we summarize our key contributions as follows.
  1. C1)
    To fully embrace the promise of analog computing, we systematically study the pipeline training on AIMC accelerators, which enables training larger models by distributing models across multiple accelerators and increasing computation density. We study the existing work and classify them as synchronous pipelines, which achieve considerable speedup, but their computation density is still sub-optimal. To further increase the computation density, we propose the first asynchronous pipeline training algorithm on AIMC accelerators.
  2. C2)
    We provide a convergence guarantee for the proposed algorithms. Theoretical results show that two strategies have pros and cons: synchronous pipeline has a better complexity while its computation density in the device is lower; asynchronous pipeline reduces device idleness but involves stale gradients. The main results are summarized in Table 1.
  3. C3)
    We conduct empirical simulation to demonstrate the efficiency of pipeline training. The results show that both pipeline algorithms significantly outperform their non-pipeline counterparts. Despite a slightly larger sample complexity, the asynchronous pipeline achieves the same accuracy with fewer clock cycles because of its high computation density. Furthermore, the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices; see Figure 1.
Pipeline Type Sample Comp. Comp. Density Clock Cycle
w.o. Pipeline $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{1}{M}$ $O\left(M\sigma^{2}\varepsilon^{-2}\right)$
Synchronous [Theorem 1] $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{B}{M+B-1}$ $O\left((M+B)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$
Asynchronous [Theorem 2] $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$ $1$ $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$
Table 1: Comparision of different pipeline strategies. Sample complexity (Sample Comp.) is the number of samples needed for achieving training loss $\leq\varepsilon$. Computation density (Comp. Density) is the time that is the portion of time when the system is computing the gradients. In the table, $M$ is the pipeline depth, and $B$ is the micro-batch size. Clock Cycle measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by clock cycle.
2 Model Training on AIMC Accelerators
In this section, we first introduce the generic gradient-based training algorithm, encompassing digital algorithms running on devices like GPU and TPU, and algorithms running on AIMC accelerators. Consider a standard model training problem on a data distribution ${\mathcal{D}}$, given by
$W^{*}:=\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}f(W)~{}~{}~{}{\rm with}~{}~{}~{}f(W):={\mathbb{E}}_{\xi\sim{\mathcal{D}}}[f(W;\xi)]$ (1)
where $W$ is all trainable weights of the model vectorized into a $D$ dimensional vector.
DNN model. Consider the multi-layer DNN, which consists only of linear layers and activation functions. Let $[M]:=\{1,2,\cdots,M\}$ denote a positive integer set. The trainable weights are separated into $M$ different blocks, i.e., $W=[W^{(1)},W^{(2)},\cdots,W^{(M)}]$, where $W^{(m)},\forall m\in[M]$ is a matrix. The training of DNN typically consists of three phases: forward, backward, and update. Suppose $x$ and $y$ are the feature and label of data $\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$, where ${\mathcal{X}}$ and ${\mathcal{Y}}$ are feature and label space, respectively. The forward pass of the model for layer $m\in[M]$ can be expressed by
$\textsf{Forward}\qquad x^{(1)}=x,~{}~{}~{}~{}$ $z^{(m)}=W^{(m)}x^{(m)},~{}~{}~{}~{}x^{(m+1)}=g^{(m)}(z^{(m)})$ (2)
where $g^{(m)}$ is the activation function of layer $m$, and $x^{(M)}$ is the feature from the last layer, which is fed together with label $y$ into the loss function $\ell(\cdot,\cdot):{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$ and yields $f(W;\xi)=\ell(x^{(M+1)},y)$. For the backward pass, define the backward error $\delta^{(m)}:=\frac{\partial f(W;\xi)}{\partial z^{(m)}}$, wherein the error of last layer is first computed by $\delta^{(M)}=\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})$. Subsequently, each layer $m\in[M-1]$ computes the next recursion according to
$\textsf{Backward}\qquad\delta^{(m)}=\frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}=(\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$ (3)
In the forward pass, the activation $z^{(m)}$ is stashed until the backward signal $\delta^{(m+1)}$ returns. Typically, the computation $\nabla g^{(m)}(z)$ is light-weight, and hence the major overhead of the backward pass lies at the MVM among $W$ and $\delta$. With the forward and backward signals, the gradient with respect to the weight $W^{(m)}$ of layer $m$ is computed by
$\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=\frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}=\delta^{(m)}\otimes x^{(m)}$ (4)
where $\otimes$ is the outer product between two vectors. Having the gradients of each layer, we can use different gradient-based algorithms to update the weights.
Gradient-based training on AIMC accelerators. After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights. In digital devices, one of the most popular training algorithms is SGD (SGD), wherein the matrix $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$ is computed and added to $W^{(m)}_{k}$, written by $W_{k+1}^{(m)}=W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})$; $\alpha>0$ is a positive learning rate. In the paper, the subscript $k$ is used to indicate the index of gradient update.
Different from digital devices, AIMC accelerator uses rank-update to update the gradient update, which constructs a series of pulses and sends them to its analog array [8]. Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, AIMC accelerators perform the update operation in a fast and energy-efficient way. Nevertheless, rank-update relies on the special structure that the gradient of $W^{(m)}_{k}$ is a rank-one matrix, which brings us further challenges.
On top of that, as existing work pointed out [9, 10], the update on AIMC accelerators suffers from asymmetric bias, and thus the dynamic of Analog SGD in single-machine training can be written as
$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})-\frac{\alpha}{\tau}|\nabla_{W^{(m)}}f(W_{k};\xi_{k})|\odot W_{k}^{(m)}$ (5)
where $\tau$ is a device-dependent constant. When $\tau\rightarrow\infty$, Analog SGD reduces to Digital SGD.
The main challenge coming from the asymmetric is the saturation phenomenon [9]. Consider the update of $(i,j)$-th element of matrix $W_{k}^{(m)}$, whose update is written as
$([W_{k+1}^{(m)}]_{ij}-[W_{k}^{(m)}]_{ij})^{2}=$ $\ \alpha^{2}(\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})-[W^{(m)}_{k}]_{ij}/\tau)^{2}(\nabla_{W^{(m)}}f(W_{k};\xi_{k}))^{2}$
$\approx$ 0if [$\ 0~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if }[W^{(m)}_{k}]_{ij}\approx\tau~{}\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})$ (6)
which means that the increment is suppressed when $[W^{(m)}_{k}]_{ij}$ gets close to the $\tau$. Therefore, we usually introduce a proper scaling coefficient so that the saturation never happens during the training.
Figure 2: Illustration of pipelines with 4 devices ($M=4$). Each mini-batch is split into $B=4$ micro-batches. Each color represents one micro-batch, and each row from bottom to top represents stages 0 to 4. Each column corresponds to a clock cycle in which one micro-batch is processed. The white square indicates the idle device. (Top) Vanilla model parallelism without pipeline. All weights are updated using gradient-based algorithms after the computation of one mini-batch is completed, which is not presented in the figure. (Middle) Synchronous pipeline with micro-batch count $B=5$. The update happens after all 5 micro-batches are processed. (Bottom) Asynchronous pipeline. The update happens once the gradient of one micro-batch is achieved. The grey square indicates the corresponding device is processing data that is not fully reflected in the figure.
3 Analog Pipeline Model Training
This section introduces pipeline parallelism and shows how it increases computational efficiency.
Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one device and motivates us to adopt model parallelism. In model parallelism, the computation is partitioned into a sequence of stages, which, in general, consists of a consecutive set of layers in the model. Each stage is mapped on one digital or analog device. For simplicity, we assume each stage contains one layer only. Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$. For simplicity, let the micro-batch size $B_{\text{micro}}=1$ in Sections 3-4 so that the subscript $b$ in $\xi_{k,b}$ can represent the index of micro-batch or data, and $B=B_{\text{mini}}$ can represent the number of micro-batch and mini-batch size $B_{\text{mini}}$. It could be easily extended to general cases.
Clock cycle. We define the maximum time for each stage to process one micro-batch as the clock cycle. Ideally, all devices would process a micro-batch in the same amount of time, avoiding idle periods during a clock cycle. Although achieving this is challenging, various task scheduling approaches have been proposed to balance computation [21, 14]. To explore the potential of pipeline training, this paper uses the clock cycle and stage as the minimal time and basic scheduling units, respectively. Based on scheduling methods, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining.
Model parallelism without pipeline. Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data, as previously mentioned) for iteration $k$. Vanilla model parallelism processes all micro-batches one by one. In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$, while stage $m$ achieves the backward error $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage in the backward pass. The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stored when the gradient for the $b$-th micro-batch is completed. After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method. Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}$, which iterates over $b\in[B_{\text{mini}}]$ by
$W_{k,b+1}^{(m)}=W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})-\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (7)
Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$. However, for the AIMC accelerators, only rank-update is permitted so that each gradient is accumulated to the weight $W^{(m)}_{k}$ one by one, rendering the analysis more difficult.
Synchronous pipeline. Existing work has already attempted to design architecture that supports the pipeline training [24, 25]. They faithfully implement the iteration (7) and overlap parts of the computation. Since they compute all gradients in a batch on the same $W_{k}$, their algorithms can be called synchronous pipeline. In the vanilla non-pipeline version, the stage $m$ is idle after its forward computation of $b$-th micro-batch until the backward error comes back from stage $m+1$. In the synchronous pipeline training, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any; see the middle of Figure 2. Similar to the vanilla non-pipeline version, the weights will be updated after processing all micro-batches. To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch (fill and drain periods), hence rendering some devices idle.
Asynchronous pipeline. To increase the usage of all devices, a more progressive strategy, asynchronous pipeline, is proposed. Instead of waiting for the gradient of one mini-batch, the asynchronous pipeline training method performs gradient-based update algorithms immediately after one micro-batch is completed. Except for short periods of fill and drain at the beginning and the end, all devices are active during the training, maximizing the training efficiency; see the bottom of Figure 2.
Despite its perfect resource usage, asynchronous pipeline parallelism suffers from the stale signal issue. Notice that in the asynchronous pipeline setting, the forward signal $x^{(m)}$ and backward signals $\delta^{(m)}$ are computed based on different stale weights. For example, after outputting forward signal $x^{(0)}$ for the $k$-th micro-batch, the first stage ($m=1$) receives the backward signal $\delta^{(0)}$ after $2(M-1)$ clock cycles, wherein $M-1$ gradient updates happen. Consequently, the gradient is stale for the current weights. Furthermore, notice both forward (2) and backward (3) passes involve the matrix $W^{(m)}$ while the $W^{(m)}$ is changed between the forward and backward passes. To depict the process more precisely, we introduce different symbols, which have tildes on top of them to indicate the staleness. Since the mini-batch is no longer needed in asynchronous pipeline, the subscript $b$ for the micro-batch index can be omitted, and we denote the micro-batch for $k$-th update as $\xi_{k}$. To make the relationship clearer, the computation dependency is plotted in Figure 3.
Since there are $M-m$ gradient updates between the forward and backward passes in stage $m$, the weight used for the forward pass has a delay of $M-m$, that is
${\tilde{x}}^{(m+1)}_{k}=$ $\ g^{(m)}({\tilde{z}}^{(m)}_{k}),\quad{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},\quad{\tilde{x}}^{(0)}_{k}=x_{k}.$ (8)
In contrast, the weights are updated following the backward pass, which means the weight used for backward is the latest one
$\tilde{\delta}^{(m)}_{k}=$ $\ \tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k}),$ (9)
$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is no longer a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partial delay gradient, making the analysis more challenging.
Remark 1 (Digital v.s. analog asynchronous pipeline).
In digital hardware, one can stash the stale weight $W^{(m)}_{k-(M-m)}$ and read it out again to compute the backward signal in (9) [26, 21]. Weight stashing ensures the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is the gradient with respect to $W^{(m)}_{k-(M-m)}$, and hence provides a better convergence guarantee. However, copying is expensive in the analog domain, as we discussed before. Consequently, the asynchronous pipeline suffers from partial delay gradient issues.
Computation density. Different pipeline strategies can process varying amounts of data in the same amount of time. In an ideal scenario, a micro-batch requires $2M$ clock cycles for forward and backward passes, distributed equally across $M$ devices, taking 2 clock cycles per micro-batch. To measure how closely an algorithm approaches this ideal, we introduce computation density. Let $N_{\text{c}}(t)$ be the number of data points processed in $t$ clock cycles. Assuming infinite runtime, we define computation density as the limit ratio of ideal processing time to real time: $\lim_{t\to\infty}\frac{2N_{\text{c}}(t)}{t}$, where we introduce the factor $2$ since each data involves forward and backward passes. For vanilla model parallelism without pipelining, it takes $2M$ cycles per micro-batch, yielding a computation density of $1/M$. In synchronous pipelining, it takes $2(M+B-1)$ clock cycles to process $B$ micro-batches, resulting in a density of $B/(M+B-1)$. Asynchronous pipelining achieves a density of $1$, as all devices remain active during training.
Figure 3: Illustration of the dynamic of asynchronous pipeline. The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each device.
4 Convergence and Wall-clock Analysis
This section provides the convergence guarantee of the proposed pipeline training algorithm on AIMC accelerators. We show that both synchronous and asynchronous pipelines enjoy the same convergence rate as the vanilla non-pipeline version. By increasing the computation density significantly, the pipeline parallelism provides notable speedup in terms of clock cycle complexity.
4.1 Assumptions on models, data, and AIMC accelerators
[…]
Assumption 1 (Unbiasness and bounded variance).
The sample $\xi_{k}$ are independently sampled from a distribution ${\mathcal{D}}$ over times $k\in[K]$, and has bounded variance, i.e., ${\mathbb{E}}_{\xi_{k}}[\nabla f(W_{k};\xi_{k})]=\nabla f(W_{k})$ and ${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}$.
Assumption 2 (Normalized feature).
The feature is normalized, i.e. $\|x\|=1,\forall x\in{\mathcal{X}}$.
Assumption 3 (Activation function).
All activation function $g^{(m)}(\cdot)$, $m\in[M]$, is (i) Lipschitz continuous, i.e. $\|g^{(m)}(z)-g^{(m)}(z^{\prime})\|\leq L_{g,0}\|z-z^{\prime}\|$; (ii) smooth, i.e. $\|\nabla g^{(m)}(z)-\nabla g^{(m)}(z^{\prime})\|\leq L_{g,1}\|z-z^{\prime}\|$; (iii) centered, i.e. $g(0)=0$.
The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid and Tanh functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
Assumption 4 (Loss function).
The loss $\ell(\cdot,\cdot)$ is smooth with respective the first input, i.e. $\|\nabla_{1}\ell(y_{1},y)-\nabla_{1}\ell(y_{2},y)\|\leq L_{\ell,1}\|y_{1}-y_{2}\|,\forall y_{1},y_{2},y$; lower bounded, i.e. $\ell(y_{1},y_{2})\geq f^{*},\forall y_{1},y_{2}$.
Note that the lower bounded loss implies the lower boundness of the objectives as well, i.e. $f(W;\xi)\geq f^{*}$, $\forall W\in{\mathbb{R}}^{D},x\in{\mathcal{X}}$. Aside from the assumptions, to avoid saturation during training, we also need proper scaling so that the saturation degree of the AIMC accelerators is bounded.
Assumption 5 (Bounded Saturation).
There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$ norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max}<\tau,\forall k\in[K]$. The ratio $W_{\max}/\tau$ is referred to as saturation degree.
Assumption 5 requires the AIMC accelerators to have a sufficiently large active region to represent weights. Besides, on digital devices, the trainable weights always remain bounded in a small neighborhood of the initialization points during training [27, 28, 29], especially for overparameterized networks. Given $W_{\max}$ is sufficiently large to include this neighborhood, Assumption 5 holds. While it is possible to prove this assumption, we adopt it to make our results more concise and to focus on revealing the key insights on the effect of pipeline training.
Under the above assumptions, we prove the objective $f(W;\xi)$ is smooth.
Lemma 1 (smoothness of the objective).
Under Assumption 25, the objective is $L_{f}$-smooth with respect to $W$, i.e. $\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$.
The proof of Lemma 1 is deferred to Appendix C. Lemma 1 enables us to adapt the tools from [30] for our analysis. However, due to the asymmetric bias of the AIMC accelerators, the convergence results for Analog SGD are still non-trivial.
4.2 Convergence rate of pipeline training algorithms
In this section, we provide the convergence rate of Analog SGD with pipelines. In non-convex optimization tasks like the DNN training, instead of measuring the objective (1), it is more common to measure the gradient square norm $\|\nabla f(W_{k})\|^{2}$.
Theorem 1 (Convergence rate, synchronous pipeline).
Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{B(f(W_{0})-f^{*})}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\frac{1}{L}$, it holds that
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S$
where $S$ denotes the amplification factor given by $S:=\frac{W_{\max}/\tau^{2}}{1-W_{\max}/\tau^{2}}$.
The proof of Theorem 1 is deferred to Appendix D. Similar to the convergence of digital SGD [30], Analog SGD converges in the rate $O(\sqrt{{\sigma^{2}}/({BK})})$. However, because of the asymmetric bias, the convergence rate slows down by a factor $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$, which reduces to $1$ when $\tau$ tends to infinity. Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error. When the AIMC accelerators can be well-designed, $\tau$ can be sufficiently large, and it does not dominate the error when $K$. Empirically, researchers also proposed some Analog SGD variants like Tiki-Taka [9, 11] to eliminate the negative impact of asymmetric bias. Next, we provide the iteration complexity of Analog SGD with asynchronous pipeline training.
Theorem 2 (Convergence rate, asynchronous pipeline).
Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\frac{1}{L_{f}}$, it holds that
[…]
where $u>0$ is any number and $S^{\prime}$ denotes the amplification factor given by $S^{\prime}:=\frac{(1+u)W_{\max}^{2}/\tau^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$.
The proof of Theorem 2 is deferred to Appendix E. Because of the stale signals, the convergence rate of the asynchronous pipeline slows down. The coefficient of the dominant term degrades from $\frac{1}{1-W_{\max}/\tau^{2}}$ to $\frac{1}{1-(1+u)W_{\max}/\tau^{2}}$, and an extra error term $O\left(\frac{1+1/u}{K}\right)$ is introduced. Fortunately, $O\left(\frac{1+1/u}{K}\right)$ is not the dominant term and we can specify $u$ as a small constant so the impact is minor.
speedup accuracy (mid) accuracy (end)
Digital w.o. pipeline 1.00 88.63  $\pm$ 0.70 94.32  $\pm$ 0.12
synchronous 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12
asynchronous 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16
Analog w.o. pipeline 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07
synchronous 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07
asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07
Figure 4: Training ResNet10 on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Right) Accuracy-vs-Clock cycle. (Bottom) Speedup and test accuracy. The speedup is computed by comparing the clock cycle to reach 93% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
Remark 2.
For digital SGD with the synchronous pipeline training, the convergence rate is given in [30]. We also provide the convergence analysis for SGD with asynchronous pipeline training on digital devices, detailed in Theorem 2 in the Appendix F. Unlike analog training, both synchronous and asynchronous pipeline training do not slow down digital SGD.
4.3 Sampling complexity and clock cycle complexity
The convergence rate alone does not indicate the total training time, as different algorithms require varying numbers of gradient updates. The bottleneck during a micro-batch process is the forward and backward computations, which depend on the number of micro-batches. Thus, we focus on the sample complexity of the algorithms, defined as the minimal gradient computation needed to achieve the desired error. Due to the existence of asymptotic error, we define the sample complexity of synchronous and asynchronous pipelines are the minimum micro-batch needed for achieving $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e. $\sigma^{2}S$ for vanilla model parallelism without pipeline/synchronous pipeline and $\sigma^{2}S^{\prime}$ for asynchronous pipeline, respectively. Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, and omitting the constant $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$ or $\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}$, we have the following corollary.
Corollary 1 (Sample complexity).
The sample complexities of model parallelism without pipeline and synchronous pipeline are $O\left(\sigma^{2}\varepsilon^{-2}\right)$, while that of the asynchronous pipeline is $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$.
Corollary 1 claims that three model parallelisms share the same dominant term $O(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while asynchronous pipeline slows down by a factor and a non-dominant term $O(\varepsilon^{-1})$. Dividing the sample complexity by the computation density, we know the number of clock cycles they need to achieve the $\varepsilon$ accuracy.
Corollary 2 (Clock cycle complexity).
The clock cycle complexities of vanilla model parallelism without pipeline, synchronous pipeline, and asynchronous pipeline are $O\left(M\sigma^{2}\varepsilon^{-2}\right)$, $O\left((M+B-1)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$, and $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$, respectively.
The results are summarized in Table 1. By overlapping the computation among micro-batches, synchronous pipeline achieves $M$ times acceleration compared to its non-pipeline counterpart without any cost. Asynchronous pipeline overlaps more computation, achieving higher computation density with slight scarification of convergence rate. When the mini-batch is split into sufficient micro-batch so that $B\gg M$, the bubble time can be omitted, and synchronous pipeline achieves the best clock cycle complexity. However, when the number of mico-batch is small, the asynchronous pipeline is able to achieve better acceleration.
speedup accuracy (mid) accuracy (end)
Digital w.o. pipeline 1.00 67.85  $\pm$ 0.52 73.99  $\pm$ 0.19
synchronous 3.69 73.91  $\pm$ 0.19 73.99  $\pm$ 0.19
asynchronous 8.19 74.64  $\pm$ 0.48 74.64  $\pm$ 0.53
Analog w.o. pipeline 1.00 61.35  $\pm$ 0.73 72.29  $\pm$ 0.30
synchronous 3.69 72.14  $\pm$ 0.31 72.29  $\pm$ 0.30
asynchronous 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
Figure 5: Training ResNet10 on CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Right) Accuracy-vs-Clock cycle. (Bottom) Speedup and test accuracy. The speedup is computed by comparing the clock cycle to reach 72% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
5 Numerical Simulations
In this section, we test our methods in an open-source AIHWKIT (AIHWKIT[31]; see github.com/IBM/aihwkit. In the simulations, we train a ResNet10 to perform an image classification task on CIFAR10 and CIFAR100 datasets; see Figure 4 and 5. The network consists of $M=6$ stages, consisting 1 convolution layer, 4 residual blocks, and 1 fully-connected layer. We map the first 5 stages on digital devices and the last one onto an AIMC (analog) device. We use digital SGD, implemented by PyTorch  library. The pipeline parallelism is also simulated on a single GPU. To get the equivalent clock cycle in the figures, we map the epoch to the clock cycle by dividing the computation density, as discussed in Section 3. Each simulation is repeated three times, and the mean and standard deviation are reported. More details are reported in Appendix G.
[…]
$B_{\text{micro}}$ Synchronous Asynchronous
speedup acc (mid) acc (end) speedup acc (mid) acc (end)
D $16$ 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16
$32$ 2.67 6.06 94.27  $\pm$ 0.18 94.28  $\pm$ 0.20
$64$ 1.71 6.00 94.09  $\pm$ 0.13 94.08  $\pm$ 0.15
$128$ 1.00 5.83 93.68  $\pm$ 0.39 93.67  $\pm$ 0.38
A $16$ 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07
$32$ 2.67 93.57  $\pm$ 0.12 93.95  $\pm$ 0.07 5.88 93.62  $\pm$ 0.13 93.61  $\pm$ 0.11
$64$ 1.71 90.86  $\pm$ 0.44 93.95  $\pm$ 0.07 5.88 93.58  $\pm$ 0.14 93.60  $\pm$ 0.17
$128$ 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07 4.89 93.36  $\pm$ 0.06 93.37  $\pm$ 0.06
Table 2: The training of Resnet10 model on CIFAR10 dataset. The test accuracy and standard deviation under different micro-batch sizes at clock cycle 300 (mid) and epoch 300 (end). The speedup is computed by comparing the clock cycle needed to reach 93% target accuracy. The accuracy of synchronous is not affected by micro-batch size.
Effect of micro-batch size. We also report the effect of micro-batch sizes in Table 2. We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size. For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog devices. Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version. The trade-off is that the asynchronous pipeline sacrifices the accuracy and can only reach lower accuracy when the training time is sufficiently long.
Effect of stage/device number. We extensively explored the impact of increasing device count on the speedup achieved through pipeline parallelism; see Figure 1. In the simulations, we train Resnet34 model on CIFAR10 dataset, which could be separated into $M=1,2,4,8,12,16$ stages. Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed. The results reveal that the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices. Even though some latency, like communication latency, are ignored in the simulations, the comparison is still fair since the asynchronous pipeline does not introduce obvious extra latency.
6 Conclusions and Limitations
This paper study the pipeline algorithms for AIMC accelerators in the convergence and computation aspects. We study the existing synchronous pipeline algorithm and propose the first asynchronous pipeline algorithm on AIMC accelerators. The synchronous pipeline achieves the best convergence rate in terms of sample complexity, but training bubbles reduce computation density. To enable more devices to be active, an asynchronous pipeline discards the synchronization at the end of one mini-batch, achieving a higher computation density, with the cost of slightly increased sample complexity. Simulations reveal the extraordinary speedup potential of our asynchronous pipeline for analog computing. A limitation of this work is that AIMC accelerator behaviors can only be simulated but not on the actual analog chips. But this is a necessary step for every new technology, and future work should verify its efficiency as the technology matures.
Supplementary Material for “Pipeline Gradient-based Model Training on Analog In-memory Accelerator”
Appendix A Analog pipeline SGD from worker and data perspectives
Involving multiple devices and data at each time step, pipeline algorithms, especially asynchronous pipeline, are usually hard to figure out. To facilitate the understanding, we list the Analog SGD under two different perspectives. In worker perspective, we pay more attention to how worker $m$ processes a flow of data; see Algorithm 1. In data perspective, we focus on how a sample $\xi_{k}$ flows through each worker; see Algorithm 2.
Algorithm 1 Analog SGD with asynchronous pipeline (Worker Perspective)
1:for iteration $k=0,1,2,\cdots,K$ do
2:// Periphery device (worker 0) at the beginning
3:     Sample $\xi_{k}=(x_{k},y_{k})$ and Send $x^{(0)}_{k}=x_{k}$ to worker $1$
4:// Worker $m=1,2,\cdots,M$ in parallel
5:     Receive input ${\tilde{x}}^{(m)}_{k}$ from worker $m-1$
6:     Compute ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$
7:     Compute ${\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$ and $\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
8:     Store input ${\tilde{x}}^{(m)}_{k}$ and $\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
9:     Receive backward error $\tilde{\delta}^{(m+1)}_{k-(M-m)}$
10:     $\tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
11:     $W_{k+1}^{(m)}=W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}$
12:// Periphery device at the end
13:     Compute $\tilde{\delta}^{(M)}_{k}=\nabla_{{\tilde{x}}^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)})$
14:     Send $\tilde{\delta}^{(M)}_{k}$ back to worker $M$
15:end for
Algorithm 2 Analog SGD with asynchronous pipeline (Data Perspective)
1:for iteration $k=0,1,2,\cdots,K$ do
2:     Sample $\xi_{k}=(x_{k},y_{k})$, $x^{(0)}_{k}=x_{k}$
3:     for all worker $m=1,2,\cdots,M$ do
4:         Compute ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$
5:         Compute ${\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$ and $\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
6:     end for
7:     Compute $\tilde{\delta}^{(M)}_{k}=\nabla_{{\tilde{x}}^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)})$
8:     for all layer $m=M,M-1,\cdots,1$ do
9:         $\tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
10:         $W_{k+1}^{(m)}=W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}$
11:     end for
12:end for
Appendix B Bounds of delay errors
In the asynchronous pipeline, the computations of both forward and backward passes involve stale weights, which induces delay error on the signals. This section provides a series of the bounds of the delay error.
Lemma 2 (Lipschitz continuity of loss function).
There exists $L_{\ell,0}$ such that for any $m\in[M]$, $\|\nabla_{x^{(m)}}\ell(x_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$.
Proof.
According to norm equivalence, there exists $W_{\max,2}=cW_{\max}$ with $c$ is defined by the dimension of $W_{k}^{(m)}$, such that $\|W_{k}^{(m)}\|\leq W_{\max,2}$. For $m\in[M]$, according to the forward recursion (2), it holds that
$\|x^{(m+1)}_{k}\|\overset{(a)}{=}$ $\ \|g^{(m)}(z^{(m)}_{k})-g^{(m)}(0)\|\overset{(b)}{\leq}L_{g,0}\|z^{(m)}_{k}\|=L_{g,0}\|W^{(m)}_{k}x^{(m)}_{k}\|$ (11)
$\leq$ $\ L_{g,0}\|W^{(m)}_{k}\|\|x^{(m)}_{k}\|{\leq}L_{g,0}W_{\max,2}\|x^{(m)}_{k}\|$
where (a) and (b) hold because of Assumption 3 (1) and (2). Then by induction, we get $\|x^{(m)}_{k}\|\leq\ (L_{g,0}W_{\max,2})^{m}$ and the same bound holds for $\|x^{(m)}_{k}\|$. Since $\ell(\cdot,\cdot)$ is differentiable so that $\nabla\ell(\cdot,\cdot)$ is continuous, then there exists $L_{f,0}$ such that $\|\nabla_{x^{(m)}}\ell(x^{(m)}_{k},y^{k})\|\leq L_{f,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}^{(m)}_{k},y^{k})\|\leq L_{f,0}$ hold for any $k,m\in[M]$. ∎
Lemma 3 (boundeds signal).
Under Assumption 24, both forward and backward signals are bounded, i.e., $\forall m\in[M+1]$
$\|x^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m},$ $\|\delta^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0},$ (12)
$\|{\tilde{x}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m},$ $\|\tilde{\delta}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}.$ (13)
Proof of Lemma 3.
This lemma is proved by induction over $m$. The statement for $\|x^{(m)}_{k}\|$ and $\|{\tilde{x}}^{(m)}_{k}\|$ at $m=1$ trivially hold because $x^{(0)}_{k}={\tilde{x}}^{(0)}_{k}=x_{k}$. For $\|\delta^{(M+1)}_{k}\|$ and $\|\tilde{\delta}^{(M+1)}_{k}\|$, we have
$\|\delta^{(M+1)}_{k}\|=$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})\nabla g^{(M)}(z^{(M)}_{k})\|$ (14)
$\leq$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}(z^{(M)}_{k})\|\leq L_{\ell,0}L_{g,0}$
$\|\tilde{\delta}^{(M+1)}_{k}\|=$ $\ \|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})\|$ (15)
$\leq$ $\ \|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})\|\leq L_{\ell,0}L_{g,0}.$
For $m\in[M]$, according to the forward recursion (2), it holds that
$\|x^{(m+1)}_{k}\|\overset{(a)}{=}$ $\ \|g^{(m)}(z^{(m)}_{k})-g^{(m)}(0)\|\overset{(b)}{\leq}L_{g,0}\|x^{(m)}_{k}\|=L_{g,0}\|W^{(m)}_{k}x^{(m)}_{k}\|$ (16)
$\leq$ $\ L_{g,0}\|W^{(m)}_{k}\|\|x^{(m)}_{k}\|{\leq}L_{g,0}W_{\max,2}\|x^{(m)}_{k}\|$
where (a) and (b) hold because of Assumption 3 (1) and (2). According to the backward recursion (3), it holds that
$\|\delta^{(m)}_{k}\|=$ $\ \|\delta^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}(z^{(m)}_{k})\|$ (17)
$\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}\|,$
$\|\tilde{\delta}^{(m)}_{k}\|=$ $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|,$ (18)
$\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\tilde{\delta}^{(m+1)}_{k}\|.$
The proof has finished. ∎
We first bound the error of the stale variables. Define $\Delta^{(m)}_{k}:=W^{(m)}_{k+1}-W^{(m)}_{k}$, which is the different between consecutive weights.
Lemma 4 (Delay error of signals).
Under Assumption 24, the difference between the latest and stale signals are bounded, i.e., $\forall m\in[M+1]$
$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (19)
$\|{\tilde{z}}^{(m-1)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\sum_{m^{\prime}=1}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (20)
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (21)
where the constants above are defined by
$C_{x}^{(m)}:=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1},$ (22)
$C_{z}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{m-1},$ (23)
$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (24)
$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
Proof of Lemma 4.
We first bound the delay of forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$. For other layers $m>1$
$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|\leq L_{g,0}\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|$ (25)
$=$ $\ L_{g,0}\|W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}-W^{(m)}_{k}x^{(m)}_{k}\|$
$=$ $\ L_{g,0}\|(W^{(m)}_{k-(M-m)}-W^{(m)}_{k}){\tilde{x}}^{(m)}_{k}+W^{(m)}_{k}({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
$\overset{(a)}{\leq}$ $\ L_{g,0}(\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}\|+\|W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
$\leq$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
where (a) follows the triangle inequality and the consistence of $\ell_{2}$-matrix norm, i.e. $\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrix and vector, respectively. Using triangle inequality over the first term of RHS in (25), we have
$\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|\leq\sum_{k^{\prime}=k-(M-m)}^{k-1}\|W^{(m)}_{k^{\prime}+1}-W^{(m)}_{k^{\prime}}\|=\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|$ (26)
and consequently,
$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|$ (27)
$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m}\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
Expanding inequality (27) from $m$ to $1$ and using the fact ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$ yield
$\ \|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$ (28)
$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{k^{\prime}=k-(M-m+1)}^{k-1}\|\Delta^{(m-1)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m-1)}_{k}-x^{(m-1)}_{k}\|$
$\leq$ $\ L_{g,0}\sum_{m^{\prime}=2}^{m}(L_{g,0}W_{\max,2})^{m^{\prime}-1}(L_{g,0}W_{\max,2})^{m-m^{\prime}}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$
$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=2}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|.$
$=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
where the last inequality reindexes $m^{\prime}$ in the first summation.
Noticing that (25)-(28) also implies that
$\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$ (29)
We then bound the delay of the backward signal $\tilde{\delta}^{(m)}_{k}$ for $m\in[M-1]$
$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|=\|\delta^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}(z^{(m)}_{k})-\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (30)
$\leq$ $\ \|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|$
$\ +\|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})-\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$
$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|+W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|$
where the last inequality comes from Assumption 3 (1)-(2) and Lemma 3. Substituting (29) into (30), we can further bound (30) by
$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|$ (31)
$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
$\ +W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$=$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|+(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\times$
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
To bound the first term in the RHS of (31), we manipulate the different $\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ by
$\ \|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ (32)
$=$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})\nabla g^{(M)}(z^{(M)}_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})\|$
$\leq$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}(z^{(M)}_{k})\|$
$\ +\|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})-\nabla g^{(M)}(z^{(M)}_{k})\|$
$\overset{(a)}{\leq}$ $\ L_{\ell,1}L_{g,0}\|x^{(M+1)}_{k}-{\tilde{x}}^{(M+1)}_{k}\|+L_{\ell,0}L_{g,1}\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
where (a) comes from Assumption 3 and Assumption 4.
To bound the second term in the RHS of (31), we change the upper bound of the second summation from $m^{\prime\prime}$ to $M$ and get
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (33)
$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (34)
$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$ (35)
Substituting (32) and (33) back into (31) we have
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (36)
where the constant $C_{\delta}^{(m)}$ is defined by
$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (37)
$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
This completes the proof. ∎
Lemma 5 (delay error of gradient).
Suppose Assumption 24 hold. Denote the stale gradient by
~$\tilde{\nabla}_{W^{(m)}}^{k}f:=\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}.$ (38)
The gradient delay is bounded by
~$\|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|.$ (39)
where the constant $C_{f}^{(m)}$ is defined by
$C_{f}^{(m)}:=(L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)}.$ (40)
Consequently, the norm square of the gradient delay is bounded by
~$\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}=$ $\ C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (41)
where the constant $C_{f}$ is defined by
$C_{f}:=M\sqrt{\sum_{m=1}^{M}(C_{f}^{(m)})^{2}}.$ (42)
Proof of Lemma 5.
By definition, the delay error is bounded by
~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (43)
$=$ $\ \|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\delta^{(m)}_{k}\otimes x^{(m)}_{k}\|$
$=$ $\ \|(\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k})\otimes{\tilde{x}}^{(m)}_{k}+\delta^{(m)}_{k}\otimes({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
$\leq$ $\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}\|+\|\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
According to Lemma 3 and Lemma 4, it follows that
~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (44)
$\leq$ $\ ((L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)})\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
It completes the proof. ∎
Appendix C Smoothness of the object
This section proves Lemma 1, i.e. the object is $L$-smooth. Before that, we introduce a lemma as the step stone.
Since the definition of function smoothness involves two weights $W$, $W^{\prime}$ and both of them generate a sequence of forward and backward signals, we require symbols to represent them. Using slightly abusing notation in this section, we denote the forward signals computed by $W$ and ${\tilde{W}}$ as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
Lemma 6 (Signal stability).
Denote the forward signals computed by $W$ and ${\tilde{W}}$ as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
$\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}\leq$ $\ C_{x^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (45)
$\|{\tilde{z}}^{(m)}-z^{(m)}\|^{2}\leq$ $\ C_{z^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (46)
$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq$ $\ C_{\delta^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (47)
where the constants $C_{x}^{(m)}$ and $C_{z}^{(m)}$ are defined by
$C_{x^{\prime}}^{(m)}:=$ $\ 2L_{g,0}(L_{g,0}W_{\max,2})^{2(m-1)},$ (48)
$C_{z^{\prime}}^{(m)}:=$ $\ 2(L_{g,0}W_{\max,2})^{2(m-1)},$ (49)
and $C_{\delta^{\prime}}^{(m)}$ is defined by
$C_{\delta^{\prime}}^{(m)}$ $\ :=3(L_{g,0}W_{\max,2})^{2}(2L_{\ell,1}^{2}L_{g,0}^{4}+2L_{\ell,0}^{2}L_{g,1}^{2})(L_{g,0}W_{\max,2})^{2(M-1)}$ (50)
$\ +3(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{4}\sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\frac{1-(L_{g,0}W_{\max,2})^{2(M-m)}}{1-L_{g,0}W_{\max,2}}$
$\ +3(L_{g,0}W_{\max,2})^{2(M-1)}L_{\ell,0}^{2}L_{g,1}^{2}\sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\frac{1-(L_{g,0}W_{\max,2})^{2(M-m)}}{1-L_{g,0}W_{\max,2}}.$
Proof of Lemma 6.
We first bound the delay of forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}=x^{(1)}=x$. For other layers $m\geq 1$
$\ \|{\tilde{x}}^{(m+1)}-x^{(m+1)}\|^{2}\leq L_{g,0}^{2}\|{\tilde{z}}^{(m)}-z^{(m)}\|^{2}$ (51)
$=$ $\ L_{g,0}^{2}\|{\tilde{W}}^{(m)}{\tilde{x}}^{(m)}-W^{(m)}x^{(m)}\|^{2}$
$=$ $\ L_{g,0}^{2}\|({\tilde{W}}^{(m)}-W^{(m)}){\tilde{x}}^{(m)}+W^{(m)}({\tilde{x}}^{(m)}-x^{(m)})\|^{2}$
$\leq$ $\ 2L_{g,0}^{2}(\|{\tilde{W}}^{(m)}-W^{(m)}\|^{2}\|{\tilde{x}}^{(m)}\|^{2}+\|W^{(m)}\|^{2}\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2})$
$\leq$ $\ 2L_{g,0}^{2}((L_{g,0}W_{\max,2})^{2m}\|{\tilde{W}}^{(m)}-W^{(m)}\|^{2}+W_{\max,2}^{2}\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}).$
Expanding inequality (51) from $m$ to $1$ and using the fact ${\tilde{x}}^{(1)}=x^{(1)}=x$ yield
$\ \|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}$ (52)
$\leq$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}^{(m-1)}-W^{(m-1)}\|^{2}+2L_{g,0}^{2}W_{\max,2}^{2}\|{\tilde{x}}^{(m-1)}-x^{(m-1)}\|^{2}$
$\leq$ $\ 2L_{g,0}^{2}\sum_{m^{\prime}=2}^{m}(L_{g,0}W_{\max,2})^{2(m^{\prime}-1)}(L_{g,0}W_{\max,2})^{2(m-m^{\prime})}\|{\tilde{W}}^{(m^{\prime}-1)}-W^{(m^{\prime}-1)}\|^{2}$
$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=2}^{m}\|{\tilde{W}}^{(m^{\prime}-1)}-W^{(m^{\prime}-1)}\|^{2}$
$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (53)
where the last equality reindexes the summation. By applying the relation
$\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}\leq\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}=\|{\tilde{W}}-W\|^{2}$ (54)
we prove (45). Inferring in the similar way of (51)-(52), we obtain (46)
$\|{\tilde{z}}^{(m)}-z^{(m)}\|^{2}\leq$ $\ 2(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}-W\|^{2}$ (55)
We then bound the difference of the backward signal $\tilde{\delta}^{(m)}$ for $m\in[M-1]$
$\ \|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}=\|\tilde{\delta}^{(m+1)}{\tilde{W}}^{(m+1)}\nabla g^{(m)}(z^{(m)})-\delta^{(m+1)}W^{(m+1)}\nabla g^{(m)}(z^{(m)})\|^{2}$ (56)
$\leq$ $\ 3\|\tilde{\delta}^{(m+1)}-\delta^{(m+1)}\|^{2}\|{\tilde{W}}^{(m+1)}\|^{2}\|\nabla g^{(m)}(z^{(m)})\|^{2}$
$\ +3\|\delta^{(m+1)}\|^{2}\|{\tilde{W}}^{(m+1)}-W^{(m+1)}\|^{2}\|\nabla g^{(m)}(z^{(m)})\|^{2}$
$\ +3\|\delta^{(m+1)}\|^{2}\|W^{(m+1)}\|^{2}\|\nabla g^{(m)}(z^{(m)})-\nabla g^{(m)}({\tilde{z}}^{(m)})\|^{2}$
$\leq$ $\ 3(L_{g,0}W_{\max,2})^{2}\|\tilde{\delta}^{(m+1)}-\delta^{(m+1)}\|^{2}+3(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{4}\|{\tilde{W}}^{(m+1)}-W^{(m+1)}\|^{2}$
$\ +3W_{\max,2}^{2}(L_{g,0}W_{\max,2})^{2(M-m-1)}L_{\ell,0}^{2}L_{g,0}^{2}L_{g,1}^{2}\|z^{(m)}-{\tilde{z}}^{(m)}\|^{2}$
where the last inequality comes from Assumption 3 (1)-(2) and Lemma 3. Substituting (55) into (56), we can further bound (56) by
$\ \|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}$ (57)
$\leq$ $\ 3L_{g,0}^{2}W_{\max,2}^{2}\|\tilde{\delta}^{(m+1)}-\delta^{(m+1)}\|^{2}+3(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{4}\|{\tilde{W}}^{(m+1)}-W^{(m+1)}\|^{2}$
$\ +3W_{\max,2}^{2}(L_{g,0}W_{\max,2})^{2(M-m-1)}L_{\ell,0}^{2}L_{g,0}^{2}L_{g,1}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
$=$ $\ 3L_{g,0}^{2}W_{\max,2}^{2}\|\tilde{\delta}^{(m+1)}-\delta^{(m+1)}\|^{2}+3(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{4}\|{\tilde{W}}^{(m+1)}-W^{(m+1)}\|^{2}$
$\ +3(L_{g,0}W_{\max,2})^{2(M-1)}L_{\ell,0}^{2}L_{g,1}^{2}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
$\leq$ $\ 3(L_{g,0}W_{\max,2})^{2(M-m)}\|\delta^{(M)}-\tilde{\delta}^{(M)}\|^{2}$
$\ +3(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{4}\sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\|{\tilde{W}}^{(m^{\prime\prime}+1)}-W^{(m^{\prime\prime}+1)}\|^{2}$
$\ +3(L_{g,0}W_{\max,2})^{2(M-1)}L_{\ell,0}^{2}L_{g,1}^{2}\sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$
To bound the first term in the RHS of (57), we manipulate the different $\|\delta^{(M)}-\tilde{\delta}^{(M)}\|$ by
$\ \|\delta^{(M)}-\tilde{\delta}^{(M)}\|^{2}$ (58)
$=$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y)\nabla g^{(M)}({\tilde{z}}^{(M)})\|^{2}$
$\leq$ $\ 2\|\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y)\|^{2}\|\nabla g^{(M)}(z^{(M)})\|^{2}$
$\ +2\|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y)\|^{2}\|\nabla g^{(M)}({\tilde{z}}^{(M)})-\nabla g^{(M)}(z^{(M)})\|^{2}$
$\overset{(a)}{\leq}$ $\ 2L_{\ell,1}^{2}L_{g,0}^{2}\|x^{(M+1)}-{\tilde{x}}^{(M+1)}\|^{2}+2L_{\ell,0}^{2}L_{g,1}^{2}\|{\tilde{z}}^{(M)}-z^{(M)}\|^{2}$
$\leq$ $\ (2L_{\ell,1}^{2}L_{g,0}^{4}+2L_{\ell,0}^{2}L_{g,1}^{2})\|{\tilde{z}}^{(M)}-z^{(M)}\|^{2}$
$\leq$ $\ (2L_{\ell,1}^{2}L_{g,0}^{4}+2L_{\ell,0}^{2}L_{g,1}^{2})(L_{g,0}W_{\max,2})^{2(M-1)}\sum_{m^{\prime}=1}^{M-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$
where (a) comes from Assumption 3 and Assumption 4.
To bound the second term in the RHS of (57), we add some terms into the summation
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\|{\tilde{W}}^{(m^{\prime\prime}+1)}-W^{(m^{\prime\prime}+1)}\|^{2}$ (59)
$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{2(M-m)}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$
To bound the third term in the RHS of (57), we change the upper bound of the second summation from $m^{\prime\prime}$ to $M$ and get
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (60)
$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{2(M-m)}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$
Substituting (58)–(60) back into (57) we have
$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq C_{\delta^{\prime}}^{(m)}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (61)
This completes the proof. ∎
Lemma 7 (smoothness of the objective, restatement of Lemma 1).
Under Assumption 25, the objective is $L$-smooth with respect to $W$, i.e. $\forall\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$,
$\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$ (62)
where the smoothness constant is defined as
$L_{f}:=\sqrt{2\sum_{m=1}^{M}((L_{g,0}W_{\max,2})^{2m}C_{\delta^{\prime}}^{(m)}+(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{2}C_{x^{\prime}}^{(m)})}.$ (63)
Proof of Lemma 1 and Lemma 7.
Analogous to the proof of Lemma (6), we use $x^{(m)}$/$\delta^{(m)}$ and ${\tilde{x}}^{(m)}$/$\tilde{\delta}^{(m)}$ to denote the forward/backward signals related to $W$ and ${\tilde{W}}$, respectively. It follows by definition of gradient (4) that
$\ \|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}$ (64)
$=$ $\ \sum_{m=1}^{M}\|\nabla_{W^{(m)}}f(W;\xi)-\nabla_{W^{(m)}}f({\tilde{W}};\xi)\|^{2}$ (65)
$=$ $\ \sum_{m=1}^{M}\|\tilde{\delta}^{(m)}\otimes{\tilde{x}}^{(m)}-\delta^{(m)}\otimes x^{(m)}\|^{2}$
$=$ $\ \sum_{m=1}^{M}\|(\tilde{\delta}^{(m)}-\delta^{(m)})\otimes{\tilde{x}}^{(m)}+\delta^{(m)}\otimes({\tilde{x}}^{(m)}-x^{(m)})\|^{2}$
$\leq$ $\ 2\sum_{m=1}^{M}\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\|{\tilde{x}}^{(m)}\|^{2}+2\|\delta^{(m)}\|^{2}\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}.$
According to Lemma 3 and Lemma 6, it follows that
$\ \|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}$ (66)
$\leq$ $\ 2\sum_{m=1}^{M}((L_{g,0}W_{\max,2})^{2m}C_{\delta^{\prime}}^{(m)}+(L_{g,0}W_{\max,2})^{2(M-m)}L_{\ell,0}^{2}L_{g,0}^{2}C_{x^{\prime}}^{(m)})\|{\tilde{W}}-W\|^{2}$
$=$ $\ L_{f}^{2}\|{\tilde{W}}-W\|^{2}.$
Taking square root on both sides on (66) completes the proof. ∎
Appendix D Proof of Theorem 1: Analog SGD with synchronous pipeline
This section provides the convergence guarantee of the Analog SGD with synchronous pipeline, whose iteration is
$W_{k,b+1}^{(m)}=W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})-\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (67)
See 1
Proof of Theorem 1.
Denote the average gradient by
¯$\bar{\nabla}f(W_{k};\xi_{k}):=\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b}),$ (68)
which satisfies that ${\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\bar{\nabla}f(W_{k};\xi_{k})]=\nabla f(W_{k})$. Furthermore, the independence of sampling (c.f. Assumption 1) ensures that
$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (69)
$=$ $\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$
$=$ $\ \frac{1}{B^{2}}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$
$\leq$ $\ \sigma^{2}/B.$
The $L$-smooth assumption of the objective (c.f. Lemma 7) implies that
$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]$ (70)
$\leq$ $\ f(W_{k})+{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]+\frac{L_{f}}{2}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}\|^{2}]$
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}-(\frac{1}{2\alpha}-L_{f})\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{2}L_{f}\mathbb{E}_{\varepsilon_{k}}[\|\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]+\frac{1}{2\alpha}\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k})+\varepsilon_{k})\|^{2}$
where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
$\ \mathbb{E}_{\varepsilon_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]$ (71)
$=$ $\ \mathbb{E}_{\varepsilon_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\right\rangle]$
$=$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}-\frac{1}{2\alpha}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\frac{1}{2\alpha}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}]$
and the following inequality
$\ \frac{L_{f}}{2}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (72)
$\leq$ $\ L_{f}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{2}L{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}].$
With the learning rate $\alpha\leq\frac{1}{2L}$ and bounded variance (c.f. (69)), (70) becomes
$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]$ (73)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}+\frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}].$
According the dynamic of Analog SGD (5), the last term in the RHS of (73) is upper bounded by
$\ \frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}]$ (74)
$=$ $\ \frac{\alpha}{2\tau^{2}}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\frac{1}{B}\sum_{b=1}^{B}|\nabla f(W_{k};\xi_{k,b})|\odot W_{k,b}\|^{2}]$
$\leq$ $\ \frac{\alpha}{2\tau^{2}}\frac{1}{B}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\||\nabla f(W_{k};\xi_{k,b})|\odot W_{k,b}\|^{2}]$
$\leq$ $\ \frac{\alpha}{2\tau^{2}}\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}\frac{1}{B}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})\|^{2}].$
By the variance decomposition and Assumption 1, the term above can be bounded by
${\mathbb{E}}_{\xi_{k,b}}[\|f(W_{k};\xi_{k,b})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$ (75)
$\leq$ $\ \|\nabla f(W_{k})\|^{2}+\sigma^{2}.$
Consequently, combining (73), (74) and (75) yields
$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]$ (76)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-\frac{\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}+\frac{\alpha\sigma^{2}}{2\tau^{2}}\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}.$
Organizing the terms, taking expectation, averaging (76) over $k$ from 00 to $K-1$, we obtain
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2(f(W_{0})-{\mathbb{E}}[f(W_{k+1})])}{1-W_{\max,\infty}^{2}/\tau^{2}}+\frac{\alpha L_{f}\sigma^{2}}{B(1-W_{\max,\infty}^{2}/\tau^{2})}$ (77)
$+\frac{\sigma^{2}}{K}\sum_{k=0}^{K-1}\frac{\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}/\tau^{2}}{1-\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}/\tau^{2}}.$
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}$
where the second inequality uses $f(W_{k+1})\geq f^{*}$ and specifies $\alpha=\sqrt{\frac{B(f(W_{0})-f^{*})}{\sigma^{2}L_{f}K}}$. Now we complete the proof. ∎
Appendix E Proof of Theorem 2: Analog SGD with asynchronous pipeline
This section provides the convergence guarantee of the Analog SGD under non-convex assumption on asymmetric linear devices. See 2
Proof of Theorem 2.
The $L$-smooth assumption of the objective (c.f. Lemma 7) implies that
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]$ (78)
$\leq$ $\ f(W_{k})+\mathbb{E}_{\xi_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]+\frac{L_{f}}{2}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}-(\frac{1}{2\alpha}-L_{f})\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{2}L_{f}\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
$\ \mathbb{E}_{\xi_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]$ (79)
$=$ $\ \mathbb{E}_{\xi_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\right\rangle]$
$=$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}-\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
and the following inequality
$\frac{L_{f}}{2}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\leq$ $\ L_{f}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (80)
$\ +\alpha^{2}L_{f}\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}].$
With the learning rate $\alpha\leq\frac{1}{4L_{f}}$ and bounded variance of noise (Assumption 1), (78) becomes
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}].$ (81)
By the dynamic of Analog SGD, the last in the RHS of (E) is bounded by
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$ (82)
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})+\frac{1}{\tau}|\tilde{\nabla}^{k}f|\odot W_{k}\|^{2}]$
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})+\frac{1}{\tau}\tilde{\nabla}^{k}f\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}]$
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})+\frac{1}{\tau}(\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k}))\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}$
$\ \hskip 100.00015pt+\frac{1}{\tau}\nabla f(W_{k};\xi_{k})\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}]$
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|({\bm{1}}+\frac{1}{\tau}\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k})(\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k}))$
$\ \hskip 100.00015pt+\frac{1}{\tau}\nabla f(W_{k};\xi_{k})\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}]$
where ${\bm{1}}$ is the all-one vector. Using inequality $\|x+y\|^{2}\leq(1+\frac{1}{u})\|x\|^{2}+(1+u)\|y\|^{2}$ we obtain
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$ (83)
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})\mathbb{E}_{\xi_{k}}[\|({\bm{1}}+\frac{1}{\tau}\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k})\odot(\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k}))\|^{2}]$
$\ +\frac{\alpha}{2\tau^{2}}(1+u)\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}]$
The first term in the RHS of (83) is bounded by
$\ \frac{\alpha}{2}(1+\frac{1}{u})\mathbb{E}_{\xi_{k}}[\|({\bm{1}}+\frac{1}{\tau}\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k})\odot(\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k}))\|^{2}]$ (84)
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})\mathbb{E}_{\xi_{k}}[\|{\bm{1}}+\frac{1}{\tau}\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}_{\infty}\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}M^{2}C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
where the last inequality comes from Lemma 5.
The second term in the RHS of (83) is bounded by
$\ \frac{\alpha}{2\tau^{2}}(1+u)\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}]$ (85)
$\leq$ $\ \frac{\alpha}{2\tau^{2}}(1+u)\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})\|^{2}]\ \|W_{k}\|^{2}_{\infty}$
$\leq$ $\ \frac{\alpha}{2\tau^{2}}(1+u)\|\nabla f(W_{k})\|^{2}\|W_{k}\|^{2}_{\infty}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
where the last inequality comes from Assumption 1
${\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (86)
$\leq$ $\ \|\nabla f(W_{k})\|^{2}+\sigma^{2}.$
Plugging (83), (84) and (85) back into (82), we have
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$ (87)
$=$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ +\frac{\alpha}{2\tau^{2}}(1+u)\|\nabla f(W_{k})\|^{2}\|W_{k}\|^{2}_{\infty}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$
Substituting (87) into (E) yields
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]$ (88)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha}{2\tau^{2}}(1+u)\|\nabla f(W_{k})\|^{2}\|W_{k}\|^{2}_{\infty}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
$=$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
Construct a Lyapunov function by
${\mathbb{V}}_{k}:=$ $\ f(W_{k})-f^{*}$ (89)
$+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
Notice that the last term above has the following recursion from $k$ to $k+1$
$\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-(k+1)+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (90)
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$\ +(M-m+1)\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
According to (88), we have
$\ {\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]$ (91)
$=$ $\ f(W_{k+1})-f^{*}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-{k+1}+M+1)\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|^{2}]$
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ -\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
By the definition of $\Delta^{(m)}_{k}$ and the dynamic of Analog SGD, we bound the last term of the RHS of (91) by
$\ \sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}=\sum_{m^{\prime}=1}^{M}\|W^{(m)}_{k+1}-W^{(m)}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}$ (92)
$\leq$ $\ 2\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}+2\alpha^{2}\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}$
$\leq$ $\ 2\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}+2\alpha^{2}\sigma^{2}.$
Organizing the terms of (91), we obtain
$\ \frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}$ (93)
$\leq$ $\ {\mathbb{V}}_{k}-{\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
$\ -\left(\frac{1}{4\alpha}-\alpha(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\right)\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}.$
Taking expectation, averaging (93) over $k$ from 00 to $K-1$, and choosing the learning rate $\alpha\leq\frac{1}{2\sqrt{1+1/u}(1+W_{\max}/\tau)C_{f}M}$ yield
$\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]$ (94)
$\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{k+1}])}{\alpha K(1-(1+u)W_{\max}^{2}/\tau^{2})}+\frac{2\alpha L_{f}\sigma^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$
$\ +\sigma^{2}\frac{1}{K}\sum_{k=0}^{K-1}\frac{(1+u)\|W_{k}\|^{2}_{\infty}/\tau}{1-(1+u)\|W_{k}\|^{2}_{\infty}/\tau^{2}}+\frac{\alpha^{2}(1+1/u)(1+W_{\max}/\tau)^{4}C_{f}^{2}M^{3}\sigma^{2}}{(1-(1+u)W_{\max}^{2}/\tau^{2})}.$
Using the fact that $V_{k}\geq 0$ for any $k\in{\mathbb{N}}$ and $V_{0}=f(W_{0})-f^{*}$, and choosing the learning rate $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$, we have
$\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]$ (95)
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}^{\prime}+O\left(\frac{1+1/u}{K}\right).$
where $S_{K}^{\prime}$ denotes the amplification factor given by
$S_{K}^{\prime}:=\frac{1}{K}\sum_{k=0}^{K}\frac{(1+u)\|W_{k}\|_{\infty}^{2}/\tau^{2}}{1-(1+u)\|W_{k}\|_{\infty}^{2}/\tau^{2}}\leq\frac{(1+u)W_{\max}^{2}/\tau^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}.$ (96)
Appendix F Convergenc of digital asynchronous pipeline SGD
Theorem 3 (Convergence rate, asynchronous pipeline).
Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$, it holds that
[…]
Proof.
The validation of inequality (78) does not rely on the any training algorithm and hence it still holds here. To bound the last term in the RHS of (78), we have
$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha C_{f}}{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|\right]^{2}$
$\leq$ $\ \frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$ (98)
where the first inequality is according to Lemma 5.
According to the descent lemma in (E), we have
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (99)
where the second inequality is earned by plugging in (F).
Construct a Lyapunov function as
${\mathbb{V}}_{k}:=f(W_{k})+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M-m^{\prime}+2)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$ (100)
According to (99), we have
$\ {\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}$
$\leq$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ +\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k}(k^{\prime}-k+M-m^{\prime}+1){\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M-m^{\prime}+2)\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\leq$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ -\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]+\frac{\alpha C_{f}M^{2}}{2}(M-m^{\prime}+1)\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}].$
$\leq$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\ -\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]+\frac{\alpha C_{f}M^{3}}{2}\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$
$\leq$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ -\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]+\alpha^{3}C_{f}M^{3}\sigma^{2}$
$\ -\left(\frac{1}{4\alpha}-\alpha C_{f}M^{3}\right){\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$\leq$ $\ -\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\alpha^{3}C_{f}M^{3}\sigma^{2}$
where the fourth inequality holds because
$\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$ $=\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}^{(m^{\prime})}-W_{k}^{(m^{\prime})}\|^{2}]$
$={\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
$~{}~{}~{}~{}+2\alpha^{2}{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$
$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]+2\alpha^{2}\sigma^{2}.$ (101)
and the last inequality is earned by letting $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$. Reorganizing the terms yield
$\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}\leq$ $\ {\mathbb{V}}_{k}-{\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]+\alpha^{2}L_{f}\sigma^{2}+\alpha^{3}C_{f}M^{3}\sigma^{2}.$
Therefore, by telescoping and taking the expectation, we get
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{K}])}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}$
$\leq$ $\ \frac{2(f(W_{0})-f^{*})}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}.$
Letting $\alpha\leq\sqrt{\frac{f(W_{0})-f^{*}}{L_{f}\sigma^{2}K}}$, we obtain that
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})L_{f}\sigma^{2}}{K}}+\frac{2C_{f}M^{3}(f(W_{0})-f^{*})}{KL_{f}}.$
The proof is completed. ∎
layer name output size 10-layer 34-layer
conv1 112$\times$112 7$\times$7, 64, stride 2
conv2_x 56$\times$56 3$\times$3 max pool, stride 2
$\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$3
conv3_x 28$\times$28 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$4
conv4_x 14$\times$14 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$6
conv5_x 7$\times$7 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$3
fc 1$\times$1 average pool, fully connected, softmax
block number 6 18
Table 3: ResNet models in the simulations. Downsampling is performed by conv3_1, conv4_1, and conv5_1 with a stride of 2.
Appendix G Simmulation details
This section provides details about the experiments in Section 5. The analog training algorithm, i.e., Analog SGD, is provided by the open-source simulation toolkit AIHWKIT [31], which has Apache-2.0 license; see github.com/IBM/aihwkit. We use Softbound device provided by AIHWKIT to simulate the asymmetric linear device (ALD), by setting its upper and lower bound as $\tau$. The digital algorithm, including SGD, and datasets used in this paper, including CIFAR10 and CIFAR-100, are provided by PyTorch, which has BSD license; see https://github.com/pytorch/pytorch.
The other setting follows the standard settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [33]. A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
Compute Resources. We conduct our simulations on an NVIDIA RTX 3090 GPU, which has 24GB memory. The simulations take from 30 minutes to 7 hours, depending on the size of the model and the dataset.
Statistical Significance. The simulations reported in Figure 2 and 4 are repeated three times. The randomness originates from the data shuffling, random initialization, and random noise in the analog device simulator. The mean and standard deviation are calculated from the final output of each run using NumPy library.
Image classification on CIFAR10/CIFAR100 with Resnet models. In the simulations, we train ResNet models on two benchmark datasets: CIFAR-10 and CIFAR-100. The CIFAR-10 dataset consists of 60,000 32$\times$32 color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds. The CIFAR-100 dataset is similar in format but includes 100 classes.
In Figure 4 and 5, ResNet10 is chosen while in Figure 1, a larger model, ResNet34 is chosen. As Table 3 shows, ResNet10 and ResNet34 are composed of 6 and 18 blocks, respectively. In Figure 4 and 5, each stage contains one block while in Figure 1, 18 blocks are separated into $M=1,2,4,8,12,16$ stages as evenly as possible.
Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default. The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage number $M=12,16$, the initial learning rate is $\alpha=0.01$. The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training. The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs. To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [34].

### v2

@@
-Pipeline Gradient-based Model Training on Analog In-memory Accelerators
+On the Convergence Theory of Pipeline Gradient-based Analog In-memory Training
 Abstract
-Aiming to accelerate the training of large deep neural models in an energy-efficient way, an analog in-memory computing (AIMC) accelerator emerges as a solution with immense potential.
-In AIMC accelerators, trainable weights are kept in memory without the need to move from memory to processors during the training, significantly reducing overhead.
-However, although the in-memory feature enables efficient computation, it also constrains the use of data parallelism since copying weights from one AIMC to another is expensive.
-To enable parallel training using AIMC, we propose synchronous and asynchronous pipeline parallelism for AIMC accelerators inspired by the pipeline in digital domains.
-This paper provides a theoretical convergence guarantee for both synchronous and asynchronous pipelines in terms of both sampling and clock cycle complexity, which is non-trivial since the physical characteristic of AIMC accelerators leads to analog updates that suffer from asymmetric bias.
-The simulations of training deep models on real datasets verify the efficiency of pipeline analog training.
-1 Introduction
+Aiming to accelerate the training of large deep neural networks (DNN) in an energy-efficient way, analog in-memory computing (AIMC) emerges as a solution with immense potential.
+AIMC accelerator keeps model weights in memory without moving them from memory to processors during training, reducing overhead dramatically.
+Despite its efficiency, scaling up AIMC systems presents significant challenges.
+Since weight copying is expensive and inaccurate, data parallelism is less efficient on AIMC accelerators.
+It necessitates the exploration of pipeline parallelism, particularly asynchronous pipeline parallelism, which utilizes all available accelerators during the training process.
+This paper examines the convergence theory of stochastic gradient descent on AIMC hardware with an asynchronous pipeline (Analog-SGD-AP).
+Although there is empirical exploration of AIMC accelerators, the theoretical understanding of how analog hardware imperfections in weight updates affect the training of multi-layer DNN models remains underexplored.
+Furthermore, the asynchronous pipeline parallelism results in stale weights issues, which render the update signals no longer valid gradients.
+To close the gap, this paper investigates the convergence properties of Analog-SGD-AP on multi-layer DNN training.
+We show that the Analog-SGD-AP converges with iteration complexity ${\mathcal{O}}\left(\varepsilon^{-2}+\varepsilon^{-1}\right)$ despite the aforementioned issues, which matches the complexities of digital SGD and Analog SGD with synchronous pipeline, except the non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$.
+It implies that AIMC training benefits from asynchronous pipelining almost for free compared with the synchronous pipeline by overlapping computation.
+I Introduction
+Figure 1: Illustration of MVM computation in AIMC accelerators.
+The weight $W^{(m)}$ on layer $m$ is stored in a crossbar tile.
+The $(i,j)$-th element of $W^{(m)}$ is represented by the conductance of the $(i,j)$-th resistive element.
+MVM operation $z^{(m)}=W^{(m)}x^{(m)}$ is performed by applying voltage $x^{(m)}]_{j}$ between $j$-th and $(j+1)$-th row.
+By Ohm’s law, the current is $I_{ij}=[W^{(m)}]_{ij}[x^{(m)}]_{j}$; and by Kirchhoff’s law, the total current on the $i$-th column is $\sum_{j}I_{ij}=\sum_{j}[W^{(m)}]_{ij}[x^{(m)}]_{j}$.
+Unlike the digital counterpart, no movement of $W^{(m)}$ is required for MVM calculation.
 Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications.
 As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive.
-For example, it requires 184 thousand GPU hours to train an LLAMA2 7 billion model, and this time increases to 1.7 million GPU hours for its 70 billion version [1].
-Stemming from the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators emerges.
-In principle, ideal specialized hardware for DNN should be efficient, support large-scale parallelism, and remain lower energy-economic simultaneously.
-As the demand for efficient processing intensifies, the fundamental backbone of both training and inference largely relies on matrix-vector multiplication (MVM) operation, which is the core component of the ubiquitous fully connected and convolutional layers.
-Typically, each forward and backward pass requires on the order of millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
+For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [1].
+In response to the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators has emerged.
+The ubiquitous fully connected and convolutional layers, whose core operation is matrix-vector multiplication (MVM), form the fundamental backbone of both training and inference as the demand for efficient processing intensifies.
+Typically, each forward and backward pass requires millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
 The promise of analog computing.
-To this end, a promising solution to accelerate the MVM operations is the analog in-memory computating (AIMC) accelerator, wherein both the input and output of MVM operation are analog signals, and the trainable weights of DNN layers are stored in analog crossbar tiles [2].
-Different from traditional Von Neumann accelerators like GPU and TPU, which require weight movement from memory to processor before computation, the AIMC accelerator keeps the weight stationary and performs MVM operations in memory.
-Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Figure 1.
+To this end, a promising solution to accelerate the MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage.
+The DNN weights are stored in analog crossbar tiles consisting of resistive elements [2] where the conductance of the resistive element represents the weight magnitude.
+Unlike traditional Von Neumann accelerators such as GPUs and TPUs, which move weights from memory to the processor during computation, AIMC accelerators keep the weights stationary and perform MVM operations in memory.
+Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Fig.
+1.
 The curse of data parallelism.
-While efficient computation is a benefit, in-memory computing imposes limitations on scalability.
+While efficient computation is beneficial, in-memory computing imposes scalability limitations.
 Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training.
-In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different computation devices to support gradient descent with large batches [4, 5].
-On the contrary, in AIMC architecture, the trainable weights are stored in the analog tiles in the form of device conductances, and thus, copying weights from one analog tile to another is expensive and prone to error.
+In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different accelerators to support gradient descent with large batches [4, 5].
+By contrast, in the AIMC architecture, trainable weights are stored in the analog tiles as the conductances of resistive elements; thus, copying weights is expensive and prone to error.
 Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
-Due to the physical limitations of the AIMC accelerator, we instead resort to pipeline parallelism [6], another prevalent parallel mechanism.
-In pipeline parallelism, a large model is partitioned into a series of smaller stages and gets distributed across multiple devices, with each stage corresponding to the computation of a few layers.
-After a device finishes the process of one datum (or one batch), it passes the intermediate outputs to the next devices and begins the process of the next data.
-In this way, the model that is too large to fit into one device can be separated into different AIMC devices, enlarging the efficient memory.
-Meanwhile, the computation on different devices can happen simultaneously, and the systems can process more data in a given period of time.
-While pipelining is a simple idea, implementing it in analog training faces a series of crucial challenges.
-First, AIMC accelerators usually suffer from noisy gradient signals and asymmetric updates [7, 8], which introduce errors on all forward, backward, and update passes.
-Second, to fully realize the potential of pipeline parallelism, the weight update pass usually involves stale gradients.
-Both of them incur errors in the training, leading to the following question:
-How to design analog pipeline parallel training approaches that do not impair the training?
-This paper answers this question by analyzing the sample complexity of different pipeline algorithms and wall-clock time.
-Before that, we will briefly review prior work that attempted to solve the issues.
-Figure 1: (Left) Illustration of MVM computation in AIMC accelerators.
-The weight $W^{(m)}$ on layer $m$ is stored in a crossbar tile consisting of an array of resistors.
-The $(i,j)$-th element of $W^{(m)}$ is represented by the conductance of the $(i,j)$-th resistor.
-To perform an MVM operation $z^{(m)}=W^{(m)}x^{(m)}$, voltage $[x^{(m)}]_{j}$ is applied between $j$-th and $(j+1)$-th row.
-By Ohm’s law, the current is $I_{ij}=[W^{(m)}]_{ij}[x^{(m)}]_{j}$; and by Kirchhoff’s law, the total current on the $i$-th column is $\sum_{j}I_{ij}=\sum_{j}[W^{(m)}]_{ij}[x^{(m)}]_{j}$.
-Unlike the digital counterpart, no movement of $W^{(m)}$ is required.
-(Right) Speedup of the proposed pipeline training methods as the number of stages increases compared with single machine training case (stage=1); see Section 5 for more details.
+Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [6].
+Pipeline parallelism partitions a large model into a series of stages distributed across multiple accelerators, with each stage corresponding to the computation of a few layers.
+After an accelerator processes one datum (or batch), it passes the intermediate outputs to the next accelerator and begins processing the next datum.
+Pipeline parallelism overlaps computation across different accelerators, increasing system throughput.
+Synchronous and asynchronous parallelism.
+As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [7, 8], Deepspeed [9], FSDP [10], Megetron-LM [11], Colossal-AI [12], just to name a few.
+In their implementation, a gradient mini-batch (e.g., with batch size 128) is split into micro-batches (e.g., 8 micro-batches with batch size 16) and filled into the pipeline.
+The model weights are updated after all the micro-batches are processed.
+Featuring the synchronization operation at the end of each mini-batch, this pipeline paradigm is referred to as synchronous pipeline.
+However, during the fill and drain phases, the pipeline is not fully occupied, which limits the device utilization.
+To overcome the issue, asynchronous pipeline is proposed, which continuously fills the micro-batch into the pipeline and updates the weights once the gradient is available.
+Asynchronous pipeline increases the device utilization, while the weights for gradient computing can be stale, leading to a biased training dynamic.
+To fully realize the promise of analog training, this paper focuses on the theoretical understanding of the viability of stochastic gradient descent on AIMC hardware with asynchronous pipeline (Analog-SGD-AP).
+Besides the stale signal issue in the asynchronous pipeline, the update dynamics on AIMC accelerators also suffer from bias since the weight update involves adjusting conductance, which is difficult to control in theory.
+Both stale signals and imperfect update dynamics lead to the following natural question:
+How do their combined effects provably impact training on the AIMC accelerator?
+Despite two sources of bias, this paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and SGD on AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
+I-A Main results
+This paper aims to understand the impact of bias and investigate the computational complexities and acceleration effect of Analog-SGD-AP.
+We highlight the unique technical challenges compared with the analysis of the asynchronous pipeline on digital accelerators as follows.
+  1. T1)
+    Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias.
+To be specific, consider an analog weight $W_{k}$ on a resistive crossbar array and a desired update $\Delta W_{k}$.
+The updated weight becomes
+    $\textsf{Analog Update}\quad W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}^{(m)}$ (1)
+    where $\tau$ is a hardware-specific constant, $|\,\cdot\,|$ is element-wise absolute value, and $\odot$ is the element-wise production.
+Unlike the accurate digital training, the last bias term in (1) is unique to analog accelerators.
+This bias can not be eliminated within a reasonable cost and leads to challenges to the analysis.
+In section II-B, we provide more background.
+  2. T2)
+    Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [13].
+However, in the asynchronous pipeline, the forward pass is performed using stale weights, while the backward pass uses the latest weights.
+Consequently, their inconsistent multiplication used for weight updates is not a valid gradient, and this error accumulates during training.
+The stale signals in DNN render the update increment used in Analog-SGD-AP no longer a valid gradient, making its convergence behavior unclear.
+  3. T3)
+    Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis.
+Without considering a specific multi-layer architecture, prior work [14] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
+Pipeline Type Sample Complexity Comp. Density Wall-clock Complexity
+Analog-SGD-WOP ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{1}{M}$ ${\mathcal{O}}\left(M\sigma^{2}\varepsilon^{-2}\right)$
+Analog-SGD-SP [Thm. 1] ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{B}{M+B-1}$ ${\mathcal{O}}\left((M+B)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$
+Analog-SGD-AP [Thm. 2] ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$ 11 ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$
+TABLE I: Comparison of different pipeline strategies.
+Sample complexity is the number of samples needed to achieve a training loss $\leq\varepsilon$.
+Computing density is the average portion of active accelerators.
+In the table, $M$ and $B$ are the number of stages and micro-batches, respectively.
+Wall-clock complexity measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by computation density.
+Overcoming the challenges, this paper answers the question above and provides both a convergence guarantee and computational complexities. The key contributions are as follows.
+  1. C1)
+    We consider the training of a multi-layer DNN, which consists of linear layers and non-linear activation functions.
+We formulate the training dynamics of Analog-SGD-AP, which reflects the stale weights in the asynchronous pipeline and the bias from the AIMC accelerators’ imperfection.
+We investigate the smoothness of the training objective as well as the robustness of the network against stale weights.
+  2. C2)
+[…]
+  3. C3)
+    We verify the speedup of Analog-SGD-AP against Analog-SGD-WOP and Analog-SGD-SP in AIMC accelerator simulator, AIHWKIT.
+Numerical simulation results show that despite a slightly larger sample complexity, Analog-SGD-AP achieves the same accuracy with fewer clock cycles due to its high computation density, which is consistent with the theoretical guarantee.
+Furthermore, empirical results show that Analog-SGD-AP achieves linear speedup at least within the range of 1-8 accelerators.
+I-B Prior art.
+To put our work in context, we review prior art that we group into the following two categories.
 Gradient-based training on AIMC accelerators.
-Analog SGD implemented by pulse update is extremely efficient [8].
+Analog training has shown promising early successes in certain tasks [15, 16, 17], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware.
 However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [18].
-To overcome these issues, a family of analog training algorithms, Tiki-Taka (TT) has been proposed.
+Numerous algorithms have been proposed to address these issues.
 TT-v1 [19] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error.
-To demonstrate its efficiency, [10] establishes a model to characterize the dynamic of the analog training and provides a theoretical convergence guarantee.
+To demonstrate its efficiency, [20] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee.
+Building on that, [21] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods.
 To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [22] introduces a digital array.
 Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [23] are proposed to mitigate the impact of inaccurate symmetric point correction [24].
-Despite the above efforts, analog training of large models on huge datasets is still a major challenge.
-Synchronous and asynchronous parallelism.
-Typically, existing pipeline designs can be classified into two categories.
-If the weights are updated using the latest gradients, it is synchronous pipeline; otherwise, it is asynchronous pipeline.
-As the computing requirement of training large models emerges, a plethora of pipeline parallel libraries has been developed, including Gpipe [14, 15], Deepspeed [16], FSDP [17], Megetron-LM [18], Colossal-AI [19], just to name a few.
-Nevertheless, they are synchronous pipeline that enjoys the same dynamic as SGD, so they have the same convergence guarantee as SGD.
-Apart from that, researchers also made some effort to explore the asynchronous pipeline.
-For instance, [20] studied the asynchronous pipeline setting and proposed weight-prediction approaches to compensate for the delay.
-PipeDream [21] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup.
-To mitigate the negative effect of stale gradients, [22] and [23] have proposed only overlapping gradient computation and communication so that delay is always one.
-However, these algorithms are not for analog pipeline training.
+To overcome the limit update granularity issue, [25] proposed a multi-tile training paradigm.
+However, the mentioned works focus on training single AIMC accelerators.
+Instead, this work aims to scale up the system by parallelizing model computation.
+Asynchronous pipeline.
+Existing works proposed methods to alleviate the stale weight issues in the asynchronous pipeline.
+For instance, [26] proposed weight-prediction approaches to compensate for the delay, PipeDream [27] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [28] and [29] propose only overlapping gradient computation and communication to ensure the delay is one.
+These algorithms are not for analog pipeline training.
 Researchers also attempted to design analog pipeline circuits to accelerate the training [30, 31], but they focus on the synchronous pipeline.
-On the theory side, there are a few studies on pipeline training on digital devices such as [26], but these results rely on weight stashing, which is expensive for analog accelerators.
-As [21] points out, delay affects both forward and backward passes in the asynchronous pipeline training.
-It renders the increment no longer a valid gradient without the weight stashing and makes the analysis challenging.
-In the context of prior work, we summarize our key contributions as follows.
-  1. C1)
-    To fully embrace the promise of analog computing, we systematically study the pipeline training on AIMC accelerators, which enables training larger models by distributing models across multiple accelerators and increasing computation density.
-We study the existing work and classify them as synchronous pipelines, which achieve considerable speedup, but their computation density is still sub-optimal.
-To further increase the computation density, we propose the first asynchronous pipeline training algorithm on AIMC accelerators.
-  2. C2)
-    We provide a convergence guarantee for the proposed algorithms.
-Theoretical results show that two strategies have pros and cons: synchronous pipeline has a better complexity while its computation density in the device is lower; asynchronous pipeline reduces device idleness but involves stale gradients.
-The main results are summarized in Table 1.
-  3. C3)
-    We conduct empirical simulation to demonstrate the efficiency of pipeline training.
-The results show that both pipeline algorithms significantly outperform their non-pipeline counterparts.
-Despite a slightly larger sample complexity, the asynchronous pipeline achieves the same accuracy with fewer clock cycles because of its high computation density.
-Furthermore, the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices; see Figure 1.
-Pipeline Type Sample Comp. Comp. Density Clock Cycle
-w.o. Pipeline $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{1}{M}$ $O\left(M\sigma^{2}\varepsilon^{-2}\right)$
-Synchronous [Theorem 1] $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{B}{M+B-1}$ $O\left((M+B)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$
-Asynchronous [Theorem 2] $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$ $1$ $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$
-Table 1: Comparision of different pipeline strategies.
-Sample complexity (Sample Comp.) is the number of samples needed for achieving training loss $\leq\varepsilon$.
-Computation density (Comp.
-Density) is the time that is the portion of time when the system is computing the gradients.
-In the table, $M$ is the pipeline depth, and $B$ is the micro-batch size.
-Clock Cycle measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by clock cycle.
-2 Model Training on AIMC Accelerators
-In this section, we first introduce the generic gradient-based training algorithm, encompassing digital algorithms running on devices like GPU and TPU, and algorithms running on AIMC accelerators.
-Consider a standard model training problem on a data distribution ${\mathcal{D}}$, given by
-$W^{*}:=\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}f(W)~{}~{}~{}{\rm with}~{}~{}~{}f(W):={\mathbb{E}}_{\xi\sim{\mathcal{D}}}[f(W;\xi)]$ (1)
-where $W$ is all trainable weights of the model vectorized into a $D$ dimensional vector.
+On the theory side, there are few studies on pipeline training for digital accelerators such as [14, 32], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms.
+As [27] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
+II Analog Training with Model Parallelism
+In this section, we first introduce the multi-layer DNN training model, and then introduce the analog stochastic gradient descent algorithm with the asynchronous pipeline.
+II-A Training problem and back-propagation
+Consider a standard model training problem on a data distribution $\mathcal{P}$} given by
+$\min_{W\in{\mathbb{R}}^{D}}\Big\{f(W):={\mathbb{E}}_{\xi\sim{\mathcal{P}}}[f(W;\xi)]\Big\}$ (2)
+where $W$ is the trainable model weights on analog crossbar tiles, which is vectorized into a $D$-dimension vector.
 DNN model.
 Consider the multi-layer DNN, which consists only of linear layers and activation functions.
-Let $[M]:=\{1,2,\cdots,M\}$ denote a positive integer set.
+Let $M]:=\{1,2,\cdots,M\}$ denote an index set.
 The trainable weights are separated into $M$ different blocks, i.e., $W=[W^{(1)},W^{(2)},\cdots,W^{(M)}]$, where $W^{(m)},\forall m\in[M]$ is a matrix.
 The training of DNN typically consists of three phases: forward, backward, and update.
-Suppose $x$ and $y$ are the feature and label of data $\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$, where ${\mathcal{X}}$ and ${\mathcal{Y}}$ are feature and label space, respectively.
-The forward pass of the model for layer $m\in[M]$ can be expressed by
-$\textsf{Forward}\qquad x^{(1)}=x,~{}~{}~{}~{}$ $z^{(m)}=W^{(m)}x^{(m)},~{}~{}~{}~{}x^{(m+1)}=g^{(m)}(z^{(m)})$ (2)
-where $g^{(m)}$ is the activation function of layer $m$, and $x^{(M)}$ is the feature from the last layer, which is fed together with label $y$ into the loss function $\ell(\cdot,\cdot):{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$ and yields $f(W;\xi)=\ell(x^{(M+1)},y)$.
-For the backward pass, define the backward error $\delta^{(m)}:=\frac{\partial f(W;\xi)}{\partial z^{(m)}}$, wherein the error of last layer is first computed by $\delta^{(M)}=\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})$.
-Subsequently, each layer $m\in[M-1]$ computes the next recursion according to
+Suppose $x$ and $y$ are the feature and label of data $\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$, where $\mathcal{X}$} and $\mathcal{Y}$} are feature and label spaces, respectively.
+The forward pass of the model for layer $m\in[M]$ is expressed by
+Forward $x^{(1)}=x,~~~~z^{(m)}=W^{(m)}x^{(m)},~~~~x^{(m+1)}=g^{(m)}(z^{(m)})$ (3)
+where $g^{(m)}(\,\cdot\,)$ is the activation function of layer $m$, and $x^{(m)}$ is the input feature.
+At the last layer, the ouput $x^{(M)}$ and label $y$ are fed together into the loss function $\ell(\,\cdot\,,\,\cdot\,):{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$ and yields $f(W;\xi)=\ell(x^{(M+1)},y)$.
+For the backward pass, define the backward signal $\delta^{(m)}:=\frac{\partial f(W;\xi)}{\partial z^{(m)}}$, wherein the error of the last layer is first computed by $\delta^{(M)}=\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})$.
+Subsequently, each layer $m\in[M-1]$ computes the next recursion by the chain rule
 $\textsf{Backward}\qquad\delta^{(m)}=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}=(\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$ (4)
 In the forward pass, the activation $z^{(m)}$ is stashed until the backward signal $\delta^{(m+1)}$ returns.
-Typically, the computation $\nabla g^{(m)}(z)$ is light-weight, and hence the major overhead of the backward pass lies at the MVM among $W$ and $\delta$.
+Typically, the computation $\nabla g^{(m)}(z)$ is light-weight, and hence the major overhead of the backward pass lies at the MVM among $W^{(m)}$ and $\delta^{(m)}$.
 With the forward and backward signals, the gradient with respect to the weight $W^{(m)}$ of layer $m$ is computed by
 $\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}=\delta^{(m)}\otimes x^{(m)}$ (5)
-where $\otimes$ is the outer product between two vectors. Having the gradients of each layer, we can use different gradient-based algorithms to update the weights.
-Gradient-based training on AIMC accelerators.
-After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights.
-In digital devices, one of the most popular training algorithms is SGD (SGD), wherein the matrix $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$ is computed and added to $W^{(m)}_{k}$, written by $W_{k+1}^{(m)}=W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})$; $\alpha>0$ is a positive learning rate.
-In the paper, the subscript $k$ is used to indicate the index of gradient update.
-Different from digital devices, AIMC accelerator uses rank-update to update the gradient update, which constructs a series of pulses and sends them to its analog array [8].
-Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, AIMC accelerators perform the update operation in a fast and energy-efficient way.
-Nevertheless, rank-update relies on the special structure that the gradient of $W^{(m)}_{k}$ is a rank-one matrix, which brings us further challenges.
-On top of that, as existing work pointed out [9, 10], the update on AIMC accelerators suffers from asymmetric bias, and thus the dynamic of Analog SGD in single-machine training can be written as
-$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})-\frac{\alpha}{\tau}|\nabla_{W^{(m)}}f(W_{k};\xi_{k})|\odot W_{k}^{(m)}$ (5)
-where $\tau$ is a device-dependent constant. When $\tau\rightarrow\infty$, Analog SGD reduces to Digital SGD.
-The main challenge coming from the asymmetric is the saturation phenomenon [9]. Consider the update of $(i,j)$-th element of matrix $W_{k}^{(m)}$, whose update is written as
-$([W_{k+1}^{(m)}]_{ij}-[W_{k}^{(m)}]_{ij})^{2}=$ $\ \alpha^{2}(\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})-[W^{(m)}_{k}]_{ij}/\tau)^{2}(\nabla_{W^{(m)}}f(W_{k};\xi_{k}))^{2}$
-$\approx$ 0if [$\ 0~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if }[W^{(m)}_{k}]_{ij}\approx\tau~{}\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})$ (6)
-which means that the increment is suppressed when $[W^{(m)}_{k}]_{ij}$ gets close to the $\tau$.
-Therefore, we usually introduce a proper scaling coefficient so that the saturation never happens during the training.
-Figure 2: Illustration of pipelines with 4 devices ($M=4$).
+where $\otimes$ is the outer product between two vectors.
+Rank-update Unlike digital accelerators, AIMC accelerators use rank-update to change the matrix, which leverages the outer product structure as shown in (5).
+To do so, independent random voltage pulses with probability proportional to the elements of $\delta^{(m)}$ and $x^{(m)}$ are applied to columns and rows, respectively.
+Consequently, the $(i,j)$-th element $W]_{i,j}$ receives a pulse coincident with probability $\propto[\delta^{(m)}]_{i}[x^{(m)}]_{j}$.
+Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, rank-update has only $O(BN)$ complexity while the update on digital accelerators has $O(BN^{2})$ complexity to update an $N\times N$ matrix.
+Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [15].
+Figure 2: Illustration of pipelines with 4 accelerators ($M=4$).
 Each mini-batch is split into $B=4$ micro-batches.
 Each color represents one micro-batch, and each row from bottom to top represents stages 0 to 4.
-Each column corresponds to a clock cycle in which one micro-batch is processed.
-The white square indicates the idle device.
-(Top) Vanilla model parallelism without pipeline.
-All weights are updated using gradient-based algorithms after the computation of one mini-batch is completed, which is not presented in the figure.
-(Middle) Synchronous pipeline with micro-batch count $B=5$.
-The update happens after all 5 micro-batches are processed.
-(Bottom) Asynchronous pipeline.
-The update happens once the gradient of one micro-batch is achieved.
-The grey square indicates the corresponding device is processing data that is not fully reflected in the figure.
-3 Analog Pipeline Model Training
-This section introduces pipeline parallelism and shows how it increases computational efficiency.
-Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one device and motivates us to adopt model parallelism.
-In model parallelism, the computation is partitioned into a sequence of stages, which, in general, consists of a consecutive set of layers in the model.
-Each stage is mapped on one digital or analog device.
-For simplicity, we assume each stage contains one layer only.
+Each column corresponds to a single clock cycle in which one micro-batch is processed.
+The white square indicates the corresponding accelerator is idle.
+(Top) Analog-SGD-WOP.
+All weights are updated after each mini-batch is computed, which is not presented in the figure.
+(Middle) Analog-SGD-SP with micro-batch count $B=5$.
+The update occurs after all 5 micro-batches are processed.
+(Bottom) Analog-SGD-AP.
+The update occurs once the gradient of a single micro-batch is achieved.
+The grey square indicates that the corresponding accelerator is processing data that is not fully reflected in the figure.
+II-B Non-ideal update on AIMC accelerator
+Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [19, 20, 21], the update on AIMC accelerators exhibits asymmetric bias.
+To be specific, for any desired update $\Delta W_{k}^{(m)}$, the update on $W_{k}$ is scaled by response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$.
+Considering the weight $W_{k}$ at iteration $k$ and the expected update $\Delta W\in{\mathbb{R}}^{D}$, we express the asymmetric update as $W_{k+1}=U(W_{k},\Delta W)$ with the update function $U:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ defined by
+$U(w,\Delta w):=\begin{cases}w+\Delta w\cdot q_{+}(w),~~~\Delta w\geq 0,\\ w+\Delta w\cdot q_{-}(w),~~~\Delta w<0.\end{cases}$ (6)
+Defining the symmetric and asymmetric components as $F(w):=\frac{1}{2}(q_{-}(w)+q_{+}(w))$ and $G(w):=\frac{1}{2}(q_{-}(w)-q_{+}(w))$, we write the update in (6) as $U(w,\Delta w)=w+\Delta w\cdot F(w)-|\Delta w|\cdot G(w)$.
+This paper consider a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$ as suggested by [20], we reach the dynamics (1).
+Figure 3: Illustration of the dynamics of the asynchronous pipeline.
+The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each accelerator.
+II-C Synchronous and asynchronous pipelines
+Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one accelerator and necessitates model parallelism.
+In model parallelism, the forward and backward computations are partitioned into a sequence of stages, which, in general, consist of a consecutive set of layers in the model.
+Each stage is mapped to one digital or analog accelerator.
 Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$.
-For simplicity, let the micro-batch size $B_{\text{micro}}=1$ in Sections 3-4 so that the subscript $b$ in $\xi_{k,b}$ can represent the index of micro-batch or data, and $B=B_{\text{mini}}$ can represent the number of micro-batch and mini-batch size $B_{\text{mini}}$.
-It could be easily extended to general cases.
-Clock cycle.
-We define the maximum time for each stage to process one micro-batch as the clock cycle.
-Ideally, all devices would process a micro-batch in the same amount of time, avoiding idle periods during a clock cycle.
-Although achieving this is challenging, various task scheduling approaches have been proposed to balance computation [21, 14].
-To explore the potential of pipeline training, this paper uses the clock cycle and stage as the minimal time and basic scheduling units, respectively.
-Based on scheduling methods, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining.
-Model parallelism without pipeline.
-Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data, as previously mentioned) for iteration $k$.
-Vanilla model parallelism processes all micro-batches one by one.
-In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$, while stage $m$ achieves the backward error $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage in the backward pass.
-The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stored when the gradient for the $b$-th micro-batch is completed.
+Without loss of generality, this paper assumes that one stage contains only one layer and the micro-batch size $B_{\text{micro}}=1$.
+Under this assumption, $\xi_{k,b}$ denotes the $b$-th micro-batch which contains only the $b$-th data.
+Consequently, the number of micro-batches is the mini-batch size, i.e., $B=B_{\text{mini}}$.
+This paper treats a stage as the minimal scheduling unit and investigates the time complexity across different scheduling strategies.
+To compare time costs, we define the basic time unit, clock cycle, as the minimum time required for all devices to process one micro-batch.
+Based on scheduling strategies, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining.
+Vanilla model parallelism without pipeline.
+Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data sample, as previously mentioned) for iteration $k$.
+Vanilla model parallelism processes micro-batches sequentially.
+In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$ according to (3), while in the backward pass stage $m$ achieves the backward signal $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage according to (4).
+The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stashed when the gradient for the $b$-th micro-batch is completed.
 After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method.
-Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}$, which iterates over $b\in[B_{\text{mini}}]$ by
+Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}^{(m)}$, which iterates over $b\in[B_{\text{mini}}]$ by
 $W_{k,b+1}^{(m)}=$ $\ W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})-\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (7)
 Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$.
-However, for the AIMC accelerators, only rank-update is permitted so that each gradient is accumulated to the weight $W^{(m)}_{k}$ one by one, rendering the analysis more difficult.
+To fully release the potential of rank-update, AIMC accelerators follow (7) to update its weight, which in turn complicates the convergence analysis.
 Synchronous pipeline.
-Existing work has already attempted to design architecture that supports the pipeline training [24, 25].
-They faithfully implement the iteration (7) and overlap parts of the computation.
-Since they compute all gradients in a batch on the same $W_{k}$, their algorithms can be called synchronous pipeline.
-In the vanilla non-pipeline version, the stage $m$ is idle after its forward computation of $b$-th micro-batch until the backward error comes back from stage $m+1$.
-In the synchronous pipeline training, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any; see the middle of Figure 2.
-Similar to the vanilla non-pipeline version, the weights will be updated after processing all micro-batches.
-To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch (fill and drain periods), hence rendering some devices idle.
+Vanilla model parallelism is simple to implement.
+However, most devices remain idle, reducing system throughput.
+Existing work has already attempted to design an architecture that supports the pipeline training [30, 31].
+They faithfully implement the iteration (7) but overlap the computing inside a mini-batch.
+Featuring synchronization at the end of a mini-batch, this stage schedule mechanism is called synchronous pipeline.
+In the vanilla non-pipeline version, stage $m$ remains idle after completing the forward computation of the $b$-th micro-batch until the backward signal returns from stage $m+1$.
+In the synchronous pipelining, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any.
+Similar to the vanilla model parallelism, the weights will be updated after processing all micro-batches.
+Consequently, for each mini-batch, most of the devices need to wait until the last micro-batch is processed.
+After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (7).
+To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch.
+During the pipeline filling and draining periods, parts of the devices are idle; see the middle of Fig.
+2
 Asynchronous pipeline.
-To increase the usage of all devices, a more progressive strategy, asynchronous pipeline, is proposed.
-Instead of waiting for the gradient of one mini-batch, the asynchronous pipeline training method performs gradient-based update algorithms immediately after one micro-batch is completed.
-Except for short periods of fill and drain at the beginning and the end, all devices are active during the training, maximizing the training efficiency; see the bottom of Figure 2.
+To fully utilize all devices, a more progressive strategy, asynchronous pipeline, is proposed.
+Instead of using mini-batch descent, asynchronous pipeline performs gradient descent on each micro-batch.
+Let $\xi_{k}$ denote the micro-batch for iteration $k$.
+In the forward pass, stage $m$ receives the output of the previous stage $x^{(m)}_{k}$ as input and outputs $x^{(m+1)}_{k}$, while in the backward pass, stage $m$ receives the backward signal $\delta^{(m)}_{k}$ from next stage and sends the error $\delta^{(m-1)}_{k}$ to its previous stage.
+Unlike the synchronous pipeline, the asynchronous pipeline updates the weights once $x^{(m)}_{k}$ and $\delta^{(m)}_{k}$ are ready.
+Without regular synchronization and filling/draining period, no device is idle during the training, maximizing the utilization.
+See Fig.
+2 for the comparison of different parallel computing strategies.
 Despite its perfect resource usage, asynchronous pipeline parallelism suffers from the stale signal issue.
 Notice that in the asynchronous pipeline setting, the forward signal $x^{(m)}$ and backward signals $\delta^{(m)}$ are computed based on different stale weights.
-For example, after outputting forward signal $x^{(0)}$ for the $k$-th micro-batch, the first stage ($m=1$) receives the backward signal $\delta^{(0)}$ after $2(M-1)$ clock cycles, wherein $M-1$ gradient updates happen.
-Consequently, the gradient is stale for the current weights.
-Furthermore, notice both forward (2) and backward (3) passes involve the matrix $W^{(m)}$ while the $W^{(m)}$ is changed between the forward and backward passes.
-To depict the process more precisely, we introduce different symbols, which have tildes on top of them to indicate the staleness.
-Since the mini-batch is no longer needed in asynchronous pipeline, the subscript $b$ for the micro-batch index can be omitted, and we denote the micro-batch for $k$-th update as $\xi_{k}$.
-To make the relationship clearer, the computation dependency is plotted in Figure 3.
-Since there are $M-m$ gradient updates between the forward and backward passes in stage $m$, the weight used for the forward pass has a delay of $M-m$, that is
-${\tilde{x}}^{(m+1)}_{k}=$ $\ g^{(m)}({\tilde{z}}^{(m)}_{k}),\quad{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},\quad{\tilde{x}}^{(0)}_{k}=x_{k}.$ (8)
-In contrast, the weights are updated following the backward pass, which means the weight used for backward is the latest one
-$\tilde{\delta}^{(m)}_{k}=$ $\ \tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k}),$ (9)
-$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
-In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is no longer a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partial delay gradient, making the analysis more challenging.
-Remark 1 (Digital v.s. analog asynchronous pipeline).
-In digital hardware, one can stash the stale weight $W^{(m)}_{k-(M-m)}$ and read it out again to compute the backward signal in (9) [26, 21].
-Weight stashing ensures the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is the gradient with respect to $W^{(m)}_{k-(M-m)}$, and hence provides a better convergence guarantee.
-However, copying is expensive in the analog domain, as we discussed before.
-Consequently, the asynchronous pipeline suffers from partial delay gradient issues.
-Computation density.
-Different pipeline strategies can process varying amounts of data in the same amount of time.
-In an ideal scenario, a micro-batch requires $2M$ clock cycles for forward and backward passes, distributed equally across $M$ devices, taking 2 clock cycles per micro-batch.
-To measure how closely an algorithm approaches this ideal, we introduce computation density.
-Let $N_{\text{c}}(t)$ be the number of data points processed in $t$ clock cycles.
-Assuming infinite runtime, we define computation density as the limit ratio of ideal processing time to real time: $\lim_{t\to\infty}\frac{2N_{\text{c}}(t)}{t}$, where we introduce the factor $2$ since each data involves forward and backward passes.
-For vanilla model parallelism without pipelining, it takes $2M$ cycles per micro-batch, yielding a computation density of $1/M$.
-In synchronous pipelining, it takes $2(M+B-1)$ clock cycles to process $B$ micro-batches, resulting in a density of $B/(M+B-1)$.
-Asynchronous pipelining achieves a density of $1$, as all devices remain active during training.
-Figure 3: Illustration of the dynamic of asynchronous pipeline.
-The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each device.
-4 Convergence and Wall-clock Analysis
-This section provides the convergence guarantee of the proposed pipeline training algorithm on AIMC accelerators.
-We show that both synchronous and asynchronous pipelines enjoy the same convergence rate as the vanilla non-pipeline version.
-By increasing the computation density significantly, the pipeline parallelism provides notable speedup in terms of clock cycle complexity.
-4.1 Assumptions on models, data, and AIMC accelerators
-[…]
+For example, after outputting forward signal $x^{(0)}$ for the $k$-th micro-batch, the first stage ($m=1$) receives the backward signal $\delta^{(0)}$ after 2(M1)2(M-1) clock cycles, wherein M1M-1 gradient updates happen.
+We introduce different symbols for precise description, which have tildes on top of them to indicate the staleness.
+Since there are MmM-m gradient updates between the forward and backward passes in stage $m$, the weight used for the forward pass has a delay of MmM-m.
+In contrast, the weights are updated by the backward pass, which means the weights used for the backward pass are the latest ones.
+The forward and backward passes of asynchronous pipeline training can be summarized as
+Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$.
+(9)
+Imperfect gradient-based training on AIMC accelerators.
+After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights.
+However, there is a bias due to the hardware imperfection [20].
+Replacing $\Delta W_{k}$ with $\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ in (1), where $\alpha>0$ is a positive learning rate, we reach the update dynamics of weights
+$\textsf{Update}\quad W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
+In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
+The main challenge coming from the AIMC training is the weight drift [20].
+Consider the case where $W_{k}^{(m)}$ is already a non-zero critical point of $f(W)$ but not the one of $f(W;\xi)$, i.e.
+${\mathbb{E}}_{\xi\sim{\mathcal{P}}}[\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)]=0$ but $\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)\neq 0$ for some $\xi$.
+After one iteration, the weight $W_{k}^{(m)}$ drifts because of the sample randomness
+${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}\neq 0$ (11)
+which prevents the weights from converging to the critical point, leading to an error.
+III Convergence & Wall-clock Analysis
+This section shows that despite stale gradient issue and the update bias in AIMC accelerators, the asynchronous pipeline converges with the same sample complexity as the counterpart without pipeline or with synchronous pipeline.
+By increasing the computation density, asynchronous pipeline parallelism provides notable speedup in terms of wall-clock complexity.
+III-A Assumptions
+Before showing the complexity of the pipeline algorithms, we first introduce a series of assumptions on the properties of DNN structures, data distribution, as well as the AIMC accelerators.
+$\|\,\cdot\,\|$ is $\ell_{2}$-norm for vectors and Frobenious norm for matrices; $\|\,\cdot\,\|_{\infty}$ takes the element with maximal absolute value of vector or matrix.
 Assumption 1 (Unbiasness and bounded variance).
-The sample $\xi_{k}$ are independently sampled from a distribution ${\mathcal{D}}$ over times $k\in[K]$, and has bounded variance, i.e., ${\mathbb{E}}_{\xi_{k}}[\nabla f(W_{k};\xi_{k})]=\nabla f(W_{k})$ and ${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}$.
+The stochastic gradient is unbiased and has bounded variance $\sigma^{2}$.
 Assumption 2 (Normalized feature).
@@
 $g(0)=0$.
-The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid and Tanh functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
+The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
 Assumption 4 (Loss function).
@@
 $\ell(y_{1},y_{2})\geq f^{*},\forall y_{1},y_{2}$.
-Note that the lower bounded loss implies the lower boundness of the objectives as well, i.e.
-$f(W;\xi)\geq f^{*}$, $\forall W\in{\mathbb{R}}^{D},x\in{\mathcal{X}}$.
-Aside from the assumptions, to avoid saturation during training, we also need proper scaling so that the saturation degree of the AIMC accelerators is bounded.
+Aside from the assumptions, to avoid saturation during training, following the setting in [20], the saturation degree is assumed to be bounded.
 Assumption 5 (Bounded Saturation).
-There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$ norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max}<\tau,\forall k\in[K]$.
-The ratio $W_{\max}/\tau$ is referred to as saturation degree.
-Assumption 5 requires the AIMC accelerators to have a sufficiently large active region to represent weights.
-Besides, on digital devices, the trainable weights always remain bounded in a small neighborhood of the initialization points during training [27, 28, 29], especially for overparameterized networks.
-Given $W_{\max}$ is sufficiently large to include this neighborhood, Assumption 5 holds.
-While it is possible to prove this assumption, we adopt it to make our results more concise and to focus on revealing the key insights on the effect of pipeline training.
-Under the above assumptions, we prove the objective $f(W;\xi)$ is smooth.
-Lemma 1 (smoothness of the objective).
-Under Assumption 25, the objective is $L_{f}$-smooth with respect to $W$, i.e. $\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$.
-The proof of Lemma 1 is deferred to Appendix C.
-Lemma 1 enables us to adapt the tools from [30] for our analysis.
-However, due to the asymmetric bias of the AIMC accelerators, the convergence results for Analog SGD are still non-trivial.
-4.2 Convergence rate of pipeline training algorithms
-In this section, we provide the convergence rate of Analog SGD with pipelines.
-In non-convex optimization tasks like the DNN training, instead of measuring the objective (1), it is more common to measure the gradient square norm $\|\nabla f(W_{k})\|^{2}$.
-Theorem 1 (Convergence rate, synchronous pipeline).
+There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$.
+The ratio $W_{\max,\infty}/\tau$ is termed saturation degree.
+From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range [20].
+It is possible since the trainable weights always remain bounded in a small neighborhood of the initialization points during training [33, 34, 35], especially for overparameterized networks in digital training.
+Given $W_{\max,\infty}$ is sufficiently large to include this neighborhood, Assumption 5 holds.
+While it is possible to prove this assumption as in [36], we adopt it to focus on revealing key insights into the effect of pipeline training.
+III-B Iteration complexity of pipeline training
+Next, we present the convergence rate of analog SGD with pipelines.
+In non-convex optimization tasks like the DNN training, it is standard to use the stationary measure as $E_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]$.
+As a warm-up, we investigate the convergence of the synchronous pipeline, Analog-SGD-SP.
+Theorem 1 (Iteration complexity, Analog-SGD-SP).
 Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{B(f(W_{0})-f^{*})}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\frac{1}{L}$, it holds that
-$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S$
-where $S$ denotes the amplification factor given by $S:=\frac{W_{\max}/\tau^{2}}{1-W_{\max}/\tau^{2}}$.
+$E_{K}\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S$
+where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}/\tau^{2}}{1-W_{\max,\infty}/\tau^{2}}$.
 The proof of Theorem 1 is deferred to Appendix D.
-Similar to the convergence of digital SGD [30], Analog SGD converges in the rate $O(\sqrt{{\sigma^{2}}/({BK})})$.
-However, because of the asymmetric bias, the convergence rate slows down by a factor $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$, which reduces to $1$ when $\tau$ tends to infinity.
-Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error.
-When the AIMC accelerators can be well-designed, $\tau$ can be sufficiently large, and it does not dominate the error when $K$.
-Empirically, researchers also proposed some Analog SGD variants like Tiki-Taka [9, 11] to eliminate the negative impact of asymmetric bias.
-Next, we provide the iteration complexity of Analog SGD with asynchronous pipeline training.
-Theorem 2 (Convergence rate, asynchronous pipeline).
+Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [37].
+Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [20], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [20, 19, 22].
+Next, we provide the iteration complexity of Analog-SGD-AP.
+Theorem 2 (Iteration complexity, Analog-SGD-AP).
 Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\frac{1}{L_{f}}$, it holds that
-[…]
+$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ (12)
 where $u>0$ is any number and $S^{\prime}$ denotes the amplification factor given by $S^{\prime}:=\frac{(1+u)W_{\max}^{2}/\tau^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$.
 The proof of Theorem 2 is deferred to Appendix E.
-Because of the stale signals, the convergence rate of the asynchronous pipeline slows down.
-The coefficient of the dominant term degrades from $\frac{1}{1-W_{\max}/\tau^{2}}$ to $\frac{1}{1-(1+u)W_{\max}/\tau^{2}}$, and an extra error term $O\left(\frac{1+1/u}{K}\right)$ is introduced.
-Fortunately, $O\left(\frac{1+1/u}{K}\right)$ is not the dominant term and we can specify $u$ as a small constant so the impact is minor.
+Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which vanishes for a well-designed AIMC accelerator or can be mitigated by error-correction algorithms [20, 19].
+Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence rate as the Analog-SGD-WOP [20].
+Remark 1.
+As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Theorem 3 in the Appendix F.
+Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [32].
+III-C Proof sketch of Theorem 2
+In this subsection, we highlight the key steps of the proof towards Theorem 2.
+(Step I) Smoothness of the objective.
+Before analyzing the training dynamic, we show that the objective is smooth.
+Recall that $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$, we first show that under the assumptions of DNN structures, both $\delta^{(m)}$ and $x^{(m)}$ are smooth.
+Using slightly abusing notations, we denote the forward and backward signals computed by $W$ as $x^{(m)}$ and $\delta^{(m)}$, respectively; and denote the ones computed by $\tilde{W}$} as ${\tilde{x}}^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
+The following lemma show that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
+Lemma 1 (Signal stability).
+Denote the forward signals computed by $W$ and $\tilde{W}$} as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
+$\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}\leq$ $\ C_{x^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (13)
+$\|{\tilde{z}}^{(m)}-z^{(m)}\|^{2}\leq$ $\ C_{z^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (14)
+$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq$ $\ C_{\delta^{\prime}}^{(m)}\|{\tilde{W}}-W\|^{2},$ (15)
+where $C_{x^{\prime}}^{(m)}$, $C_{z^{\prime}}^{(m)}$, and $C_{\delta^{\prime}}^{(m)}$ are constants that depend on the DNN architecture.
+Based on this lemma, we prove the objective $f(W;\xi)$ is smooth.
+Lemma 2 (Smoothness of the objective).
+Under Assumption 25, the objective is $L_{f}$-smooth with respect to $W$, i.e.
+$\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$, where $L_{f}$ is a constant depending on the DNN archetecture.
+Lemma 2 enables us to adapt the tools from [37] to obtain the sufficient descent of the function value.
+However, the bias from the asymmetric update and stale signals from the asynchronous pipelining accumulates layer by layer within the DNN and eventually lead to an extra term in the upper bound.
+Next step we quantify the error of the stale signals.
+(Step II) Error of stale signals.
+Similar to Step I, we begin from analyzing the error of $x^{(m)}$ and $\delta^{(m)}$.
+However, the stale signal involves the weights at different iterations and introduces additional errors.
+Intuitively, the error of forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal.
+It motivated us to bound the error by the delay.
+Since the error between latest $W^{(m^{\prime})}_{k}$ and the stale weight $W^{(m^{\prime})}_{k-(M-m)}$ can be decomposed as
+$W^{(m^{\prime})}_{k}-W^{(m^{\prime})}_{k-(M-m)}=\!\!\!\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\!\!\!(W^{(m^{\prime})}_{k^{\prime}+1}-W^{(m^{\prime})}_{k^{\prime}})$, (16)
+the following accumulated delay of layer $m$ can be used for analyzing the upper bound of the stale signals’ error
+$\psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (17)
+where $\Delta^{(m)}_{k}:=W^{(m)}_{k+1}-W^{(m)}_{k}$ is the difference between consecutive weights at iteration $k$.
+Lemma 3 (Delay error of signals).
+Under Assumption 24, the difference between the latest and stale signals generated by Analog-SGD-AP are bounded, i.e., $\forall m\in[M+1]$
+$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\psi_{k},$ (18)
+$\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\psi_{k},$ (19)
+$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\psi_{k},$ (20)
+where $C_{x}^{(m)}$, $C_{z}^{(m)}$, and $C_{\delta}^{(m)}$ are constants that depends on the DNN architecture.
+Remark 2.
+The inequalities in Lemma 1 and 3 have a similar form, but they provide upper bounds for different error: in Lemma 1, two signals are computed on two arbitrary weights $W$ and $\tilde{W}$}, while in Lemma 3, two signals are generated by Analog-SGD-AP.
+(Step III) Sufficient descent of the objective.
+Building on the results in Steps I and II, we will analyze the sufficient descent of the objective function.
+However, due to the asymmetric bias of the AIMC accelerators, the convergence results for analog SGD are still non-trivial.
+Lemma 4 (Sufficient descent).
+Under Assumption 15, the objective function $f(W_{k})$ satisfies
+$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$ (21)
+$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)+C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
+where $\lesssim$ performs non-essential approximation on the upper bound of descent.
+From Lemma 4, we notice that the accumulated delays $\psi_{k}$ appear in the descent lemma, but intuitively they vanish when the algorithm converges according to its definition.
+(Step IV) Construction of the Lyapunov function. To compensate for this term in equation (21), we design the following Lyapunov function.
+${\mathbb{V}}_{k}:=f(W_{k})-f^{*}+C_{\psi}\Psi_{k}$ (22)
+where the auxiliary function $\Psi_{k}$ is defined as
+$\Psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (23)
+and $C_{\psi}$ is a constant. To achieve the dynamics of $\Psi_{k}$, we derive that
+$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (24)
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+Let the index of the inner summation of the first term in the RHS of (24) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
+$\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (25)
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}+M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
+$=$ $\ \Psi_{k}+M\|W_{k+1}-W_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
+where the last equality holds by definition of $\Delta^{(m^{\prime})}_{k}$, i.e.,
+$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=$ $\ \sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}.$ (26)
+Plugging (25) into (24) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$.
+We then show that the Lyapunov functions have sufficient descent if $C_{\psi}$ is properly chosen, i.e.,
+$\mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (27)
+Reorganizing the inequality above yields Theorem 2.
+Figure 4: Training ResNet models on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
+(Left) Accuracy-vs-Epoch reflects sample complexity.
+(Middle) Accuracy-vs-Clock cycle.
+(Right) Speedup of the proposed pipeline training methods on AIMC accelerators as the number of stages increases compared with the single machine training case (stage=1).
+We ignore the communication latency since the asynchronous pipeline does not introduce obvious extra latency.
+III-D Sampling and wall-clock complexity
+Sample complexity measures the computational cost required to achieve the desired accuracy, whereas in model parallelism, the computation of different stages overlaps.
+Therefore, we are also interested in the wall-clock time it takes.
+In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\epsilon$.
+Due to asymptotic error, we define the sample complexity in analog training as the minimum micro-batch needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e.
+$\sigma^{2}S$ for Analog-SGD-WOP/Analog-SGD-SP and $\sigma^{2}S^{\prime}$ for Analog-SGD-AP, respectively.
+Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, we have the following conclusion.
+Corollary 1 (Sample complexity).
+The sample complexities of Analog-SGD-WOP or Analog-SGD-SP are ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$, while that of the Analog-SGD-AP is ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$.
+This shows that three model parallelisms share the same dominant term ${\mathcal{O}}(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while Analog-SGD-AP shows down by a non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$.
+Since the computing is overlapped, the computation density is also needed to yield the number of clock cycles needed to achieve the $\varepsilon$ accuracy.
+Computation density.
+In an ideal scenario, a micro-batch requires $2M$ clock cycles for forward and backward passes, distributed equally across $M$ accelerators, taking 2 clock cycles per micro-batch.
+Computation density measures how closely an algorithm approaches this ideal.
+Let $N_{\text{c}}(t)$ be the number of micro-batches processed in $t$ clock cycles.
+We define computation density as the limit ratio of ideal processing time to real time: $\lim_{t\to\infty}\frac{2N_{\text{c}}(t)}{t}$, where the factor $2$ indicates that micro-batch processing involves forward and backward passes.
+Analog-SGD-WOP takes $2M$ cycles per micro-batch, yielding computation density $1/M$.
+In Analog-SGD-SP, it takes 2(M+B1)2(M+B-1) clock cycles to process $B$ micro-batches, resulting in a density of B/(M+B1)B/(M+B-1).
+Analog-SGD-AP achieves a density of $1$, as all accelerators remain active during training.
+Similar to sample complexity, wall-clock complexity is defined as the minimum clock cycle needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error.
+Wall-clock complexity is achieved by dividing the sample complexity by the computation density.
+The following corollary summarizes the clock cycle complexity of different pipeline strategies.
+Corollary 2 (Wall-clock complexity).
+The wall-clock complexities of Analog-SGD-WOP, Analog-SGD-SP, and Analog-SGD-AP are ${\mathcal{O}}\left(M\sigma^{2}\varepsilon^{-2}\right)$, ${\mathcal{O}}\left((M+B-1)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$, and ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$, respectively.
+Takeaway. Analog-SGD-AP maximizes the computation density of DNN training only at the cost of a higher-order error term compared with the synchronous pipeline.
+CIFAR10
 speedup accuracy (mid) accuracy (end)
@@
 asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07
-Figure 4: Training ResNet10 on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
-(Left) Accuracy-vs-Epoch reflects sample complexity.
-(Right) Accuracy-vs-Clock cycle.
-(Bottom) Speedup and test accuracy.
-The speedup is computed by comparing the clock cycle to reach 93% accuracy.
-The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
-Remark 2.
-For digital SGD with the synchronous pipeline training, the convergence rate is given in [30].
-We also provide the convergence analysis for SGD with asynchronous pipeline training on digital devices, detailed in Theorem 2 in the Appendix F.
-Unlike analog training, both synchronous and asynchronous pipeline training do not slow down digital SGD.
-4.3 Sampling complexity and clock cycle complexity
-The convergence rate alone does not indicate the total training time, as different algorithms require varying numbers of gradient updates.
-The bottleneck during a micro-batch process is the forward and backward computations, which depend on the number of micro-batches.
-Thus, we focus on the sample complexity of the algorithms, defined as the minimal gradient computation needed to achieve the desired error.
-Due to the existence of asymptotic error, we define the sample complexity of synchronous and asynchronous pipelines are the minimum micro-batch needed for achieving $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e.
-$\sigma^{2}S$ for vanilla model parallelism without pipeline/synchronous pipeline and $\sigma^{2}S^{\prime}$ for asynchronous pipeline, respectively.
-Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, and omitting the constant $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$ or $\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}$, we have the following corollary.
-Corollary 1 (Sample complexity).
-The sample complexities of model parallelism without pipeline and synchronous pipeline are $O\left(\sigma^{2}\varepsilon^{-2}\right)$, while that of the asynchronous pipeline is $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$.
-Corollary 1 claims that three model parallelisms share the same dominant term $O(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while asynchronous pipeline slows down by a factor and a non-dominant term $O(\varepsilon^{-1})$.
-Dividing the sample complexity by the computation density, we know the number of clock cycles they need to achieve the $\varepsilon$ accuracy.
-Corollary 2 (Clock cycle complexity).
-The clock cycle complexities of vanilla model parallelism without pipeline, synchronous pipeline, and asynchronous pipeline are $O\left(M\sigma^{2}\varepsilon^{-2}\right)$, $O\left((M+B-1)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$, and $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$, respectively.
-The results are summarized in Table 1.
-By overlapping the computation among micro-batches, synchronous pipeline achieves $M$ times acceleration compared to its non-pipeline counterpart without any cost.
-Asynchronous pipeline overlaps more computation, achieving higher computation density with slight scarification of convergence rate.
-When the mini-batch is split into sufficient micro-batch so that $B\gg M$, the bubble time can be omitted, and synchronous pipeline achieves the best clock cycle complexity.
-However, when the number of mico-batch is small, the asynchronous pipeline is able to achieve better acceleration.
+CIFAR100
 speedup accuracy (mid) accuracy (end)
@@
 asynchronous 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
-Figure 5: Training ResNet10 on CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
-(Left) Accuracy-vs-Epoch reflects sample complexity.
-(Right) Accuracy-vs-Clock cycle.
-(Bottom) Speedup and test accuracy.
-The speedup is computed by comparing the clock cycle to reach 72% accuracy.
+TABLE II: Training ResNet10 on CIFAR10/CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
+The speedup is computed by comparing the number of clock cycles required to achieve 93% accuracy.
 The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
-5 Numerical Simulations
-In this section, we test our methods in an open-source AIHWKIT (AIHWKIT[31]; see github.com/IBM/aihwkit.
-In the simulations, we train a ResNet10 to perform an image classification task on CIFAR10 and CIFAR100 datasets; see Figure 4 and 5.
-The network consists of $M=6$ stages, consisting 1 convolution layer, 4 residual blocks, and 1 fully-connected layer.
-We map the first 5 stages on digital devices and the last one onto an AIMC (analog) device.
-We use digital SGD, implemented by PyTorch  library.
-The pipeline parallelism is also simulated on a single GPU.
-To get the equivalent clock cycle in the figures, we map the epoch to the clock cycle by dividing the computation density, as discussed in Section 3.
-Each simulation is repeated three times, and the mean and standard deviation are reported.
-More details are reported in Appendix G.
-[…]
 $B_{\text{micro}}$ Synchronous Asynchronous
@@
 $128$ 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07 4.89 93.36  $\pm$ 0.06 93.37  $\pm$ 0.06
-Table 2: The training of Resnet10 model on CIFAR10 dataset.
-The test accuracy and standard deviation under different micro-batch sizes at clock cycle 300 (mid) and epoch 300 (end).
-The speedup is computed by comparing the clock cycle needed to reach 93% target accuracy.
+TABLE III: Training Resnet10 model on CIFAR10 dataset under different micro-batch size setting.
+Test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
+The speedup is computed by comparing the number of clock cycles required to achieve 93% target accuracy.
 The accuracy of synchronous is not affected by micro-batch size.
+IV Numerical Simulations
+In this section, we aim to numerically verify the theoretical benefits of Analog-SGD-AP.
+We test the analog methods in an open-source AIHWKIT (AIHWKIT[38]; see github.com/IBM/aihwkit.
+The pipeline parallelism is also simulated on a single GPU.
+To get the equivalent clock cycle in the figures, we map the epoch to clock cycle by dividing the computation density.
+Simulation settings.
+In the simulations, we train ResNet models on two benchmark datasets: CIFAR10 and CIFAR100.
+The CIFAR10 dataset consists of 60,000 $\times32$ color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds.
+The CIFAR100 dataset is similar in format but includes 100 classes.
+Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default.
+The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage numbers $M=12,16$, the initial learning rate is $\alpha=0.01$.
+The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training.
+The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs.
+To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [39].
+The other setting follows the default settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits).
+Noise and bound management techniques are used in [40].
+A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
+Sample complexity and wall-clock complexity.
+We train a ResNet10 for an image classification task on CIFAR10.
+We map the first 5 stages on digital accelerators while the last one onto an AIMC (analog) accelerator.
+The results on the left of Fig.
+4 show that the accuracy-epoch curves of all three pipeline strategies on AIMC accelerators almost overlap and even match the performance of digital SGD, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1.
+Moreover, as indicated by Corollary 2, the walk-clock complexity of the Analog-SGD-AP pipeline is the lowest owing to the highest computation density.
+The results on the CIFAR100 training are reported in Table III, whose conclusion is the same as above.
+Effect of stage/accelerator number.
+We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Figure 4, Right.
+In the simulations, we train a ResNet34 model [41] on the CIFAR10 dataset.
+The model is separated into $M=1,2,4,8,12,16$ stages.
+Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed.
+The results reveal that the asynchronous pipeline achieves linear speedup at least within the range of 181-8 accelerators.
 Effect of micro-batch size.
-We also report the effect of micro-batch sizes in Table 2.
+We also report the effect of micro-batch sizes in Table III.
 We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size.
-For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog devices.
+For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog accelerators.
 Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version.
-The trade-off is that the asynchronous pipeline sacrifices the accuracy and can only reach lower accuracy when the training time is sufficiently long.
-Effect of stage/device number.
-We extensively explored the impact of increasing device count on the speedup achieved through pipeline parallelism; see Figure 1.
-In the simulations, we train Resnet34 model on CIFAR10 dataset, which could be separated into $M=1,2,4,8,12,16$ stages.
-Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed.
-The results reveal that the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices.
-Even though some latency, like communication latency, are ignored in the simulations, the comparison is still fair since the asynchronous pipeline does not introduce obvious extra latency.
-6 Conclusions and Limitations
-This paper study the pipeline algorithms for AIMC accelerators in the convergence and computation aspects.
-We study the existing synchronous pipeline algorithm and propose the first asynchronous pipeline algorithm on AIMC accelerators.
-The synchronous pipeline achieves the best convergence rate in terms of sample complexity, but training bubbles reduce computation density.
-To enable more devices to be active, an asynchronous pipeline discards the synchronization at the end of one mini-batch, achieving a higher computation density, with the cost of slightly increased sample complexity.
-Simulations reveal the extraordinary speedup potential of our asynchronous pipeline for analog computing.
-A limitation of this work is that AIMC accelerator behaviors can only be simulated but not on the actual analog chips.
-But this is a necessary step for every new technology, and future work should verify its efficiency as the technology matures.
-Supplementary Material for “Pipeline Gradient-based Model Training on Analog In-memory Accelerator”
-Appendix A Analog pipeline SGD from worker and data perspectives
-Involving multiple devices and data at each time step, pipeline algorithms, especially asynchronous pipeline, are usually hard to figure out.
+The trade-off is that the asynchronous pipeline sacrifices accuracy and can only reach lower accuracy when the training time is sufficiently long.
+V Conclusions and Limitations
+This paper theoretically investigates the pipeline algorithms for AIMC accelerators with respect to convergence and computation complexity.
+To unlock the full potential of pipeline training, asynchronous pipeline discards the synchronization at the end of one mini-batch to maximize the accelerator usage.
+We theoretically demonstrate that Analog-SGD-AP achieves maximal computational density with only a slight sacrifice in sample complexity through higher-order terms, which brings extraordinary speedup on wall-clock complexity.
+Simulations validate our theoretical findings.
+A limitation of this work is the lack of real hardware validation, but this is a necessary step for every new technology and is widely used in current research regarding the algorithm development for AIMC accelerators.
+Future work should verify its efficiency as the technology becomes mature.
+Appendix
+Appendix A Analog SGD with Asynchronous Pipeline in Different Perspective
+Involving multiple accelerators and data at each time step, pipeline algorithms, especially the asynchronous pipeline, are usually hard to figure out.
 To facilitate the understanding, we list the Analog SGD under two different perspectives.
-In worker perspective, we pay more attention to how worker $m$ processes a flow of data; see Algorithm 1.
-In data perspective, we focus on how a sample $\xi_{k}$ flows through each worker; see Algorithm 2.
-Algorithm 1 Analog SGD with asynchronous pipeline (Worker Perspective)
+In accelerator perspective, we pay more attention to how accelerator $m$ processes a flow of data; see Algorithm 1.
+In data perspective, we focus on how a sample $\xi_{k}$ flows through each accelerator; see Algorithm 2.
+To make the relationship clearer, the computation dependency is plotted in Fig.
+3.
+Algorithm 1 Analog SGD with asynchronous pipeline (accelerator Perspective)
 1:for iteration $k=0,1,2,\cdots,K$ do
-2:// Periphery device (worker 0) at the beginning
-3:     Sample $\xi_{k}=(x_{k},y_{k})$ and Send $x^{(0)}_{k}=x_{k}$ to worker $1$
-4:// Worker $m=1,2,\cdots,M$ in parallel
-5:     Receive input ${\tilde{x}}^{(m)}_{k}$ from worker $m-1$
+2:// Periphery device at the beginning
+3:  Sample $\xi_{k}=(x_{k},y_{k})$ and Send $x^{(0)}_{k}=x_{k}$ to accelerator 11
+4:// Accelerator $m=1,2,\cdots,M$ in parallel
+5:  Receive input ${\tilde{x}}^{(m)}_{k}$ from accelerator m1m-1
 6:  Compute ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$
@@
 8:  Store input ${\tilde{x}}^{(m)}_{k}$ and $\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
-9:     Receive backward error $\tilde{\delta}^{(m+1)}_{k-(M-m)}$
+9:  Receive backward signal $\tilde{\delta}^{(m+1)}_{k-(M-m)}$
 10:  $\tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$
@@
 13:  Compute $\tilde{\delta}^{(M)}_{k}=\nabla_{{\tilde{x}}^{(M+1)}}\ell({\tilde{x}}^{(M+1)},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)})$
-14:     Send $\tilde{\delta}^{(M)}_{k}$ back to worker $M$
+14:  Send $\tilde{\delta}^{(M)}_{k}$ back to accelerator $M$
 15:end for
@@
 2:  Sample $\xi_{k}=(x_{k},y_{k})$, $x^{(0)}_{k}=x_{k}$
-3:     for all worker $m=1,2,\cdots,M$ do
+3:  for all accelerator $m=1,2,\cdots,M$ do
 4:   Compute ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$
@@
 12:end for
-Appendix B Bounds of delay errors
-In the asynchronous pipeline, the computations of both forward and backward passes involve stale weights, which induces delay error on the signals.
-This section provides a series of the bounds of the delay error.
+Appendix B Useful Lemmas and their proof
+We first bound the error of the stale variables. Define $\Delta^{(m)}_{k}:=W^{(m)}_{k+1}-W^{(m)}_{k}$, which is the different between consecutive weights.
+Lemma 5 (Delay error of signals, restatement of Lemma 3).
+Under Assumption 24, the difference between the latest and stale signals generated by Analog-SGD-AP are bounded, i.e., $\forall m\in[M+1]$
+$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (28)
+$\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\sum_{m^{\prime}=1}^{m}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (29)
+$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (30)
+where the constants above are defined by
+$C_{x}^{(m)}:=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1},$ (31)
+$C_{z}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{m-1},$ (32)
+$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (33)
+$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
+Proof of Lemma 3 and Lemma 5.
+We first bound the delay of the forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$. For other layers $m>1$
+$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|\leq L_{g,0}\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|$ (34)
+$=$ $\ L_{g,0}\|W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}-W^{(m)}_{k}x^{(m)}_{k}\|$
+$=$ $\ L_{g,0}\|(W^{(m)}_{k-(M-m)}-W^{(m)}_{k}){\tilde{x}}^{(m)}_{k}+W^{(m)}_{k}({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
+$\overset{(a)}{\leq}$ $\ L_{g,0}(\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}\|+\|W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
+$\leq$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
+where (a) follows the triangle inequality and the consistency of $\ell_{2}$-matrix norm, i.e.
+$\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively.
+Using triangle inequality over the first term of RHS in (34), we have
+$\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|\leq\sum_{k^{\prime}=k-(M-m)}^{k-1}\|W^{(m)}_{k^{\prime}+1}-W^{(m)}_{k^{\prime}}\|=\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|$ (35)
+and consequently,
+$\|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m}\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$.
+(36)
+Expanding inequality (36) from $m$ to $1$ and using the fact ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$ yield
+$\ \|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$ (37)
+$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{k^{\prime}=k-(M-m+1)}^{k-1}\|\Delta^{(m-1)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m-1)}_{k}-x^{(m-1)}_{k}\|$
+$\leq$ $\ L_{g,0}\sum_{m^{\prime}=2}^{m}(L_{g,0}W_{\max,2})^{m^{\prime}-1}(L_{g,0}W_{\max,2})^{m-m^{\prime}}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$
+$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=2}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$.
+$=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
+where the last inequality reindexes $m^{\prime}$ in the first summation.
+Noticing that (34)-(37) also implies that
+$\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$. (38)
+We then bound the delay of the backward signal $\tilde{\delta}^{(m)}_{k}$ for $m\in[M-1]$
+$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|=\|\delta^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}(z^{(m)}_{k})-\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (39)
+$\leq$ $\ \|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|$
+$\ +\|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})-\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$
+$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|+W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|$
+where the last inequality comes from Assumption 3 (1)-(2) and Lemma 8. Substituting (38) into (39), we can further bound (39) by
+$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|$ (40)
+$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
+$\ +W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
+$=$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|+(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
+$\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|+(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\times$
+$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
+To bound the first term in the RHS of (40), we manipulate the different $\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ by
+$\ \|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ (41)
+$=$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})\nabla g^{(M)}(z^{(M)}_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})\|$
+$\leq$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}(z^{(M)}_{k})\|$
+$\ +\|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})-\nabla g^{(M)}(z^{(M)}_{k})\|$
+$\overset{(a)}{\leq}$ $\ L_{\ell,1}L_{g,0}\|x^{(M+1)}_{k}-{\tilde{x}}^{(M+1)}_{k}\|+L_{\ell,0}L_{g,1}\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
+$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
+$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
+where (a) comes from Assumption 3 and Assumption 4.
+To bound the second term in the RHS of (40), we change the upper bound of the second summation from $m^{\prime\prime}$ to $M$ and get
+$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (42)
+$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
+$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
+Substituting (41) and (42) back into (40) we have
+$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (43)
+where the constant $C_{\delta}^{(m)}$ is defined by
+$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (44)
+$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
+This completes the proof. ∎
+Lemma 6 (Delayed error of gradient).
+Denote the stale gradient as
+$\tilde{\nabla}_{W^{(m)}}^{k}f:=\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$. (45)
+The gradient delay is bounded by
+$\|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|\leq$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$. (46)
+where the constant $C_{f}^{(m)}$ is defined by
+$C_{f}^{(m)}:=(L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)}.$ (47)
+Consequently, the norm square of the gradient delay is bounded by
+$\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}\leq$ $\ C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (48)
+where the constant $C_{f}$ is defined by
+$C_{f}:=M\sqrt{\sum_{m=1}^{M}(C_{f}^{(m)})^{2}}.$ (49)
+Proof of Lemma 6.
+By definition, the delay error is bounded by
+$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|\|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\delta^{(m)}_{k}\otimes x^{(m)}_{k}\|$ (50)
+$=$ $\ \|(\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k})\otimes{\tilde{x}}^{(m)}_{k}+\delta^{(m)}_{k}\otimes({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|\leq\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}\|+\|\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$.
+According to Lemma 8 and Lemma 5, it follows that
+$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (51)
+$\leq$ $\ ((L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)})\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
+$=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
+It completes the proof. ∎
+Bounds of delay errors.
+In Analog-SGD-AP, the computations in both the forward and backward passes involve stale weights, which introduce delay error in the signals.
+This section presents bounds on the delay error.
 Lemma 7 (Lipschitz continuity of loss function).
@@
 $\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}\|$,
-$\|\tilde{\delta}^{(m)}_{k}\|=$ $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|,$ (18)
+$\|\tilde{\delta}^{(m)}_{k}\|$= $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (59)
 $\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\tilde{\delta}^{(m+1)}_{k}\|$.
 The proof has finished. ∎
-We first bound the error of the stale variables. Define $\Delta^{(m)}_{k}:=W^{(m)}_{k+1}-W^{(m)}_{k}$, which is the different between consecutive weights.
-Lemma 4 (Delay error of signals).
-Under Assumption 24, the difference between the latest and stale signals are bounded, i.e., $\forall m\in[M+1]$
-$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (19)
-$\|{\tilde{z}}^{(m-1)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\sum_{m^{\prime}=1}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (20)
-$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (21)
-where the constants above are defined by
-$C_{x}^{(m)}:=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1},$ (22)
-$C_{z}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{m-1},$ (23)
-$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (24)
-$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
-Proof of Lemma 4.
-We first bound the delay of forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$. For other layers $m>1$
-$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|\leq L_{g,0}\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|$ (25)
-$=$ $\ L_{g,0}\|W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}-W^{(m)}_{k}x^{(m)}_{k}\|$
-$=$ $\ L_{g,0}\|(W^{(m)}_{k-(M-m)}-W^{(m)}_{k}){\tilde{x}}^{(m)}_{k}+W^{(m)}_{k}({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
-$\overset{(a)}{\leq}$ $\ L_{g,0}(\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}\|+\|W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
-$\leq$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
-where (a) follows the triangle inequality and the consistence of $\ell_{2}$-matrix norm, i.e.
-$\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrix and vector, respectively.
-Using triangle inequality over the first term of RHS in (25), we have
-$\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|\leq\sum_{k^{\prime}=k-(M-m)}^{k-1}\|W^{(m)}_{k^{\prime}+1}-W^{(m)}_{k^{\prime}}\|=\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|$ (26)
-and consequently,
-$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|$ (27)
-$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m}\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
-Expanding inequality (27) from $m$ to $1$ and using the fact ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$ yield
-$\ \|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$ (28)
-$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{k^{\prime}=k-(M-m+1)}^{k-1}\|\Delta^{(m-1)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m-1)}_{k}-x^{(m-1)}_{k}\|$
-$\leq$ $\ L_{g,0}\sum_{m^{\prime}=2}^{m}(L_{g,0}W_{\max,2})^{m^{\prime}-1}(L_{g,0}W_{\max,2})^{m-m^{\prime}}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$
-$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=2}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|.$
-$=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
-where the last inequality reindexes $m^{\prime}$ in the first summation.
-Noticing that (25)-(28) also implies that
-$\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$ (29)
-We then bound the delay of the backward signal $\tilde{\delta}^{(m)}_{k}$ for $m\in[M-1]$
-$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|=\|\delta^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}(z^{(m)}_{k})-\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (30)
-$\leq$ $\ \|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|$
-$\ +\|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})-\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$
-$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|+W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}\|z^{(m)}_{k}-{\tilde{z}}^{(m)}_{k}\|$
-where the last inequality comes from Assumption 3 (1)-(2) and Lemma 3. Substituting (29) into (30), we can further bound (30) by
-$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|$ (31)
-$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
-$\ +W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
-$=$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
-$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
-$\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|+(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\times$
-$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
-To bound the first term in the RHS of (31), we manipulate the different $\|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ by
-$\ \|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ (32)
-$=$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})\nabla g^{(M)}(z^{(M)}_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})\|$
-$\leq$ $\ \|\nabla_{x^{(M+1)}}\ell(x^{(M+1)}_{k},y_{k})-\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}(z^{(M)}_{k})\|$
-$\ +\|\nabla_{x^{(M+1)}}\ell({\tilde{x}}^{(M+1)}_{k},y_{k})\|\|\nabla g^{(M)}({\tilde{z}}^{(M)}_{k})-\nabla g^{(M)}(z^{(M)}_{k})\|$
-$\overset{(a)}{\leq}$ $\ L_{\ell,1}L_{g,0}\|x^{(M+1)}_{k}-{\tilde{x}}^{(M+1)}_{k}\|+L_{\ell,0}L_{g,1}\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
-$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})\|{\tilde{z}}^{(M)}_{k}-z^{(M)}_{k}\|$
-$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
-where (a) comes from Assumption 3 and Assumption 4.
-To bound the second term in the RHS of (31), we change the upper bound of the second summation from $m^{\prime\prime}$ to $M$ and get
-$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (33)
-$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (34)
-$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$ (35)
-Substituting (32) and (33) back into (31) we have
-$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (36)
-where the constant $C_{\delta}^{(m)}$ is defined by
-$C_{\delta}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{M-m+1}(L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}$ (37)
-$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}.$
-This completes the proof. ∎
-Lemma 5 (delay error of gradient).
-Suppose Assumption 24 hold. Denote the stale gradient by
-~$\tilde{\nabla}_{W^{(m)}}^{k}f:=\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}.$ (38)
-The gradient delay is bounded by
-~$\|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|.$ (39)
-where the constant $C_{f}^{(m)}$ is defined by
-$C_{f}^{(m)}:=(L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)}.$ (40)
-Consequently, the norm square of the gradient delay is bounded by
-~$\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}=$ $\ C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (41)
-where the constant $C_{f}$ is defined by
-$C_{f}:=M\sqrt{\sum_{m=1}^{M}(C_{f}^{(m)})^{2}}.$ (42)
-Proof of Lemma 5.
-By definition, the delay error is bounded by
-~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (43)
-$=$ $\ \|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\delta^{(m)}_{k}\otimes x^{(m)}_{k}\|$
-$=$ $\ \|(\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k})\otimes{\tilde{x}}^{(m)}_{k}+\delta^{(m)}_{k}\otimes({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
-$\leq$ $\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}\|+\|\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
-According to Lemma 3 and Lemma 4, it follows that
-~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (44)
-$\leq$ $\ ((L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)})\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
-$=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
-It completes the proof. ∎
 Appendix C Smoothness of the object
-This section proves Lemma 1, i.e. the object is $L$-smooth. Before that, we introduce a lemma as the step stone.
+This section proves Lemma 2, i.e.
+the object is $L$-smooth.
+Before that, we introduce a lemma as the step stone.
 Since the definition of function smoothness involves two weights $W$, $W^{\prime}$ and both of them generate a sequence of forward and backward signals, we require symbols to represent them.
 Using slightly abusing notation in this section, we denote the forward signals computed by $W$ and $\tilde{W}$} as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
-Lemma 6 (Signal stability).
+Lemma 9 (Signal stability, restatement of Lemma 1).
 Denote the forward signals computed by $W$ and $\tilde{W}$} as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
@@
 $\ +3(L_{g,0}W_{\max,2})^{2(M-1)}L_{\ell,0}^{2}L_{g,1}^{2}\sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{2(M-1-m^{\prime\prime})}\frac{1-(L_{g,0}W_{\max,2})^{2(M-m)}}{1-L_{g,0}W_{\max,2}}.$
-Proof of Lemma 6.
+Proof of Lemma 1 and Lemma 9.
 We first bound the delay of forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}=x^{(1)}=x$. For other layers $m\geq 1$
@@
 Expanding inequality (66) from $m$ to $1$ and using the fact ${\tilde{x}}^{(1)}=x^{(1)}=x$ yield
-$\ \|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}$ (52)
-$\leq$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}^{(m-1)}-W^{(m-1)}\|^{2}+2L_{g,0}^{2}W_{\max,2}^{2}\|{\tilde{x}}^{(m-1)}-x^{(m-1)}\|^{2}$
+$\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}\leq$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}^{(m-1)}-W^{(m-1)}\|^{2}+2L_{g,0}^{2}W_{\max,2}^{2}\|{\tilde{x}}^{(m-1)}-x^{(m-1)}\|^{2}$ (67)
 $\leq$ $\ 2L_{g,0}^{2}\sum_{m^{\prime}=2}^{m}(L_{g,0}W_{\max,2})^{2(m^{\prime}-1)}(L_{g,0}W_{\max,2})^{2(m-m^{\prime})}\|{\tilde{W}}^{(m^{\prime}-1)}-W^{(m^{\prime}-1)}\|^{2}$
 $=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=2}^{m}\|{\tilde{W}}^{(m^{\prime}-1)}-W^{(m^{\prime}-1)}\|^{2}$
-$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (53)
+$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
 where the last equality reindexes the summation. By applying the relation
@@
 Substituting (72)–(74) back into (71) we have
-$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq C_{\delta^{\prime}}^{(m)}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (61)
+$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq C_{\delta^{\prime}}^{(m)}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$ (75)
 This completes the proof. ∎
@@
 Proof of Lemma 2 and Lemma 10.
-Analogous to the proof of Lemma (6), we use $x^{(m)}$/$\delta^{(m)}$ and ${\tilde{x}}^{(m)}$/$\tilde{\delta}^{(m)}$ to denote the forward/backward signals related to $W$ and ${\tilde{W}}$, respectively.
+Analogous to the proof of Lemma (9), we use $x^{(m)}$$\delta^{(m)}$ and ${\tilde{x}}^{(m)}$$\tilde{\delta}^{(m)}$ to denote the forward/backward signals related to $W$ and $\tilde{W}$}, respectively.
 It follows by definition of gradient (5) that
-$\ \|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}$ (64)
-$=$ $\ \sum_{m=1}^{M}\|\nabla_{W^{(m)}}f(W;\xi)-\nabla_{W^{(m)}}f({\tilde{W}};\xi)\|^{2}$ (65)
+$\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}=$ $\ \sum_{m=1}^{M}\|\nabla_{W^{(m)}}f(W;\xi)-\nabla_{W^{(m)}}f({\tilde{W}};\xi)\|^{2}$ (78)
 $=$ $\ \sum_{m=1}^{M}\|\tilde{\delta}^{(m)}\otimes{\tilde{x}}^{(m)}-\delta^{(m)}\otimes x^{(m)}\|^{2}$
@@
 $=$ $\ L_{f}^{2}\|{\tilde{W}}-W\|^{2}.$
-Taking square root on both sides on (66) completes the proof. ∎
+Taking the square root on both sides of (79) completes the proof. ∎
 Appendix D Proof of Theorem 1: Analog SGD with synchronous pipeline
@@
 Denote the average gradient by
-¯$\bar{\nabla}f(W_{k};\xi_{k}):=\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b}),$ (68)
+$\bar{\nabla}f(W_{k};\xi_{k}):=\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b}),$ (81)
 which satisfies that ${\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\bar{\nabla}f(W_{k};\xi_{k})]=\nabla f(W_{k})$. Furthermore, the independence of sampling (c.f. Assumption 1) ensures that
@@
 $=$ $\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$
-$=$ $\ \frac{1}{B^{2}}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$
-$\leq$ $\ \sigma^{2}/B.$
+$=$ $\ \frac{1}{B^{2}}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}/B.$
 The $L$-smooth assumption of the objective (c.f. Lemma 10) implies that
@@
 $\ +\alpha^{2}L_{f}\mathbb{E}_{\varepsilon_{k}}[\|\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]+\frac{1}{2\alpha}\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k})+\varepsilon_{k})\|^{2}$
-where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
+where the second inequality comes from the assumption that the expectation of the noise is 0 (Assumption 1)
 $\ \mathbb{E}_{\varepsilon_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]$ (84)
@@
 and the following inequality
-$\ \frac{L_{f}}{2}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (72)
-$\leq$ $\ L_{f}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
+$\frac{L_{f}}{2}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}\|^{2}]\leq$ $\ L_{f}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (85)
 $\ +\alpha^{2}L{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}].$
 With the learning rate $\alpha\leq\frac{1}{2L}$ and bounded variance (c.f. (82)), (83) becomes
-$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]$ (73)
-$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}+\frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}].$
-According the dynamic of Analog SGD (5), the last term in the RHS of (73) is upper bounded by
+${\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}$ (86)
+$\ +\frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}].$ (87)
+According the dynamics of Analog-SGD-SP, the last term in the RHS of (86) is bounded by
 $\ \frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}]$ (88)
@@
 By the variance decomposition and Assumption 1, the term above can be bounded by
-${\mathbb{E}}_{\xi_{k,b}}[\|f(W_{k};\xi_{k,b})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$ (75)
-$\leq$ $\ \|\nabla f(W_{k})\|^{2}+\sigma^{2}.$
+${\mathbb{E}}_{\xi_{k,b}}[\|f(W_{k};\xi_{k,b})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]\leq\|\nabla f(W_{k})\|^{2}+\sigma^{2}.$ (89)
 Consequently, combining (86), (88) and (89) yields
@@
 $\leq$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-\frac{\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}+\frac{\alpha\sigma^{2}}{2\tau^{2}}\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}.$
-Organizing the terms, taking expectation, averaging (76) over $k$ from 00 to $K-1$, we obtain
+Organizing the terms, taking expectation, averaging (90) over $k$ from 0 to K1K-1, we obtain
 $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2(f(W_{0})-{\mathbb{E}}[f(W_{k+1})])}{1-W_{\max,\infty}^{2}/\tau^{2}}+\frac{\alpha L_{f}\sigma^{2}}{B(1-W_{\max,\infty}^{2}/\tau^{2})}$ (91)
 $+\frac{\sigma^{2}}{K}\sum_{k=0}^{K-1}\frac{\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}/\tau^{2}}{1-\max_{b\in[B]}\|W_{k,b}\|^{2}_{\infty}/\tau^{2}}.$
-$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}$
+$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S_{K}$
 where the second inequality uses $f(W_{k+1})\geq f^{*}$ and specifies $\alpha=\sqrt{\frac{B(f(W_{0})-f^{*})}{\sigma^{2}L_{f}K}}$. Now we complete the proof. ∎
@@
 $\ +\alpha^{2}L_{f}\mathbb{E}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
-where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
+where the second inequality comes from the assumption that the expectation of the noise is 0 (Assumption 1)
 $\ \mathbb{E}_{\xi_{k}}[\left\langle\nabla f(W_{k}),W_{k+1}-W_{k}\right\rangle]$ (93)
@@
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}].$ (95)
-By the dynamic of Analog SGD, the last in the RHS of (E) is bounded by
+By the dynamics of Analog SGD, the last in the RHS of (E) is bounded by
 $\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$ (96)
@@
 $\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})\mathbb{E}_{\xi_{k}}[\|{\bm{1}}+\frac{1}{\tau}\odot\text{sign}(\tilde{\nabla}^{k}f)\odot W_{k}\|^{2}_{\infty}\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
-$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
-$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}M^{2}C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
+$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
+$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}M^{2}C_{f}^{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}\right].$
 where the last inequality comes from Lemma 6.
@@
 where the last inequality comes from Assumption 1
-${\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (86)
+${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (100)
 $\leq$ $\ \|\nabla f(W_{k})\|^{2}+\sigma^{2}.$
@@
 $\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$ (101)
-$=$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
+$=$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
 $\ +\frac{\alpha}{2\tau^{2}}(1+u)\|\nabla f(W_{k})\|^{2}\|W_{k}\|^{2}_{\infty}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$
@@
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
+$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
 $=$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
+$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\mathbb{E}_{\xi_{k}}\Big[\underbrace{\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}}_{=:\psi_{k}}\Big].$
+To deal with $\psi_{k}$ in the RHS of (102), we define an auxiliary function
+$\Psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (103)
+which has the following recursion from $k$ to $k+1$
+$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (104)
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+$\ +M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
+$\leq$ $\ \Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$
+where the last inequality holds by definition of $\Delta^{(m^{\prime})}_{k}$, i.e.,
+$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=\sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}.$ (105)
 Construct a Lyapunov function by
-${\mathbb{V}}_{k}:=$ $\ f(W_{k})-f^{*}$ (89)
-$+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
-Notice that the last term above has the following recursion from $k$ to $k+1$
-$\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-(k+1)+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (90)
-$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
-$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
-$\ +(M-m+1)\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
+${\mathbb{V}}_{k}:=f(W_{k})-f^{*}+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\Psi_{k}.$ (106)
 According to (102), we have
-$\ {\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]$ (91)
-$=$ $\ f(W_{k+1})-f^{*}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-{k+1}+M+1)\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|^{2}]$
+${\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]=$ $\ f(W_{k+1})-f^{*}+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\Psi_{k+1}$ (107)
 $\leq$ $\ f(W_{k})-\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-$\ -\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
-$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
-By the definition of $\Delta^{(m)}_{k}$ and the dynamic of Analog SGD, we bound the last term of the RHS of (91) by
-$\ \sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}=\sum_{m^{\prime}=1}^{M}\|W^{(m)}_{k+1}-W^{(m)}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}$ (92)
+$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
+we bound the last term of the RHS of (107) by
+$\ \|W_{k+1}-W_{k}\|^{2}$ (108)
 $\leq$ $\ 2\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}+2\alpha^{2}\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}$
@@
 $\leq$ $\ {\mathbb{V}}_{k}-{\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}$
-$\ -\left(\frac{1}{4\alpha}-\alpha(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\right)\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}.$
-Taking expectation, averaging (93) over $k$ from 00 to $K-1$, and choosing the learning rate $\alpha\leq\frac{1}{2\sqrt{1+1/u}(1+W_{\max}/\tau)C_{f}M}$ yield
+$\ -\left(\frac{1}{4\alpha}-\alpha(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\right)\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
+$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}.$
+Taking expectation, averaging (109) over $k$ from 0 to K1K-1, and choosing the stepsize $\alpha\leq\frac{1}{2\sqrt{1+1/u}(1+W_{\max,\infty}/\tau)C_{f}M}$ yield
 $\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]$ (110)
-$\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{k+1}])}{\alpha K(1-(1+u)W_{\max}^{2}/\tau^{2})}+\frac{2\alpha L_{f}\sigma^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$
-$\ +\sigma^{2}\frac{1}{K}\sum_{k=0}^{K-1}\frac{(1+u)\|W_{k}\|^{2}_{\infty}/\tau}{1-(1+u)\|W_{k}\|^{2}_{\infty}/\tau^{2}}+\frac{\alpha^{2}(1+1/u)(1+W_{\max}/\tau)^{4}C_{f}^{2}M^{3}\sigma^{2}}{(1-(1+u)W_{\max}^{2}/\tau^{2})}.$
+$\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{k+1}])}{\alpha K(1-(1+u)W_{\max,\infty}^{2}/\tau^{2})}+\frac{2\alpha L_{f}\sigma^{2}}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$
+$\ +\sigma^{2}\frac{1}{K}\sum_{k=0}^{K-1}\frac{(1+u)\|W_{k}\|^{2}_{\infty}/\tau}{1-(1+u)\|W_{k}\|^{2}_{\infty}/\tau^{2}}+\frac{\alpha^{2}(1+1/u)(1+W_{\max,\infty}/\tau)^{4}C_{f}^{2}M^{3}\sigma^{2}}{(1-(1+u)W_{\max,\infty}^{2}/\tau^{2})}.$
 Using the fact that $V_{k}\geq 0$ for any $k\in{\mathbb{N}}$ and $V_{0}=f(W_{0})-f^{*}$, and choosing the learning rate $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$, we have
 $\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]$ (111)
-$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}^{\prime}+O\left(\frac{1+1/u}{K}\right).$
+$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S_{K}^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right).$
 where $S_{K}^{\prime}$ denotes the amplification factor given by
 $S_{K}^{\prime}:=\frac{1}{K}\sum_{k=0}^{K}\frac{(1+u)\|W_{k}\|_{\infty}^{2}/\tau^{2}}{1-(1+u)\|W_{k}\|_{\infty}^{2}/\tau^{2}}\leq\frac{(1+u)W_{\max}^{2}/\tau^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}.$ (112)
-Appendix F Convergenc of digital asynchronous pipeline SGD
-Theorem 3 (Convergence rate, asynchronous pipeline).
+The proof is completed now. ∎
+Appendix F Convergence of Digital Asynchronous Pipeline SGD
+Theorem 3 (Iteration complexity, asynchronous pipeline).
 Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$, it holds that
-[…]
+$\ \frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]\leq\ 4\sqrt{\frac{(f(W_{0})-f^{*})L_{f}\sigma^{2}}{K}}+\frac{2C_{f}M^{3}(f(W_{0})-f^{*})}{KL_{f}}.$ (113)
 Proof.
-The validation of inequality (78) does not rely on the any training algorithm and hence it still holds here. To bound the last term in the RHS of (78), we have
-$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
-$=$ $\ \frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
-$\leq$ $\ \frac{\alpha C_{f}}{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|\right]^{2}$
-$\leq$ $\ \frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$ (98)
+The validation of inequality (92) does not rely on any training algorithm, and hence it still holds here. To bound the last term in the RHS of (92), we have
+$\ \frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]=\frac{\alpha}{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
+$\leq$ $\ \frac{\alpha C_{f}}{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|\right]^{2}\leq\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$ (114)
 where the first inequality is according to Lemma 6.
 According to the descent lemma in (E), we have
-$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
+$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]$ (115)
+$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{1}{2\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha\nabla f(W_{k};\xi_{k})\|^{2}]$
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
 $\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
-$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (99)
-where the second inequality is earned by plugging in (F).
-Construct a Lyapunov function as
+$\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
+where the second inequality is earned by plugging in (F). Construct a Lyapunov function as
 ${\mathbb{V}}_{k}:=f(W_{k})+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M-m^{\prime}+2)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$ (116)
@@
 where the fourth inequality holds because
-$\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$ $=\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}^{(m^{\prime})}-W_{k}^{(m^{\prime})}\|^{2}]$
-$={\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
+$\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k}\|^{2}]$ $=\sum_{m^{\prime}=1}^{M}{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}^{(m^{\prime})}-W_{k}^{(m^{\prime})}\|^{2}]={\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (117)
 $\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
 $~~~~+2\alpha^{2}{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$
-$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]+2\alpha^{2}\sigma^{2}.$ (101)
-and the last inequality is earned by letting $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$. Reorganizing the terms yield
+$\leq 2{\mathbb{E}}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]+2\alpha^{2}\sigma^{2}.$
+and the last inequality is earned by letting $\alpha\leq\sqrt{\frac{1}{4C_{f}M^{3}}}$. Reorganizing the terms yields
 $\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}\leq$ $\ {\mathbb{V}}_{k}-{\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]+\alpha^{2}L_{f}\sigma^{2}+\alpha^{3}C_{f}M^{3}\sigma^{2}.$
 Therefore, by telescoping and taking the expectation, we get
-$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{K}])}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}$
+$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ \frac{2({\mathbb{V}}_{0}-{\mathbb{E}}[{\mathbb{V}}_{K}])}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}$ (118)
 $\leq\frac{2(f(W_{0})-f^{*})}{\alpha K}+2\alpha L_{f}\sigma^{2}+2\alpha^{2}C_{f}M^{3}\sigma^{2}.$
@@
 $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})L_{f}\sigma^{2}}{K}}+\frac{2C_{f}M^{3}(f(W_{0})-f^{*})}{KL_{f}}.$
-The proof is completed. ∎
-layer name output size 10-layer 34-layer
-conv1 112$\times$112 7$\times$7, 64, stride 2
-conv2_x 56$\times$56 3$\times$3 max pool, stride 2
-$\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 64}\\[-1.00006pt] \text{3$\times$3, 64}\end{array}\right]$$\times$3
-conv3_x 28$\times$28 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 128}\\[-1.00006pt] \text{3$\times$3, 128}\end{array}\right]$$\times$4
-conv4_x 14$\times$14 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 256}\\[-1.00006pt] \text{3$\times$3, 256}\end{array}\right]$$\times$6
-conv5_x 7$\times$7 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$1 $\left[\begin{array}[]{c}\text{3$\times$3, 512}\\[-1.00006pt] \text{3$\times$3, 512}\end{array}\right]$$\times$3
-fc 1$\times$1 average pool, fully connected, softmax
-block number 6 18
-Table 3: ResNet models in the simulations. Downsampling is performed by conv3_1, conv4_1, and conv5_1 with a stride of 2.
-Appendix G Simmulation details
-This section provides details about the experiments in Section 5.
-The analog training algorithm, i.e., Analog SGD, is provided by the open-source simulation toolkit AIHWKIT [31], which has Apache-2.0 license; see github.com/IBM/aihwkit.
-We use Softbound device provided by AIHWKIT to simulate the asymmetric linear device (ALD), by setting its upper and lower bound as $\tau$.
-The digital algorithm, including SGD, and datasets used in this paper, including CIFAR10 and CIFAR-100, are provided by PyTorch, which has BSD license; see https://github.com/pytorch/pytorch.
-The other setting follows the standard settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits).
-Noise and bound management techniques are used in [33].
-A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
-Compute Resources. We conduct our simulations on an NVIDIA RTX 3090 GPU, which has 24GB memory. The simulations take from 30 minutes to 7 hours, depending on the size of the model and the dataset.
-Statistical Significance.
-The simulations reported in Figure 2 and 4 are repeated three times.
-The randomness originates from the data shuffling, random initialization, and random noise in the analog device simulator.
-The mean and standard deviation are calculated from the final output of each run using NumPy library.
-Image classification on CIFAR10/CIFAR100 with Resnet models.
-In the simulations, we train ResNet models on two benchmark datasets: CIFAR-10 and CIFAR-100.
-The CIFAR-10 dataset consists of 60,000 32$\times$32 color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds.
-The CIFAR-100 dataset is similar in format but includes 100 classes.
-In Figure 4 and 5, ResNet10 is chosen while in Figure 1, a larger model, ResNet34 is chosen.
-As Table 3 shows, ResNet10 and ResNet34 are composed of 6 and 18 blocks, respectively.
-In Figure 4 and 5, each stage contains one block while in Figure 1, 18 blocks are separated into $M=1,2,4,8,12,16$ stages as evenly as possible.
-Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default.
-The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage number $M=12,16$, the initial learning rate is $\alpha=0.01$.
-The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training.
-The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs.
-To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [34].

### v3

@@
 Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications.
-As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive.
-For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [1].
+As the model sizes and dataset volumes grow exponentially, DNN training becomes time-consuming and resource-intensive.
+For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [touvron2023llama2].
 In response to the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators has emerged.
@@
 The promise of analog computing.
-To this end, a promising solution to accelerate the MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage.
-The DNN weights are stored in analog crossbar tiles consisting of resistive elements [2] where the conductance of the resistive element represents the weight magnitude.
+To this end, a promising solution to accelerate MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage.
+The DNN weights are stored in analog crossbar tiles consisting of resistive elements [jain2019neural] where the conductance of the resistive element represents the weight magnitude.
 Unlike traditional Von Neumann accelerators such as GPUs and TPUs, which move weights from memory to the processor during computation, AIMC accelerators keep the weights stationary and perform MVM operations in memory.
@@
 Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training.
-In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different accelerators to support gradient descent with large batches [4, 5].
+In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [li2020pytorch], where DNN weights are replicated to different accelerators to support gradient descent with large batches [goyal2017accurate, you2020Large].
 By contrast, in the AIMC architecture, trainable weights are stored in the analog tiles as the conductances of resistive elements; thus, copying weights is expensive and prone to error.
 Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
-Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [6].
+Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [xu2021parallelizing].
 Pipeline parallelism partitions a large model into a series of stages distributed across multiple accelerators, with each stage corresponding to the computation of a few layers.
@@
 Synchronous and asynchronous parallelism.
-As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [7, 8], Deepspeed [9], FSDP [10], Megetron-LM [11], Colossal-AI [12], just to name a few.
+As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [huang2019gpipe, kim2020torchgpipe], Deepspeed [ren2021zero], FSDP [zhao2023pytorch], Megetron-LM [shoeybi2019megatron], Colossal-AI [li2023colossal], just to name a few.
 In their implementation, a gradient mini-batch (e.g., with batch size 128) is split into micro-batches (e.g., 8 micro-batches with batch size 16) and filled into the pipeline.
@@
 How do their combined effects provably impact training on the AIMC accelerator?
-Despite two sources of bias, this paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and SGD on AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
+This paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
 I-A Main results
@@
 We highlight the unique technical challenges compared with the analysis of the asynchronous pipeline on digital accelerators as follows.
-  1. T1)
-    Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias.
+T1) Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias.
 To be specific, consider an analog weight $W_{k}$ on a resistive crossbar array and a desired update $\Delta W_{k}$.
 The updated weight becomes
-    $\textsf{Analog Update}\quad W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}^{(m)}$ (1)
+$W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}$ (1)
 where $\tau$ is a hardware-specific constant, $|\,\cdot\,|$ is element-wise absolute value, and $\odot$ is the element-wise production.
-Unlike the accurate digital training, the last bias term in (1) is unique to analog accelerators.
+Unlike digital training, the last term in (1) is unique to analog accelerators.
 This bias can not be eliminated within a reasonable cost and leads to challenges to the analysis.
 In section II-B, we provide more background.
-  2. T2)
-    Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [13].
+T2) Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [sun2017asynchronous].
 However, in the asynchronous pipeline, the forward pass is performed using stale weights, while the backward pass uses the latest weights.
@@
 The stale signals in DNN render the update increment used in Analog-SGD-AP no longer a valid gradient, making its convergence behavior unclear.
-  3. T3)
-    Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis.
-Without considering a specific multi-layer architecture, prior work [14] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
+T3) Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis.
+Without considering a specific multi-layer architecture, prior work [huo2018decoupled] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
 Pipeline Type Sample Complexity Comp. Density Wall-clock Complexity
@@
 TABLE I: Comparison of different pipeline strategies.
-Sample complexity is the number of samples needed to achieve a training loss $\leq\varepsilon$.
+Sample complexity is the number of samples needed to achieve a training loss $\leq E+\varepsilon$, where $E$ is an inevitable asymptotic error due to hardware imperfections.
 Computing density is the average portion of active accelerators.
@@
 Wall-clock complexity measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by computation density.
-Overcoming the challenges, this paper answers the question above and provides both a convergence guarantee and computational complexities. The key contributions are as follows.
-  1. C1)
-    We consider the training of a multi-layer DNN, which consists of linear layers and non-linear activation functions.
-We formulate the training dynamics of Analog-SGD-AP, which reflects the stale weights in the asynchronous pipeline and the bias from the AIMC accelerators’ imperfection.
-We investigate the smoothness of the training objective as well as the robustness of the network against stale weights.
-  2. C2)
+See Theorems 1 and 2 for iteration complexity analysis; and Corollary 1 and 2 for sample and wall-clock complexity analysis.
 […]
-  3. C3)
-    We verify the speedup of Analog-SGD-AP against Analog-SGD-WOP and Analog-SGD-SP in AIMC accelerator simulator, AIHWKIT.
-Numerical simulation results show that despite a slightly larger sample complexity, Analog-SGD-AP achieves the same accuracy with fewer clock cycles due to its high computation density, which is consistent with the theoretical guarantee.
-Furthermore, empirical results show that Analog-SGD-AP achieves linear speedup at least within the range of 1-8 accelerators.
 I-B Prior art.
-To put our work in context, we review prior art that we group into the following two categories.
+Prior art can be grouped into two categories.
 Gradient-based training on AIMC accelerators.
-Analog training has shown promising early successes in certain tasks [15, 16, 17], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware.
-However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [18].
-Numerous algorithms have been proposed to address these issues.
-TT-v1 [19] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error.
-To demonstrate its efficiency, [20] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee.
-Building on that, [21] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods.
-To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [22] introduces a digital array.
-Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [23] are proposed to mitigate the impact of inaccurate symmetric point correction [24].
-To overcome the limit update granularity issue, [25] proposed a multi-tile training paradigm.
+Analog SGD implemented by pulse update is extremely efficient [gokmen2016acceleration, yao2017face, wang2019situ], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware.
+However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [burr2015experimental].
+Numerous algorithms have been proposed proposed to address these issues.
+TT-v1 [gokmen2020] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error.
+To demonstrate its efficiency, [wu2024towards] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee.
+Building on that, [wu2025analog] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods.
+To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [gokmen2021] introduces a digital array.
+Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [rasch2024fast] are proposed to mitigate the impact of inaccurate symmetric point correction [onen2022neural].
+To overcome the limit update granularity issue, [li2025memory] proposed a multi-tile training paradigm.
 However, the mentioned works focus on training single AIMC accelerators.
@@
 Existing works proposed methods to alleviate the stale weight issues in the asynchronous pipeline.
-For instance, [26] proposed weight-prediction approaches to compensate for the delay, PipeDream [27] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [28] and [29] propose only overlapping gradient computation and communication to ensure the delay is one.
+For instance, [kosson2021pipelined] proposed weight-prediction approaches to compensate for the delay, PipeDream [narayanan2019pipedream] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [chen2022sapipe] and [li2018pipe] propose only overlapping gradient computation and communication to ensure the delay is one.
 These algorithms are not for analog pipeline training.
-Researchers also attempted to design analog pipeline circuits to accelerate the training [30, 31], but they focus on the synchronous pipeline.
-On the theory side, there are few studies on pipeline training for digital accelerators such as [14, 32], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms.
-As [27] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
+Researchers also attempted to design analog pipeline circuits to accelerate the training [shafiee2016isaac, song2017pipelayer], but they focus on the synchronous pipeline.
+On the theory side, there are few studies on pipeline training for digital accelerators such as [huo2018decoupled, xu2020acceleration], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms.
+As [narayanan2019pipedream] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
 II Analog Training with Model Parallelism
@@
 The forward pass of the model for layer $m\in[M]$ is expressed by
-Forward $x^{(1)}=x,~~~~z^{(m)}=W^{(m)}x^{(m)},~~~~x^{(m+1)}=g^{(m)}(z^{(m)})$ (3)
+Forward $x^{(1)}=x,~~~~z^{(m)}=W^{(m)}x^{(m)},$ (3)
+$x^{(m+1)}=g^{(m)}(z^{(m)})$
 where $g^{(m)}(\,\cdot\,)$ is the activation function of layer $m$, and $x^{(m)}$ is the input feature.
@@
 Subsequently, each layer $m\in[M-1]$ computes the next recursion by the chain rule
-$\textsf{Backward}\qquad\delta^{(m)}=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}=(\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$ (4)
+$\textsf{Backward}\qquad\delta^{(m)}=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}$ (4)
+$=$ $\ (\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$
 In the forward pass, the activation $z^{(m)}$ is stashed until the backward signal $\delta^{(m+1)}$ returns.
@@
 With the forward and backward signals, the gradient with respect to the weight $W^{(m)}$ of layer $m$ is computed by
-$\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}=\delta^{(m)}\otimes x^{(m)}$ (5)
+$\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}$ (5)
+$=$ $\ \delta^{(m)}\otimes x^{(m)}$
 where $\otimes$ is the outer product between two vectors.
@@
 Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, rank-update has only $O(BN)$ complexity while the update on digital accelerators has $O(BN^{2})$ complexity to update an $N\times N$ matrix.
-Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [15].
+Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [gokmen2016acceleration].
 Figure 2: Illustration of pipelines with 4 accelerators ($M=4$).
@@
 II-B Non-ideal update on AIMC accelerator
-Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [19, 20, 21], the update on AIMC accelerators exhibits asymmetric bias.
+Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [gokmen2020, wu2024towards, wu2025analog], the update on AIMC accelerators exhibits asymmetric bias.
 To be specific, for any desired update $\Delta W_{k}^{(m)}$, the update on $W_{k}$ is scaled by response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$.
@@
 Defining the symmetric and asymmetric components as $F(w):=\frac{1}{2}(q_{-}(w)+q_{+}(w))$ and $G(w):=\frac{1}{2}(q_{-}(w)-q_{+}(w))$, we write the update in (6) as $U(w,\Delta w)=w+\Delta w\cdot F(w)-|\Delta w|\cdot G(w)$.
-This paper consider a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$ as suggested by [20], we reach the dynamics (1).
-Figure 3: Illustration of the dynamics of the asynchronous pipeline.
-The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each accelerator.
+This paper considers a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$, which leads to the dynamics (1).
+The linear forms of $q_{+}(w)$ and $q_{-}(w)$ can be viewed as a first-order approximation of more general nonlinear device response functions along the training trajectory.
+In this model, $\tau$ controls how strongly the update magnitude depends on the current conductance state.
+Physically, a larger $\tau$ means weaker state dependence (more symmetric updates), while a smaller $\tau$ means stronger asymmetry and stronger bias during training.
+Therefore, $\tau$ is a hardware-specific constant.
+Refer to [wu2024towards] for more details about the background of this model.
+Remark 1.
+As established in [gokmen2020, wu2024towards], the asymmetric update is one of the most significant challenges in analog training, and our theoretical framework explicitly captures this dominant source of non-ideality through the bias term in equation (1).
+In the theoretical analysis, we focus on the asymmetric update as the primary source of hardware imperfection to balance between faithfulness to the hardware and analytical tractability.
+Beyond the asymmetric update, the simulations additionally incorporate a comprehensive set of hardware imperfections to more accurately reflect real device behavior; see Section IV.
+Beyond the update model in (6), additional hardware effects (e.g., output noise) can be viewed as bounded perturbations to the idealized update dynamics.
+Under such perturbations, the dominant convergence trend is expected to be preserved, while constants and the asymptotic error may increase by perturbation-dependent terms.
+A full study that explicitly incorporates these non-idealities is left for future work.
 II-C Synchronous and asynchronous pipelines
@@
 This paper treats a stage as the minimal scheduling unit and investigates the time complexity across different scheduling strategies.
-To compare time costs, we define the basic time unit, clock cycle, as the minimum time required for all devices to process one micro-batch.
+To compare time costs, we define the basic time unit, clock cycle, which is the minimum time required for all devices to process one micro-batch.
+Note that this definition is based on the assumption that per-stage computation loads are balanced across all devices, so that all stages complete each micro-batch in approximately the same amount of time, which can be achieved via balanced partitioning [huang2019gpipe, narayanan2019pipedream, shoeybi2019megatron].
 Based on scheduling strategies, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining.
@@
 Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}^{(m)}$, which iterates over $b\in[B_{\text{mini}}]$ by
-$W_{k,b+1}^{(m)}=$ $\ W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})-\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (7)
-Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$.
-To fully release the potential of rank-update, AIMC accelerators follow (7) to update its weight, which in turn complicates the convergence analysis.
+$W_{k,b+1}^{(m)}=$ $\ W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$
+$\ -\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (7)
+Notice that for digital SGD, the last term in the RHS (RHS) of (II-C) disappears, and hence (II-C) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$.
+To fully release the potential of rank-update, AIMC accelerators follow (II-C) to update its weight, which in turn complicates the convergence analysis.
 Synchronous pipeline.
@@
 However, most devices remain idle, reducing system throughput.
-Existing work has already attempted to design an architecture that supports the pipeline training [30, 31].
-They faithfully implement the iteration (7) but overlap the computing inside a mini-batch.
+Existing work has already attempted to design an architecture that supports the pipeline training [shafiee2016isaac, song2017pipelayer].
+They faithfully implement the iteration (II-C) but overlap the computing inside a mini-batch.
 Featuring synchronization at the end of a mini-batch, this stage schedule mechanism is called synchronous pipeline.
@@
 Consequently, for each mini-batch, most of the devices need to wait until the last micro-batch is processed.
-After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (7).
+After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (II-C).
 To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch.
@@
 The forward and backward passes of asynchronous pipeline training can be summarized as
-Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$.
+Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$,~~ Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$.
 (9)
@@
 After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights.
-However, there is a bias due to the hardware imperfection [20].
+However, there is a bias due to the hardware imperfection [wu2024towards].
 Replacing $\Delta W_{k}$ with $\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ in (1), where $\alpha>0$ is a positive learning rate, we reach the update dynamics of weights
 $\textsf{Update}\quad W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ $\ -\frac{\alpha}{\tau}|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}|\odot W_{k}^{(m)}.$ (10)
-In (10), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
-The main challenge coming from the AIMC training is the weight drift [20].
+In (II-C), the outer product $\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ is not a valid gradient with respect to any weight $W^{(m)}$, which is referred to as partially stale gradient and makes the analysis more challenging.
+The main challenge coming from the AIMC training is the weight drift [wu2024towards].
 Consider the case where $W_{k}^{(m)}$ is already a non-zero critical point of $f(W)$ but not the one of $f(W;\xi)$, i.e.
@@
 After one iteration, the weight $W_{k}^{(m)}$ drifts because of the sample randomness
-${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}\neq 0$ (11)
+${\mathbb{E}}[W_{k+1}^{(m)}]-W_{k}^{(m)}=$ $\ -\frac{\alpha}{\tau}{\mathbb{E}}[|\nabla_{W^{(m)}}f(W_{k}^{(m)};\xi)|]\odot W_{k}^{(m)}$
+$\neq$ $\ 0$ (11)
 which prevents the weights from converging to the critical point, leading to an error.
+Remark 2 (Communication overhead).
+In all three pipeline strategies considered in this paper, inter-accelerator communication consists of passing activation vectors $x^{(m)}$ and backward signals $\delta^{(m)}$ between adjacent stages.
+The per-stage communication cost remains O(D) as M grows.
+Accordingly, deeper pipelines do not increase the asymptotic per-stage communication burden, although the realized wall-clock benefit still depends on bandwidth and latency remaining secondary to per-stage computation.
+Furthermore, the communication overhead can be reduced by overlapping gradient communication with computation as proposed in [chen2022sapipe, li2018pipe], thereby hiding inter-stage communication latency within the computation time of each clock cycle.
 III Convergence & Wall-clock Analysis
@@
 Assumption 3 (Activation function).
-All activation function $g^{(m)}(\,\cdot\,)$, $m\in[M]$, is (i) Lipschitz continuous, i.e.
+All activation functions are (i) Lipschitz continuous, i.e.
 $\|g^{(m)}(z)-g^{(m)}(z^{\prime})\|\leq L_{g,0}\|z-z^{\prime}\|$; (ii) smooth, i.e.
@@
 $g(0)=0$.
-The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
+The conditions (i) and (ii) in Assumption 3 hold for many activation functions like Sigmoid, Tanh, or smoothed ReLU functions, while (iii) can be achieved by adding an offset mapping before the activation function if there is a point $x$ such that $g(x)=0$.
+Remark 3 (Non-smooth activation functions).
+Assumption 3 requires the activation functions to be smooth.
+For non-smooth activations such as ReLU, the signal smoothness (see Lemmas 1, 2 later) cannot be guaranteed in general.
+In this case, one can establish that the functions are semi-smooth [zou2019improved].
+In the overparameterized regime, such a semi-smoothness property can be established by showing that the training trajectory remains within a small neighborhood of initialization where the set of active neurons is nearly fixed.
+Extending the convergence analysis of Analog-SGD-AP to non-smooth activations within this semi-smoothness framework is an interesting direction we leave for future work.
 Assumption 4 (Loss function).
-The loss $\ell(\cdot,\cdot)$ is smooth with respective the first input, i.e.
-$\|\nabla_{1}\ell(y_{1},y)-\nabla_{1}\ell(y_{2},y)\|\leq L_{\ell,1}\|y_{1}-y_{2}\|,\forall y_{1},y_{2},y$; lower bounded, i.e.
-$\ell(y_{1},y_{2})\geq f^{*},\forall y_{1},y_{2}$.
-Aside from the assumptions, to avoid saturation during training, following the setting in [20], the saturation degree is assumed to be bounded.
+The loss $\ell(\cdot,\cdot)$ is: (i) smooth with respect to the first input, i.e., $\|\nabla_{1}\ell(y_{1},y)-\nabla_{1}\ell(y_{2},y)\|\leq L_{\ell,1}\|y_{1}-y_{2}\|$; (ii) lower bounded, i.e., $\ell(y_{1},y_{2})\geq f^{*}$.
+Aside from these assumptions, to avoid saturation during training, following the setting in [wu2024towards], the saturation degree is assumed to be bounded.
 Assumption 5 (Bounded Saturation).
-There exists a positive constant $W_{\max}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$.
-The ratio $W_{\max,\infty}/\tau$ is termed saturation degree.
-From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range [20].
-It is possible since the trainable weights always remain bounded in a small neighborhood of the initialization points during training [33, 34, 35], especially for overparameterized networks in digital training.
-Given $W_{\max,\infty}$ is sufficiently large to include this neighborhood, Assumption 5 holds.
-While it is possible to prove this assumption as in [36], we adopt it to focus on revealing key insights into the effect of pipeline training.
+There exists a positive constant $W_{\max,\infty}<\tau$ such that all the weight $W^{(m)}_{k}$, $m\in[M]$ is $\ell_{\infty}$-norm bounded, i.e., $\|W^{(m)}_{k}\|_{\infty}\leq W_{\max,\infty}<\tau,\forall k\in[K]$.
+The ratio $W_{\max,\infty}/\tau=\Theta(1)<1$ is termed saturation degree.
+From the physical perspective, Assumption 5 requires that the weight can be represented by a conductance state in the dynamic range $(-\tau,\tau)$.
+Actually, the analog update dynamic induced by the hardware naturally constrains the weight, as shown by the following theorem.
+Proposition 1 ([wu2024towards, Theorem 1]).
+Given $\|W_{0}\|_{\infty}<\tau$ and for any sequence $\{\Delta W_{k}:k\in{\mathbb{N}}\}$, which satisfies $\|\Delta W_{k}\|_{\infty}\leq\tau$, it holds for any $W_{k}$ following (1) that $\|W_{k}\|_{\infty}<\tau,\,\forall k\in{\mathbb{N}}$.
+Proposition 1 shows that if the magnitude of the update increment $\Delta W_{k}$ is bounded by $\tau$, then the entire trajectory remains within the dynamic range $(-\tau,\tau)$.
+We further establish a stronger result in Lemma 8: with a properly chosen stepsize $\alpha$, the Analog SGD update with asynchronous pipeline in (II-C) guarantees that the trajectory of $W_{k}$ stays within the dynamic range $(-\tau,\tau)$, without requiring any additional assumption on bounded update increments or bounded delay effects.
+However, we note that both Proposition 1 and Lemma 8 are established only for the finite-time setting, so that they do not rule out the extreme limit case where $\lim_{k\to\infty}|W_{k}|=\tau$.
+This is because the proof is built upon a worst-case analysis that accounts only for the gradient magnitude, rather than its direction.
+Under additional over-parameterization assumptions for DNNs [oymak2019overparameterized, liu2022loss, zou2019improved], it is possible to rule out this extreme case.
+However, a formal derivation is beyond the scope of this paper, whose primary focus is on the impact of pipeline parallelism.
+Moreover, in practice, the ratio $W_{\max,\infty}/\tau$ is much smaller than $1$.
+We empirically monitor the saturation degree, defined as $\|W_{k}^{(m)}\|_{\infty}/\tau$, throughout the training process.
+As illustrated in Section IV, the saturation degree for all layers remains remarkably stable and stays well below 10% across the full trajectory.
+Therefore, Assumption 5 is sensible in practice.
 III-B Iteration complexity of pipeline training
@@
 $E_{K}\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S$
-where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}/\tau^{2}}{1-W_{\max,\infty}/\tau^{2}}$.
-The proof of Theorem 1 is deferred to Appendix D.
-Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [37].
-Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [20], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [20, 19, 22].
+where $B$ is the number of micro-batches, and the constant $S$ is given by $S:=\frac{W_{\max,\infty}^{2}/\tau^{2}}{1-W_{\max,\infty}^{2}/\tau^{2}}$.
+Analog-SGD-SP converges in the rate ${\mathcal{O}}(\sqrt{{\sigma^{2}}/({BK})})$, which matches the rate of digital SGD [bottou2018optimization].
+Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error, which also occurs in Analog-SGD-WOP [wu2024towards], and can be mitigated by a well-designed AIMC accelerator or error-correction algorithms [wu2024towards, gokmen2020, gokmen2021].
 Next, we provide the iteration complexity of Analog-SGD-AP.
@@
 Under Assumption 15, if the learning rate is set as $\alpha=\sqrt{\frac{f(W_{0})-f^{*}}{\sigma^{2}L_{f}K}}$ and $K$ is sufficiently large such that $\alpha\leq\frac{1}{L_{f}}$, it holds that
-$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ (12)
+$E_{K}\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{K}}\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$
+$\ +\sigma^{2}S^{\prime}+{\mathcal{O}}\left(\frac{M^{3}(1+1/u)}{K}\right)$ (12)
 where $u>0$ is any number and $S^{\prime}$ denotes the amplification factor given by $S^{\prime}:=\frac{(1+u)W_{\max}^{2}/\tau^{2}}{1-(1+u)W_{\max}^{2}/\tau^{2}}$.
-The proof of Theorem 2 is deferred to Appendix E.
-Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which vanishes for a well-designed AIMC accelerator or can be mitigated by error-correction algorithms [20, 19].
-Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence rate as the Analog-SGD-WOP [20].
+Similar to Analog-SGD-WOP or Analog-SGD-SP, we have asymptotic error $\sigma^{2}S^{\prime}$, which can be mitigated by error-correction algorithms [wu2024towards, gokmen2020].
+Moreover, since ${\mathcal{O}}\left(\frac{1+1/u}{K}\right)$ is not the dominant term, one can set a small $u$ so that Analog-SGD-AP achieves the same convergence order as the Analog-SGD-WOP [wu2024towards] owing to our proof techniques in Section III-C.
 Remark 4.
-As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Theorem 3 in the Appendix F.
-Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [32].
+As a side product, we also provide an enhanced convergence analysis for SGD with asynchronous pipeline training on digital accelerators, detailed in Appendix F.
+Our analysis, based on a novel Lyapunov function, shows that asynchronous pipeline effects appear only in higher-order terms, which does not slow down digital SGD, and the analysis is tighter than [xu2020acceleration].
 III-C Proof sketch of Theorem 2
-In this subsection, we highlight the key steps of the proof towards Theorem 2.
+In this subsection, we highlight the key steps of the proof towards Theorem 2. We only summarize the key lemmas used for analysis. The proofs for each lemma can be found in E.
 (Step I) Smoothness of the objective.
@@
 Using slightly abusing notations, we denote the forward and backward signals computed by $W$ as $x^{(m)}$ and $\delta^{(m)}$, respectively; and denote the ones computed by $\tilde{W}$} as ${\tilde{x}}^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
-The following lemma show that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
+The following lemma shows that for each layer $m$, if we regard the signals as functions of weights, they are smooth functions.
 Lemma 1 (Signal stability).
@@
 where $C_{x^{\prime}}^{(m)}$, $C_{z^{\prime}}^{(m)}$, and $C_{\delta^{\prime}}^{(m)}$ are constants that depend on the DNN architecture.
-Based on this lemma, we prove the objective $f(W;\xi)$ is smooth.
+Based on this lemma, we prove the objective is smooth.
 Lemma 2 (Smoothness of the objective).
@@
 $\forall\xi=(x,y)$, $\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|\leq L_{f}\|W-{\tilde{W}}\|$, where $L_{f}$ is a constant depending on the DNN archetecture.
-Lemma 2 enables us to adapt the tools from [37] to obtain the sufficient descent of the function value.
-However, the bias from the asymmetric update and stale signals from the asynchronous pipelining accumulates layer by layer within the DNN and eventually lead to an extra term in the upper bound.
-Next step we quantify the error of the stale signals.
+Lemma 2 enables us to adapt the tools from [bottou2018optimization] to obtain sufficient descent of the function value.
+However, the bias from the asymmetric update and stale signals from asynchronous pipelining accumulate layer by layer within the DNN and eventually lead to an extra term in the upper bound.
+Next step, we quantify the error of the stale signals.
 (Step II) Error of stale signals.
@@
 However, the stale signal involves the weights at different iterations and introduces additional errors.
-Intuitively, the error of forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal.
+Intuitively, the error of the forward signal ${\tilde{x}}^{(m)}_{k}$ comes from the stale weight $W^{(m)}_{k-(M-m)}$ as shown by (8); while the error of the backward signal arises from the error of the forward signal.
 It motivated us to bound the error by the delay.
@@
 Under Assumption 15, the objective function $f(W_{k})$ satisfies
-$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$ (21)
-$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)+C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
+$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})$ (21)
+$\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$
+$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)$
+$\ +C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
 where $\lesssim$ performs non-essential approximation on the upper bound of descent.
 From Lemma 4, we notice that the accumulated delays $\psi_{k}$ appear in the descent lemma, but intuitively they vanish when the algorithm converges according to its definition.
-(Step IV) Construction of the Lyapunov function. To compensate for this term in equation (21), we design the following Lyapunov function.
+(Step IV) Construction of the Lyapunov function.
+To address the key challenge that the accumulated delay error $\psi_{k}$ appears in the descent lemma (Lemma 4) as a positive term that prevents a clean telescoping argument, we augment the objective $f(W_{k})$ with an auxiliary function $\Psi_{k}$ that satisfies two properties: (i) $\Psi_{k+1}-\Psi_{k}$ produces a term that cancels $\psi_{k}$ in the descent of $f(W_{k})$ while other terms can be also canceled; and (ii) $\Psi_{k}$ itself is non-negative so that it works as an upper bound of $f(W_{k})$.
+Following this, construct a Lyapunov function as
 ${\mathbb{V}}_{k}:=f(W_{k})-f^{*}+C_{\psi}\Psi_{k}$ (22)
@@
 $\Psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (23)
-and $C_{\psi}$ is a constant. To achieve the dynamics of $\Psi_{k}$, we derive that
-$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (24)
-$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
-Let the index of the inner summation of the first term in the RHS of (24) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
+and $C_{\psi}$ is a constant.
+The weighting scheme $(k^{\prime}-k+M)$ in equation (23) is specifically designed to satisfy property (i).
+To see this, we view $\Psi_{k}$ as a weighted version of $\psi_{k}$, which yields a telescoping structure: it assigns larger weights to more recent iterates, reflecting that recent weight updates contribute more to the current staleness error.
+To be specific, the closer $k^{\prime}$ is to the current iteration $k$, the larger the effect of $\Delta_{k^{\prime}}^{(m^{\prime})}$ on the current iteration; if $k^{\prime}$ is farther from $k$, much of its delayed effect has already been consumed, so it receives a smaller weight.
+Enabled by this structure, we can prove the property of $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\left\|W_{k+1}-W_{k}\right\|^{2}$, which cancels the $\psi_{k}$ term in Lemma 4, while the extra $\left\|W_{k+1}-W_{k}\right\|^{2}$ term can be canceled by the $\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)$ in (21).
+To achieve the dynamics of $\Psi_{k}$, we derive that
+$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (24)
+$\ -\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+Let the index of the inner summation of the first term in the RHS of (III-C) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
 $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (25)
-$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}+M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
+$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
+$\ +M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
 $=$ $\ \Psi_{k}+M\|W_{k+1}-W_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
 where the last equality holds by definition of $\Delta^{(m^{\prime})}_{k}$, i.e.,
-$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=$ $\ \sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}.$ (26)
-Plugging (25) into (24) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$.
+$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=$ $\ \sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}$ (26)
+$=$ $\ \|W_{k+1}-W_{k}\|^{2}.$
+Plugging (25) into (III-C) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$.
 We then show that the Lyapunov functions have sufficient descent if $C_{\psi}$ is properly chosen, i.e.,
-$\mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (27)
+$\ \mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}$ (27)
+$\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$
+$\ +\alpha^{3}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sigma^{2}$
+$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (28)
 Reorganizing the inequality above yields Theorem 2.
-Figure 4: Training ResNet models on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
+Figure 3: ResNet10 @ CIFAR10 via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
 (Left) Accuracy-vs-Epoch reflects sample complexity.
@@
 Therefore, we are also interested in the wall-clock time it takes.
-In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\epsilon$.
+In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\varepsilon$.
 Due to asymptotic error, we define the sample complexity in analog training as the minimum micro-batch needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e.
@@
 This shows that three model parallelisms share the same dominant term ${\mathcal{O}}(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while Analog-SGD-AP shows down by a non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$.
-Since the computing is overlapped, the computation density is also needed to yield the number of clock cycles needed to achieve the $\varepsilon$ accuracy.
+Since the computation is overlapped, the computation density is also needed to determine the number of clock cycles required to achieve $\varepsilon$ accuracy.
 Computation density.
@@
 Takeaway. Analog-SGD-AP maximizes the computation density of DNN training only at the cost of a higher-order error term compared with the synchronous pipeline.
-CIFAR10
-speedup accuracy (mid) accuracy (end)
-Digital w.o. pipeline 1.00 88.63  $\pm$ 0.70 94.32  $\pm$ 0.12
-synchronous 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12
-asynchronous 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16
-Analog w.o. pipeline 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07
-synchronous 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07
-asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07
-CIFAR100
-speedup accuracy (mid) accuracy (end)
-Digital w.o. pipeline 1.00 67.85  $\pm$ 0.52 73.99  $\pm$ 0.19
-synchronous 3.69 73.91  $\pm$ 0.19 73.99  $\pm$ 0.19
-asynchronous 8.19 74.64  $\pm$ 0.48 74.64  $\pm$ 0.53
-Analog w.o. pipeline 1.00 61.35  $\pm$ 0.73 72.29  $\pm$ 0.30
-synchronous 3.69 72.14  $\pm$ 0.31 72.29  $\pm$ 0.30
-asynchronous 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
-TABLE II: Training ResNet10 on CIFAR10/CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
+Remark 6 (Load balancing in pipeline parallelism).
+The computation density analysis above assumes that the DNN model is evenly partitioned across $M$ pipeline stages, such that each stage incurs the same computational cost per micro-batch.
+In practice, uneven partitioning would introduce pipeline bubbles at the bottleneck stage, reducing the effective computation density.
+However, this load-balancing issue is orthogonal to the convergence analysis of this paper.
+Theorems 1 and 2, as well as Corollary 1, remain valid regardless of how the model is partitioned, as they bound the sample complexity rather than the wall-clock complexity directly.
+For Corollary 2, a computation density of 1 is an achievable upper bound under balanced partitioning.
+If the partition is unbalanced, the density should be scaled down by a partition-specific factor, which approaches 1 as partitioning becomes more balanced.
+In practice, balanced partitioning can be achieved using existing pipeline-parallelism frameworks such as [huang2019gpipe, narayanan2019pipedream, shoeybi2019megatron]; a detailed investigation of load-balancing strategies for AIMC pipelines is beyond the scope of this paper.
+Remark 7 (Comparison with digital-SGD pipelines).
+In fact, the digital-SGD counterpart is a special case of (1) by forcing $\tau\to\infty$, which removes the asymmetric bias term.
+In particular, the analog-specific denominator inflation, $\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}$ and $\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$, and asymptotic error terms $S$ and $S^{\prime}$ vanish in the digital limit.
+Consequently, the pipelining benefit established in Corollaries 1 and 2 is not unique to AIMC dynamics and can be reproduced in the digital domain.
+CIFAR10 CIFAR100
+speedup accuracy (mid) accuracy (end) speedup accuracy (mid) accuracy (end)
+Digital w.o. pipeline 1.00 88.63  $\pm$ 0.70 94.32  $\pm$ 0.12 1.00 67.85  $\pm$ 0.52 73.99  $\pm$ 0.19
+synchronous 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12 3.69 73.91  $\pm$ 0.19 73.99  $\pm$ 0.19
+asynchronous 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16 8.19 74.64  $\pm$ 0.48 74.64  $\pm$ 0.53
+Analog w.o. pipeline 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07 1.00 61.35  $\pm$ 0.73 72.29  $\pm$ 0.30
+synchronous 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07 3.69 72.14  $\pm$ 0.31 72.29  $\pm$ 0.30
+asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
+TABLE II: ResNet10 @ CIFAR10/CIFAR100 via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn).
 The speedup is computed by comparing the number of clock cycles required to achieve 93% accuracy.
@@
 $128$ 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07 4.89 93.36  $\pm$ 0.06 93.37  $\pm$ 0.06
-TABLE III: Training Resnet10 model on CIFAR10 dataset under different micro-batch size setting.
+TABLE III: ResNet10 @ CIFAR10 under different micro-batch size setting.
 Test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
@@
 In this section, we aim to numerically verify the theoretical benefits of Analog-SGD-AP.
-We test the analog methods in an open-source AIHWKIT (AIHWKIT[38]; see github.com/IBM/aihwkit.
+We test the analog methods in an open-source AIHWKIT (AIHWKIT[Rasch2021AFA]; see github.com/IBM/aihwkit.
 The pipeline parallelism is also simulated on a single GPU.
@@
 The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage numbers $M=12,16$, the initial learning rate is $\alpha=0.01$.
-The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training.
 The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs.
-To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [39].
-The other setting follows the default settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits).
-Noise and bound management techniques are used in [40].
-A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
+To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [cubuk2018autoaugment].
+AIMC hardware configurations.
+The simulations use the SoftBoundsReference device in AIHWKIT, which directly instantiates the asymmetric update model in (6) with $\tau=0.9$ by default.
+Beyond the asymmetric update, the simulations additionally incorporate a comprehensive set of hardware imperfections to more accurately reflect real device behavior, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits).
+In addition, 30% device and cycle variation are introduced.
+Noise and bound management techniques are used in [gokmen2017cnn].
+A learnable scaling factor is applied after each analog layer and updated via digital SGD.
 Sample complexity and wall-clock complexity.
-We train a ResNet10 for an image classification task on CIFAR10.
+We train a ResNet10 on CIFAR10.
 We map the first 5 stages on digital accelerators while the last one onto an AIMC (analog) accelerator.
 The results on the left of Fig.
-4 show that the accuracy-epoch curves of all three pipeline strategies on AIMC accelerators almost overlap and even match the performance of digital SGD, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1.
+3 show that the final accuracy differences between the three strategies are below 1%, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1.
 Moreover, as indicated by Corollary 2, the walk-clock complexity of the Analog-SGD-AP pipeline is the lowest owing to the highest computation density.
@@
 Effect of stage/accelerator number.
-We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Figure 4, Right.
-In the simulations, we train a ResNet34 model [41] on the CIFAR10 dataset.
+We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Fig.
+3, Right.
+In the simulations, we train a ResNet34 model [wu2019wider] on the CIFAR10 dataset.
 The model is separated into $M=1,2,4,8,12,16$ stages.
@@
 The results reveal that the asynchronous pipeline achieves linear speedup at least within the range of 181-8 accelerators.
+As the $M$ increases, the speedup deviates from linear scaling, which can be explained by Theorem 2.
+As shown by (2), the iteration complexity of Analog-SGD-AP contains a higher-order term $O(M^{3}(1+1/u)/K)$, whose coefficient depends on $M$.
+Although this term is higher-order and does not affect the dominant sample complexity $O(\sigma^{2}\varepsilon^{-2})$, it has a non-negligible practical impact when $M$ increases while iterations $K$ remain small.
+Consequently, it results in the observed deviation from linear speedup.
+Developing acceleration strategies to mitigate this depth-dependent overhead and recover linear speedup at larger scales ($M>8$) is an interesting direction we leave for future work.
 Effect of micro-batch size.
 We also report the effect of micro-batch sizes in Table III.
-We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size.
-For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog accelerators.
-Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version.
+We set the mini-batch size to $B_{\text{mini}}=128$ and vary the micro-batch size.
+For synchronous pipeline, increasing the batch size improves the speedup on both digital and analog accelerators.
+Moreover, asynchronous pipeline training consistently yields greater speedup than the synchronous version.
 The trade-off is that the asynchronous pipeline sacrifices accuracy and can only reach lower accuracy when the training time is sufficiently long.
+Effect of hardware imperfections.
+To disentangle the contributions of analog imperfections and asynchronous pipeline staleness to the convergence error, we compare the performance of digital and analog pipelines under identical configurations, as shown in Tables III and III.
+The results in Table III show that (i) the accuracy gap between digital and analog is consistent ($1\%-2\%$) across all three pipeline strategies; (ii) within each hardware setting (digital or analog), the accuracy difference between synchronous and asynchronous pipelines is negligible at convergence ($1\%$).
+These results indicate that pipeline staleness introduces only marginal additional error, and the gap is attributable to analog imperfections.
+Table III further shows that this conclusion holds robustly across different micro-batch sizes: the accuracy gap between digital and analog remains stable regardless of micro-batch size, while the gap between synchronous and asynchronous pipelines remains small in both settings.
+Taken together, these results confirm that analog imperfection is the dominant source of convergence error, while the asynchronous pipeline staleness has a secondary effect.
+Effect of asymmetry parameter $\tau$.
+We further examine how the asymmetry parameter $\tau$ affects both convergence and acceleration; see Fig.
+4a/b.
+The left figure shows that when $\tau>0.5$, Analog-SGD-AP still achieves above 90% test accuracy.
+Furthermore, as shown in the right figure, Analog-SGD-AP maintains a speedup well above the synchronous pipeline baseline (shown as the dashed line at $3.69\times$) for all $\tau\geq 0.4$.
+This confirms that the wall-clock gains of the asynchronous pipeline are preserved across a wide range of $\tau$ values.
+In contrast, as $\tau$ decreases, training quality degrades significantly, with asynchronous training affected more strongly.
+Figure 4: Analog-SGD-AP for CIFAR10 with $B_{\text{micro}}=16$ and delay =0=0.
+(a) Test accuracy versus $\tau$.
+(b) Speedup versus $\tau$.
+(c) Speedup versus device-to-device variation.
+(d) Saturation degree evolution over the trajectory.
+Effect of device variation on speedup.
+We next evaluate whether device-to-device variation changes the acceleration trend; see Fig.
+4c.
+In the simulation, we set $\tau=1.0$, $B_{\text{micro}}=16$, and delay =0=0.
+Across the typical variation levels tested here, the asynchronous pipeline speedup changes only mildly, suggesting that device variation is not the primary limiter of pipeline acceleration in this regime.
+Saturation degree evolution along training trajectory.
+We examine the saturation degree over the full training trajectory to justify the bounded saturation condition; see Fig.
+4d.
+The saturation degree of $m$-th layer is defined as $\|W_{k}^{(m)}\|_{\infty}/\tau$.
+To ensure the formulation is well-defined, we set the device variation to 0.
+The results show that the saturation remains nearly constant throughout training, with variations less than 10%.
+This observation is consistent with existing results, such as [zou2019improved], which demonstrate that the trajectory is bounded in a neighborhood of the initial point.
+It provides empirical evidence for the Assumption 5 in the studied configuration; a general theoretical characterization is left for future work.
 V Conclusions and Limitations
@@
 Future work should verify its efficiency as the technology becomes mature.
-Appendix
-Appendix A Analog SGD with Asynchronous Pipeline in Different Perspective
+Supplementary
+Appendix A Analog SGD with Asynchronous Pipeline in Different Perspectives
 Involving multiple accelerators and data at each time step, pipeline algorithms, especially the asynchronous pipeline, are usually hard to figure out.
@@
 In data perspective, we focus on how a sample $\xi_{k}$ flows through each accelerator; see Algorithm 2.
-To make the relationship clearer, the computation dependency is plotted in Fig.
-3.
+To clarify the relationship, the computational dependency is plotted in Fig.
+5.
 Algorithm 1 Analog SGD with asynchronous pipeline (accelerator Perspective)
@@
 12:end for
+Figure 5: Illustration of the dynamics of the asynchronous pipeline.
+The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each accelerator.
 Appendix B Useful Lemmas and their proof
@@
 $\overset{(a)}{\leq}$ $\ L_{g,0}(\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}\|+\|W^{(m)}_{k}\|_{2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
-$\leq$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
+$\overset{(b)}{\leq}$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
 where (a) follows the triangle inequality and the consistency of $\ell_{2}$-matrix norm, i.e.
-$\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively.
+$\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively; and (b) is because of Lemma 8.
 Using triangle inequality over the first term of RHS in (35), we have
@@
 Lemma 7 (Lipschitz continuity of loss function).
-There exists $L_{\ell,0}$ such that for any $m\in[M]$, $\|\nabla_{x^{(m)}}\ell(x_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$.
+Under same conditions with Lemma 8, there exists $L_{\ell,0}$ such that for any $m\in[M]$, $\|\nabla_{x^{(m)}}\ell(x_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$.
 Proof.
-According to norm equivalence, there exists $W_{\max,2}=cW_{\max}$ with $c$ is defined by the dimension of $W_{k}^{(m)}$, such that $\|W_{k}^{(m)}\|\leq W_{\max,2}$.
+According to norm equivalence, there exists $W_{\max,2}=cW_{\max}$ with $c$ is defined by the dimension of $W_{k}^{(m)}$, such that $\|W_{k}^{(m)}\|\leq W_{\max,2}\leq\tau$.
 For $m\in[M]$, according to the forward recursion (3), it holds that
@@
 ∎
-Lemma 8 (boundeds signal).
-Under Assumption 24, both forward and backward signals are bounded, i.e., $\forall m\in[M+1]$
+Lemma 8 (Bounded signal and iterates).
+Under Assumption 24, selecting stepsize $\alpha\leq\frac{\tau}{(L_{g,0}\tau)^{M}L_{\ell,0}L_{g,0}}$ and initializing $\|W_{0}^{(m)}\|<\tau,\forall m\in[M+1]$, then both forward and backward signals are bounded, i.e., $\forall m\in[M+1]$
 $\|x^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m},$ $\|\delta^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0},$ (54)
 $\|{\tilde{x}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{m},$ $\|\tilde{\delta}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}.$ (55)
+Moreover, for any $k\leq K$ and $m\in[M+1]$, the iterates $\|W_{k}^{(m)}\|<\tau$ are bounded within the dynamic range, i.e. $W_{\max,2}:=\max_{k\leq K,m\in[M+1]}\{\|W_{k}^{(m)}\|\}<\tau$.
 Proof of Lemma 8.
-This lemma is proved by induction over $m$.
-The statement for $\|x^{(m)}_{k}\|$ and $\|{\tilde{x}}^{(m)}_{k}\|$ at $m=1$ trivially hold because $x^{(0)}_{k}={\tilde{x}}^{(0)}_{k}=x_{k}$.
+This lemma is proved by induction over $m$ and $k$. We can initialize such that $\|W_{0}^{(m)}\|<\tau$ holds for any $m\in[M+1]$.
+At each iteration $k$, assume $\|W_{k}^{(m)}\|<\tau$ holds for any $m\in[M+1]$ and denote $W_{\max,2}^{(m)}=\max_{m}\{\|W_{k}^{(m)}\|\}$.
+Clearly, $W_{\max,2}^{(m)}<\tau$.
+Then the statement for $\|x^{(m)}_{k}\|$ and $\|{\tilde{x}}^{(m)}_{k}\|$ at $m=1$ trivially hold because $x^{(0)}_{k}={\tilde{x}}^{(0)}_{k}=x_{k}$.
 For $\|\delta^{(M+1)}_{k}\|$ and $\|\tilde{\delta}^{(M+1)}_{k}\|$, we have
@@
 $\|x^{(m+1)}_{k}\|\overset{(a)}{=}$ $\ \|g^{(m)}(z^{(m)}_{k})-g^{(m)}(0)\|\overset{(b)}{\leq}L_{g,0}\|x^{(m)}_{k}\|=L_{g,0}\|W^{(m)}_{k}x^{(m)}_{k}\|$ (58)
-$\leq$ $\ L_{g,0}\|W^{(m)}_{k}\|\|x^{(m)}_{k}\|{\leq}L_{g,0}W_{\max,2}\|x^{(m)}_{k}\|$
+$\leq$ $\ L_{g,0}\|W^{(m)}_{k}\|\|x^{(m)}_{k}\|{\leq}L_{g,0}W_{\max,2}^{(m)}\|x^{(m)}_{k}\|$
 where (a) and (b) hold because of Assumption 3 (1) and (2). According to the backward recursion (4), it holds that
 $\|\delta^{(m)}_{k}\|$= $\ \|\delta^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}(z^{(m)}_{k})\|$ (59)
-$\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}\|$,
+$\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}^{(m)}\|\delta^{(m+1)}_{k}\|$,
 $\|\tilde{\delta}^{(m)}_{k}\|$= $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (60)
-$\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\tilde{\delta}^{(m+1)}_{k}\|$.
-The proof has finished. ∎
+$\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}^{(m)}\|\tilde{\delta}^{(m+1)}_{k}\|$.
+Therefore, by induction over $m$, we obtain,
+$\|x^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{m},$ $\|\delta^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{M-m}L_{\ell,0}L_{g,0},$ (61)
+$\|{\tilde{x}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{m},$ $\|\tilde{\delta}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{M-m}L_{\ell,0}L_{g,0}.$ (62)
+On the other hand, if we select the stepsize $\alpha$ such that $\alpha\leq\frac{\tau}{(L_{g,0}\tau)^{M}L_{\ell,0}L_{g,0}}$, we have the magnitude of each update increment
+$\|\Delta\tilde{W}_{k}^{(m)}\|=\alpha\|{\tilde{x}}^{(m)}_{k}\|\|\tilde{\delta}^{(m)}_{k}\|\leq\tau.$
+Owing to the induction nature proof for Theorem 1 (see [wu2024towards, Theorem 1]), we know $\|W_{k+1}^{(m)}\|<\tau$ holds for any $m\in[M+1]$.
+By induction, the proof is completed. ∎
 Appendix C Smoothness of the object
@@
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
+$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}].$
 We bound the last term of the RHS of (110) by
@@
 $\ -\frac{1}{4\alpha}\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}+\alpha(\nabla f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
-where the second inequality is earned by plugging in (F). Construct a Lyapunov function as
+where the second inequality is earned by plugging in (F).
+Construct a Lyapunov function as
 ${\mathbb{V}}_{k}:=f(W_{k})+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M-m^{\prime}+2)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$ (119)